A Selected Bibliography of Publications by, and about, George E. Uhlenbeck

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 April 2024 Version 1.07

Title word cross-reference

(α, q) [Sza12, Sza13]. (N, d) [WC97]. $12.50 [Uhl63a]. $13.75 [Bri62, Uhl60a]. ^{6} [WU34b]. α [HP14, Pat07, ZZ13]. β [Fra05, KU35a, KU41]. N = 2 [Mas15, Mas16a]. D [KS08, WX15]. e/m [UY30]. ℓ [KLS16]. γ [RU35]. H [RD05]. l^{2} [CL90, CCS92, IM89, IMM^{+}90, CLS93]. L^{p} [Fat09, Sch90]. L_{1} [KP94]. l_{p}, p ∈ [1, ∞[ [Fer92]. μ [UC52]. N [KL06, KL08, RE13]. p > 0 [Fat09]. q [Sza12, Sza13, Wan16]. S^{a} [Ive10]. U [AM14, Har05]. U(3) [ASA89]. Z^{a} [Ive10].

-conformal [KLS16]. -conjecture [Har05]. -Dimensional [WX15]. -Functionals [Fat09]. -Meson [UC52]. -norm [CL90, CLS93, KP94]. -optimal [KS08]. -Ornstein [Sza13, Sza12, Wan16]. -Particle [KL06, KL08, RE13]. -Radiation [RU35]. -Radioactivity [KU35a, KU41]. -Stable [Pat07, HP14, ZZ13]. -Statistics [AM14]. -theorem [RD05]. -Valued [IM89, IMM^{+}90, CCS92, Sch90]. -Wiener [Sza12, Sza13].

1 [McL62]. 1939 [Uhl40]. 1958 [Uhl59b]. 1960 [Kac60]. 1962 [McL62]. 1964 [Mei65]. 1969 [WH72]. 1972 [Meh73]. 1988 [dB89].

20th [Meh73].

Aachen [Mei65]. aan [van35]. aangeboden [van35]. absolute [Sim11]. Absorption [Dir75]. Academy [WH72]. accept [Ano54a]. According [EU27a, WU34a]. Accuracy [Bis01]. accurate [Yan17]. achievements [Woo80]. Action [PU50, SSMS17]. Adaptive [GMS93]. Added [Ano38b]. Addison [Uhl59b]. Addison-Wesley [Uhl59b]. Additive [MM05]. Admission [BPS15]. Aeronautics [Ano74]. Again [Rai85]. Age [Sul78]. Aggregation [OV04]. Albert [Woo80]. Aléatoires [Fer92, Fer90]. algebra [KLS16]. algebras [And14a]. Almost [KL06, KL08, Xin12]. Alte [Uhl62a]. alternating [CU70a]. alternative [DMTUV09, Mas16b]. American [WH72, SK13, Rom91]. Ammonia [DU32]. Amsterdam [Bri62, Uhl60a, Uhl63a]. Analysis [CK19, LU31, Mia13, Sim91, WZ11]. analytical [CF17]. Angular [FU50a]. Anharmonic [MM05]. Anomalous [Ben38]. Ansatz [KU35b]. Answered [Lau56]. Application [Eng10, LU31]. Applications [CT73, S.64, BCNP15, ELY99, Fas13, PSLL15, UF62, WSW15, ZWB17]. Appointments [Ano38b]. approach [ACVV17, BBD^{+}14, CF17, DMTUV09]. Approximate [YU30, BB01, Bis06]. Approximation [CP07, LP00, Bis01, Uhl60c]. Approximations [Tho75]. Arts [WH72]. aspects [Uhl62b]. Assistants [Ano38a]. associated [BDT11]. Associates [Ano38a]. Astronautics [Ano74]. Asymptotic [APvZ01, Bee75, BW94, Fas10, HA13, Kaw13, KK11, Sim95, ZZ19, vB09, Shi12, Xin12]. Asymptotics [Fat09, Kol01, WX15, HP14, Jeg09]. atom [GU26a, vCH18, GU26a]. Atomic [Ano54b, Lau56, RU53]. atoombom [vC16]. Attraction [LS00]. autocorrelation [HA13]. autocovariance [Fas10]. autoregression [APvZ01]. Average [Lac05]. Award [Ano54b, Ano54a].

B [Bus60, Uhl60a]. Banach [ÁLBRM16, Mas07, Sch93]. bankruptcy [WZ11]. Barnard [Ano38a]. based [DNO18, RS12]. basic [Uhl50]. basis [Jør78]. Bayes [Bis01]. Bayesian [GS10, RPD04]. Be [Ano38b, Lau56, Ano54a]. before [BS85]. behavior [CS95, FZ99, UG25, UB37]. Behaviour [Ben05, ZZ19]. behoud [Uhl36]. Berry [ES13, Jia12]. Bessel [DMRYZ04]. Beta [FU50a, KU36]. Beth [BBD^{+}14, LLAT16]. Between [Uhl80a]. bezüglich [UG25]. biography [Coh90b]. Birth [BS85]. Boer [Bri62, Buf63, Koz66, McL62, Rei63]. Boltzmann [CT73, EU27a, FU70b, FU76, Uhl55a, Uhl73b]. Boltzmannschen [EU27a]. bomb [vCH18]. Book [Bri62, Buf63, Koz66, McL62, S.64, Uhl40, Uhl41, Uhl59b, Uhl60a, Uhl63a, Uhl73a]. Bose [SK99, UU33, UG32, UU32, ZUK77]. Bougerol [Vak12]. Boulder [Kac60, KU59]. Bound [Tho11, ZZ16]. Boundaries [BWY11]. Boundary [CKS99, MS09b, RS87, Roc92, Mor94]. Bounds [Jia12, Bor85, ES13]. branched [SSMS17]. Branching [AM14, Eng10, Li00, AM15a, CM18, EH82]. bridges [Maz17]. Brillouin [YU30, HM75]. Broglie [EU26a]. Broglieschen [EU26a]. Brownian [BW94, CP07, DMLF^{+}17, FMW06, Igl05, Jør78, UG29, UO30, Vak12, WU45b, WU45c, WU45a, vLU31]. buffer [CS95].

C [Uhl40]. calculation [VU37, Yan17]. Canonical [SSMS17, DMTUV09, RU50, Uhl35]. Capacitors [EK05]. Capacity [CKS99]. Carlo [Gri68]. case [Sus89]. Casimir [Uhl60a]. Catalyst [FMW06]. catastrophes [GNdC12]. Cauchy [GO00, Min92a, Pol90, Sus86]. caused [Mia13]. celebrate [Woo80]. centennial [Woo80]. Center [Eng10, Gil13]. Centre [Meh73]. century [Meh73, FW60]. Certain [DMM62]. certaines [Fer90]. Chang [HM75, Sté03, Ani17, Che02, Del03, FH66]. change [CMN18]. change-point [CMN18]. Chapman [Uhl41]. Characteristic [Vri16, Mit87]. Characterization [Har05, SG15]. Characterize [Tay89]. Charged [PU49]. Check [FK89, FK90]. chemistry [BS85]. chiral [ASA89]. Choh [GCF66]. circle [Nor98]. claim [Sch14]. Class [PS05]. Classes [BOGP12, XZW09, GP11]. classical [BP07, UB36, Uhl60c]. Clinical [Ano38a]. CLT [AM15a]. cluster [LLAT16]. Co [Sul88, Fas13]. Co-Discoverer [Sul88]. co-integration [Fas13]. Codiscoverer [Sul78]. Coefficient [PU59, LLAT16]. coefficients [Mit87]. coherent [Bag11, PG13]. Collected [Uhl60a]. College [Ano38a, Ano38b]. Colleges [ESU68]. collision [FH66, Sus86, Sus89, Sus93]. Colorado [Kac60, KU59]. Columbia [Ano38a, Ano38b]. Combinatorial [FU56a, FNU56, FU56b, FU57]. Comment [Cer73]. Comments [BCG13, Smi09]. commodity [ZGO12]. common [DG13]. Company [Uhl59b]. Comparison [DM72]. Complete [Ano38a]. complex [ABI99]. complexified [DMTUV09]. Component [Nov04]. concept [Ano54a, BS85, SK99]. conception [Meh73]. condensation [KU37, KU38, SK99]. Conditional [Sal84]. Conditions [WTZ15, Mor94]. Conductivity [CUdB64]. Conference [Lau58]. conferences [WH72]. Conformal [And14a, AGGM14, KLS16]. conjecture [Har05]. connected [Row86]. connection [EU26b]. conservation [Uhl36]. Consistency [ÁLBRM16]. constant [CS95]. Constraint [Maz17]. Construction [Jør78, KP04, TM08]. contingent [Sch14]. contingent-claim [Sch14]. Continuity [IMM^{+}90, LiS09, CCS92, CL90, CLS93, MS92, Sim11]. continuous [GS10, Sch14]. continuous-time [Sch14]. Continuum [WTZ15]. contrast [Bis06]. Contributions [FU70a, FU70b, FU76, SU42]. Control [FP03, BPS15]. Convergence [Bis08b, EM05, Fan04, AHR17, BB01, Bis06, OV04]. Conversion [RU35]. convex [BDT11, Har05]. Corp. [Ano54a]. Corporation [Ano54b]. corrections [Row86]. Correlated [FK02]. Correlation [FU50a, FU50b, Har05]. Cosmic [SU43, NLU40, SU42]. cosmic-ray [NLU40, SU42]. cosmological [PSLL15]. Couette [Pie91]. Coupling [Del03, Wan11a, GU26a]. Course [Uhl59b]. Covariant [UL31]. Cowling [Uhl41]. Cox [CW93]. credit [CS09]. criterion [Shi90]. Critical [Deu89, HKU64, SK99, UH65, Uhl78]. Crossings [CKS99]. curves [JM14]. Cut [Lac05]. Cut-off [Lac05].

D [Ano38a, Ano38b, Uhl59b]. Dans [Fer92]. Dead [Sul78]. Death [CW93]. Decade [Coh71]. Decay [Fra05, KU36]. december [dB89]. Decomposability [App06]. decomposition [KYH16, MS09b]. definition [PS09]. Deflection [UY30]. Deformation [BMMR15, KLS16]. degree [JDD13]. demeaned [Ai16]. Dense [GCF66, BBD^{+}14, UF63a]. Densities [KK11, NRS85, RS87, Jeg09]. Density [RS88, Tho11, VU37, XXY12]. dependence [IT99]. dependent [CS95, CKX09, KYH16]. derivation [dBvW78]. derivative [Üst95]. Derivatives [KU35b, vB09]. designs [BSS13, BS15, KS08]. Deutons [WU34b]. Development [Meh73, S.64, RU53, UF62]. Deviation [GJ09a, Gas11]. Deviations [BR02, BCS11, Deu89, FLP99, Hec99, IM89, dRdC17, BR15, BR17, Bis08a, DMRYZ04, GJ09a, GJ09b, UB36]. Dies [Sul88]. different [EU26b]. differential [FK11]. Diffusion [DR98, FK02, Kol01, Mor94, AHR17]. Diffusions [Sim95]. Digit [LG78, Rai85, Rai76]. dilemma [SK99]. dilute [Uhl74, Yan17]. dimension [Bak82]. Dimensional [EU26a, GZ10, KUH63, WX15, AC87, App15, EU26b, Sha93, Uhl25]. dimensions [MS92]. diquarks [BBD^{+}14]. Dirac [LU31, MD05, Sus89, UU33, UL31, UG32, UU32]. Directional [FU50b]. Dirichlet [Sch93]. Discoverer [Ano54a, Sul88]. Discoveries [Lau58]. discovery [Gou71, Sul78]. Discrete [BS98, KM12, Kaw13, Ren87, Glo01, KM11, Mia13, Min92b, Min92a]. discretely [Bis06, ES13, Lon09, Ma10, Shi12, ZZ13]. Discussion [HKU64, KUH63, UHK63]. Disintegration [WU34a, WU34b]. distance [DY10]. Distribution [BD13, Dan91, DMM62, JP99, KK11, MYZ67, Sim95, Yur17, AKS96, BG91, PS09, TL09, UG35, UHK63, ZW17]. distribution-valued [BG91]. Distributions [Bee75, Fat09, LS00, PS05, XZW09, Lef97, LiS09]. divergence [Üst95]. dividends [WZ11]. divisibility [LiS09]. Division [McL62]. Does [SK99]. Domain [LS00]. D’Ornstein [Fer92, Bak82, Fer90, Jac85]. doubly [SWW15]. Dr. [Ano38b]. Drift [AM15b, Jia12, GJ09a, HW97, JD15, QH95]. drifts [FZ99]. Driven [BDY07, KR94, MM05, Ped02, PS05, Wan12, Abd15, AIK18, App15, BPS14, BN08, DY10, FK11, HP14, HLT16, JVV05, Mai14, Mas04, MY13, TL09, Vri16, Wan11b, ZZ13, ZWB17]. durch [UG25]. during [KU36]. Dutch [Ano54b, GU25, GU26b, Uhl24, Uhl25, Uhl27, Uhl36, Uhl55b]. Dying [Kol01]. Dynamics [Ano74, Smi09].

E. [Ano38a]. each [UG25]. early [ZGO12]. early-exercise [ZGO12]. easy [SK99]. Economics [Ano38a]. Ed [Uhl60a]. Edgeworth [MY13]. Editorial [Rom91]. Editors [McL62]. Eds [Bri62]. effect [Dio16, Mia13]. effects [Mor94]. Efficiency [Str58]. Efficient [Mai14, ZGO12]. Ehrenfest [Uhl60a, MUG^{+}33, Uhl56]. Eigenfunction [AC87]. Eigenfunctions [Tay89]. Einstein [EU27b, SK99, UU33, UG32, UU32, Uhl80b, Woo80, ZUK77]. Einsteinschen [EU27b]. einzelnen [UG25]. Elastic [vLU31]. electric [Ive10]. Electromagnetic [PU49]. Electron [Ano54a, RU35, Sul78, Sul88, Tho26, GU26b, Gou71, UG25]. Electrons [FU33a, UG26, UG67, UGB84, BS85, FU33b]. Elektrons [UG25]. elettroni [FU33b]. Ely [Ano38b]. Emission [KU36]. Empirical [Deu89, Hec99]. energie [Uhl36]. energies [Pav16]. Energy [KL15, Seg85, UG35, Uhl36, VU37]. ensemble [RU50, Uhl35]. Entropy [WX15]. Envisions [Lau56]. equality [Har05]. Equation [CT73, DR98, HM75, KL06, KL08, UG32, Uhl55a, Uhl73b, YU30, Ani17, BDT11, Che02, EU26b, EMV07, EMV08, FU70b, FU76, Min92b, Min92a, Pie91, PSV18, RU53, Sus86, Sus89, Yan17, dBvW78]. Equations [Kol01, LU31, Mas07, UL31, AL97, Del03, FK11, Pie92, Sté03, Sus93, dBvW78]. Equidistant [KS08]. Equilibrium [CM97, HKU64, KUH63, DMLF^{+}17, FU70a, Mei65, PU69, UHK63]. Equivalence [CW93]. ergodic [Tra17]. Ergodicity [HLT16, Wan12]. Ersetzung [UG25]. Esseen [Jia12, ES13]. estimate [KP94]. estimated [Abd15]. Estimates [CKS99, Str58]. Estimating [Gri68]. Estimation [BPvZ03, BDY07, FLP99, HKST17, WTZ15, AHR17, AM15b, BY12, Dio16, DY10, Fas13, Glo01, HN10, JD15, Mai14, ZZ16]. Estimator [ZZ19, Bis01, ES13, Lon09, Ma10, ZZ13]. Estimators [Bis08b, Jia12, AKS96, Bis01, BB01, Bis06, GJ09a]. Euclidean [ASA89]. Eugene [Cas88, OR01, dB89, vB13]. Euler [AL97, Del03]. European [Sul78]. Evolution [Kol01]. Exact [AKS96, Fat09, KM11, KK11, Zha11]. exactly [CM18]. examples [KMiS06]. exercise [ZGO12]. Existence [Sus89, Sus93]. Exit [Sim95, BN08, Pat07]. Expansion [Igl05, MY13]. Expansions [LP09, Ai16, AC87]. Experiments [UY30]. Explained [Lau58]. Explaining [Lau56]. Exploring [WH72]. Exponential [NRS85, Wan12, KMiS06, ZWB17]. Exponentials [MS09a]. extension [DMTUV09, FK89, FK90, Uhl25]. extremes [Fas10].

F [Uhl59b]. Faculty [Ano38b]. family [KP04]. Fast [Kol01, Yan17]. features [BCG13]. Fermi [KU35a, KU35b, KU41, Sus89, UU33, UG32, UU32, WU34a]. Fernique [CCS92]. Ferromagnets [Dys69, Dys71]. Ferziger [Uhl73a]. Field [CM97, PU49, PU50, JDD13]. Fields [CGXM96, Fan04, CGXM97, CU70a]. Fifty [Uhl76]. Figures [GF44]. filtering [CKX09]. filtration [Jac85]. financial [WSW15]. Finite [Gri68, KM11]. Finite-Time [Gri68]. First [BWY11, Lef97, LG78, NRS85, Nov04, Rai85, RS88, SWW15, Tho75, Tho11, Dit07, HP14, Rai76, Vee15]. First-Passage [Nov04, Tho75, Dit07]. First-Passage-Time [NRS85, RS88]. fit [AIK18]. Fitting [ZZ19]. Five [EU26a]. Five-Dimensional [EU26a]. Flow [Ben05, Pie91]. Flows [FK02, Fan04, vB09]. Fluctuation [FU76, NLU40, SU42, FU70b]. Fluctuations [FMW06, GZ10, FU70a]. Fluid [KR94, Sim91, Uhl80a, CS95]. Fluids [ESU68]. Fokker [KL06, KL08]. Following [UL31]. Fonctions [Fer92, Fer90]. Forbidden [KU41]. Force [CM97, CU70a, UG25]. Ford [FK89, FK90, Row86, S.64]. Forderung [UG25]. form [Min92b]. formalisms [And14b]. Formation [RU35]. forms [Sch93]. formula [LLAT16, SG15, Vee15]. formulation [Mas16b, Mos10, SSMS17]. four [Pie91, Pie92]. four-velocity [Pie91, Pie92]. fourth [BMMR15, MD05]. Fractional [BCS11, CKM03, CK19, FK11, Igl05, MS09a, AM15b, Bis08a, BPS15, CP07, ES13, HN10, SG15]. Fragen [Uhl62a]. fragmentation [BB15]. Frames [LP09]. free [Mar81]. freedom [JDD13]. French [Bak82, Fer90, Fer92, Jac85]. Frontiers [Lau58]. function [Vri16]. Functional [DMM62, ÁLBRM16]. Functionals [ABI99, Fat09, KK11]. Functions [DM72, DG13, Fas10, Fer90, Fer92, Har05, Pat07, UHK63]. Fundamental [Coh68, EMV07, Uhl60b]. fünfdimensionalen [EU26a]. Furry [NLU40]. Further [SU42].

G [Ano38b, Bri62, Buf63, Bus60, Koz66, McL62, Rei63, S.64, Uhl41, Uhl60a, Uhl73a]. G. [Ano89, Coh90b]. Galilei [KLS16]. Gallery [Pai00]. Gamma [FU50a, IT99, KU36]. Gamma-Angular [FU50a]. gamma-mixed [IT99]. Gangolli [Jør78]. Gas [Ano74, UG32, UU32, CU70a, UB36, UB37, UH65, Yan17, ZUK77]. Gases [CUdB64, CUFF70, FU67, GCF66, UU33, CU70b, CU70c, FUF72, RU53, UF62, UF63a, Uhl74, S.64, Uhl41, Uhl73a]. gauge [JDD13]. Gaussian [BG91, CL90, GZ10, GS06, GS10, Gri68, LP09, RS12, RPD04]. genealogy [CM18]. general [Mas04, RD05, ZZ16]. généralisé [Jac85]. Generalised [EM05, CMN18]. Generalization [GCF66, Sza12, Sza13, CU70a]. Generalized [And15, BOGP12, BMW^{+}06, BBD^{+}14, Cha90, HS81, LZ16, Mit87, Yur17, BY12, EM08, Fas10, Ive10, LLAT16, LiS09, Pat07, Wit84]. generated [CL90]. generation [FK89, FK90]. generators [CMG02]. Genius [Pai00]. Geometric [Bor85]. George [vB13, Cas88, Coh90a, Dre89, OR01, Sul88, dB89]. German [EU26a, EU27b, EU27a, GU26a, UG25, Uhl62a]. Germany [Mei65]. Gets [Ano38a]. ghost [DMTUV09]. Gideonse [Ano38a]. given [TL09]. Gives [Ano54b]. Goal [Lau56]. Goodness [AIK18]. Goodness-of-fit [AIK18]. Goudsmit [Ano54a, BS85, Ive10, Sul78, vC16, vCH18]. Graphical [EU26a]. Graphische [EU26a]. Graphs [S.64, FK89, FK90, FU56a, FNU56, FU56b, FU57, Row86, UF62]. gravity [Mar81, oT06]. Gross [Üst95]. Grossissement [Jac85]. Growth [vB09].

H [Uhl60a, Uhl73a]. H. [Ano38a]. haar [Uhl25, TM08]. Haar-like [TM08]. Hamiltonian [And14b, Kol01, Mas16b, Mos10, SSMS17]. happened [Ano54a]. harmonic [LP00]. Harnack [Wan11b]. Harvard [Lau58]. having [Mit87]. Heat [CUdB64]. Heavy [BD13, BPS15]. Heckius [Uhl24]. Held [Meh73, Mei65]. helium [GU25]. Hibbs [Bus60]. hierarchical [HA13]. Higher [KU35b, PSLL15]. Hilbert [ÁLBRM16, App06, BDT11, FP03, Jeg09, Liu08, Liu12, LZ16]. Historical [Seg85, Uhl78]. history [KHFA67, WH72]. Hitler [vC16, vCH18]. Hitting [Lac05, Vak12, Lef97, SWW15]. Hodgkin [HLT16]. Holland [Bri62, Uhl60a, Uhl63a, Ano38b]. homogeneous [Sus86]. Honored [Ano54a, Ano54b]. Hooke [And14a, AGGM14]. hot [BBD^{+}14]. hunt [vCH18]. Husimi [HU53]. Huxley [HLT16]. Hydrodynamic [FMW06]. hydrodynamical [FU70a]. hydrogen [Ani17, GU25]. Hypercontractivity [CS08, HW97, QH95]. hyperfine [Bel82]. hypergeometric [Pat07]. Hypothese [UG25]. Hypotheses [LS19]. hypothesis [Ive10, UG25].

idea [Ano54a]. Ideal [UU32, UG32, UB36, UB37, ZUK77]. identity [Vak12]. II [Koz66, BDT11, FNU56, FU70b, FU76, KU41, PKU72, SU42, UB37, UHK63, WU45b, WU45c, WU45a, dBvW78]. III [FU56b, HKU64]. Illus [Uhl63a]. Illustration [EU26a]. im [GU26a]. Immigration [Li00]. importance [ELY99]. indexed [BH12, Nor98]. Indirect [RS12]. inequalities [CCS92, GJ09a, Wan11b]. inequality [Har05]. inertial [Mor94]. Inference [CMN18, GS06, RPD04, APvZ01, GS10, JVV05, RS12]. infinie [Bak82]. Infinite [App15, KM12, AC87, Bak82, BB15, Cha90, LiS09, MS92, Sha93]. infinite-dimensional [Sha93]. infinite-particle [Cha90]. inhomogeneous [Sus89, Vri16]. injection [Jør78]. inner [UG25]. inneren [UG25]. Input [Sim91, BPS15, BCNP15, CS95, HLT16]. inspired [PSLL15]. Institute [WH72]. integrability [AT90]. integral [Sus86, Sus89, Sus93]. Integrals [BOGP12, EM05, KMiS06]. Integrated [Dan91, Fas10, Glo01, HP14, HKST17, MY13]. integration [Fas13, SG15]. Intensity [CW93, CS09]. Interacting [RE13, Gil13, HS81]. Interaction [Eng10, KU35b, PU49]. interactions [PSV18]. Intermittency [GLST16]. Internal [RU35]. International [Ano64, Meh73, Mei65]. Interpretation [EU27a]. Interscience [Bri62, McL62, Uhl60a, Uhl63a]. interval [WC97]. Intervals [FP03]. intrinsic [BS85]. Introduction [Uhl60a]. Invariance [Deu89]. Invariant [FP03]. inventory [KHFA67]. investigating [CS09]. investigations [AF02]. Isolated [KL06, KL08]. isomers [HU38]. Isotropic [vB09]. issue [SSMS17]. Italian [FU33b]. Italy [Meh73]. Iterated [Jia12, Sha93]. Itiro [Uhl63a]. IV [FU57].

J [Bri62, Buf63, Koz66, McL62, Rei63, Uhl60a, Uhl73a]. jacht [vC16]. Japanese [Uhl63a]. jedes [UG25]. Johannes [Uhl24, Uhl24]. John [McL62]. Journal [Rom91]. Jump [BWY11, Nov04, Cha90]. jump-type [Cha90]. Jumps [App06, CS09, Kaw13, Tra17, Wan11a, Wan16, ZWB17]. June [Mei65].

Kac [Bus60]. Kaper [Uhl73a]. Karhunen [Ai16]. Kernel [FH66]. Key [Ano54b]. Killing [Wen90]. kind [AM15b]. Kinetic [CUFF70, GCF66, KL06, KL08, AL97, CU70a, Che02, Sus86, Sus89, Sus93, Uhl62b, UF63a, Uhl74, dBvW78]. Klein [Uhl60a, EU26a]. Known [Sza12, Sza13]. Kolmogorov [BDT11]. Konopinski [Fra05]. Kopplungsmöglichkeiten [GU26a]. Koshiba [Uhl63a]. Kraichnan [Fan04]. Kramers [YU30].

L [Uhl59b]. Lagrangians [PSLL15]. LAN [Tra17]. Landau [Uhl59b]. Laplace [Jac96]. Laporte [Blo44]. Large [BR02, BCS11, BR15, BR17, Bis08a, Deu89, DMRYZ04, FLP99, Hec99, IM89, Wan16, dRdC17, Gas11, Min92a]. Law [Ben38, Jia12, ZZ09, Sha93, UK37]. Least [Lon09, Ma10, ES13, ZZ13]. Lectures [UF60, UF63b, UF86, UF65]. Lee [Ano38a]. Leiden [Uhl80b]. Lekcii [UF65]. levels [VU37]. Lévy [Abd15, AIK18, App15, BN08, BDY07, DY10, DNO18, EM05, FK11, HP14, JVV05, KK11, KMiS06, LZ16, Lon09, Ma10, Mai14, Mas04, MY13, Ped02, PS05, TL09, Vri16, Wan11b, Wan12, ZWB17]. Lévy-driven [Abd15, AIK18, BN08, FK11, JVV05, Mai14, TL09, Vri16, ZWB17]. Li [WU34b]. Lifshitz [Uhl59b]. like [TM08]. likelihood [Bis01, BB01, BY12, Mai14]. Limit [CM97, FMW06, RE13, Yur17, BP07, EH82]. Limitations [Uhl73b]. limits [AL97, Cha90, HS81]. Linear [DM72, FP03, Jia12, KL06, KL08, S.64, FZ99, FK89, FK90, GJ09a, HKST17, JD15, UF62]. linearized [EMV07, FH66]. Liquid [HKU64, KUH63, UHK63]. List [Ano38a, FK89, FK90, Row86]. Local [Kaw13, Shi12, ELY99, LP00, TLZ12]. Localized [PU50]. Loève [Ai16]. Logarithm [Jia12, Sha93]. London [Uhl59b]. Long [FK02, IT99, MY03]. Long-Memory [MY03]. Long-Range [FK02, IT99]. longitudinal [JDD13]. Lorentz [Uhl25]. low [UB37]. Lower [Tho11, ZZ16]. Lyapunov [DG13].

M [Uhl59b]. magnetic [BS85, Ive10]. magnetism [BS85]. make [SK99]. man [Ano54a]. Manifold [CM97, Jør78]. Many [Sul78]. Marc [Bus60]. marginal [TL09]. Markov [CS95, HJM^{+}16, Shi90, TLZ12, ZW17]. Markov-modulated [HJM^{+}16]. Markovian [SK13]. Martin [Uhl60a, Roc92]. Martingale [App06]. Martingale-Valued [App06]. Martingales [Nov04, ELY99]. Mass [Eng10, Gil13]. massive [oT06]. Mathematical [Uhl41, Uhl73a]. matter [BBD^{+}14]. Maxima [Gri68]. maximum [Bis01, BB01, BY12, Mai14]. Maxwell [LU31]. may [Ano54a]. McKean [Jør78]. Mean [Deu89, Tho75, BPvZ03, DNO18]. measured [BS15]. Measures [App06, FP03]. Mechanical [EU27a]. Mechanics [Coh68, Coh71, McL62, Mei65, UF60, Uhl63a, UF63b, Uhl63b, Uhl71b, Uhl80a, UF86, dBU62, dBU64, dBU65, Coh90a, Uhl50, Uhl60c, Uhl60d, Uhl62a, Uhl66, Uhl68, Uhl71a, UF65, Bri62, Buf63, Koz66, Rei63, Uhl40]. mechanik [UF65, Uhl62a]. Medal [Ano64]. meerdimensionale [Uhl25].

Mei [van35]. Memorable [Rom91]. memorial [FW60]. memoriam [Ano89]. Memory [MY03]. Meson [UC52]. mesons [BBD^{+}14]. mesoscopic [SK99]. Method [GCF66, Cic05]. Methoden [Uhl27]. Methods [Uhl27, EU26b, RS12, Uhl60c]. Minima [DU32, Gri68]. Minimal [KLS16, Tay89]. Minimum [DY10, Bis06, KP94]. Miramare [Meh73]. Mischungsparadoxon [EU27b]. Mixed [CK19, Dio16, HKST17, IT99, Shi12]. mixed-effect [Dio16]. Mixing [EU27b]. mixtures [FUF72]. Model [Abd15, CW93, Fan04, FLP99, HM75, KUH63, KR94, ASA89, AT90, Bag11, Dio16, DNO18, HLT16, HKST17, Min92b, NLU40, Pie91, Pie92, SK13, Yan17]. modeling [CS09]. models [Bis08a, BCNP15, CS09, CS95, GS10, HA13, Min92a, RS12, Sch14]. Moderate [GJ09b, GJ09a]. modulated [HJM^{+}16]. moduli [CCS92, CL90, CLS93]. molecular [Ani17]. Molecule [DU32]. Moment [WTZ15, BS85]. Moments [RS88, Vee15]. Monatomic [FU67, CU70b, FUF72]. mono [RU53]. mono-atomic [RU53]. monotonic [BS15]. Monte [Gri68]. Moser [AT90]. Motion [CM97, FMW06, Igl05, Tho26, UG29, UO30, Vak12, WU45b, WU45c, vLU31, BW94, DMLF^{+}17, Jør78, WU45a]. motions [ZZ13]. Multi [GZ10, EU26b, Uhl25]. Multi-Dimensional [GZ10, EU26b, Uhl25]. multidimensional [BPS15, EM08, Mas04, Sim11]. multiparameter [GP11]. multispecies [Del03]. Multitype [GMS93]. Multivariate [BNS11, Fas13].

Named [Ano38a, Ano38b]. Names [Ano38a]. nature [Meh73]. natuurkunde [Uhl55b]. Navier [AL97]. neben [EU27a]. Negative [Pat07, Pav16]. neue [Uhl62a]. neueren [EU27a]. Neumann [AT90]. neuronal [BCNP15]. Neutral [UC52]. Neutron [WU34a]. Newton [And14a, AGGM14]. Next [Ano38b]. Niederer [And14a, And15]. nieuwe [Uhl55b]. Noise [MM05, Sim95, Wil10, CKX09, HP14]. Noises [LZ16, Lon09, Ma10]. Non [DM72, JP99, Kol01, PU50, RPD04, UG32, Uhl41, BG91, DMLF^{+}17, FU70a, GS06, GS10, JD15, Mei65, RS12, Shi12, UB36, UB37]. non-equilibrium [DMLF^{+}17, FU70a, Mei65]. Non-Gaussian [RPD04, BG91, GS06, GS10, RS12]. Non-Hamiltonian [Kol01]. Non-ideal [UG32, UB36, UB37]. Non-Linear [DM72]. Non-Localized [PU50]. Non-Normal [JP99]. non-recurrent [Shi12]. non-stationary [JD15]. Non-uniform [Uhl41]. nonequilibrium [FU70b, FU76]. Nonlinear [CKX09, DMM62, KL06, KL08, Pol90]. Nonlocal [dBvW78]. Nonnegative [BDY07]. Nonparametric [Dio16, JVV05]. nonstationary [BPvZ03]. Nonsymmetric [Sch93]. norm [CL90, CLS93, KP94]. Normal [HM75, JP99, Bis01]. Normality [Kaw13, Shi12]. normings [Bis01]. North [Bri62, Uhl60a, Uhl63a]. North-Holland [Bri62, Uhl60a, Uhl63a]. Note [UK37, Ai16, Bag11, Bel82, Ma10, XXY12, GU25]. Notes [Uhl80a]. notion [RU50]. Nuclear [FU50b, Seg85, HU38, WH72]. nuclei [KU36, VU37]. number [HU53, UG35]. Numbers [Ben38, GF44]. Numerical [PSV18].

O [EU26a]. O. [EU26a]. Obituary [Dre89, MUG^{+}33]. Obligations [Lau56]. observation [CKX09]. Observations [KM12]. observed [Bis06, ES13, Lon09, Ma10, Shi12, ZZ13]. Occupation [Fat09, ZWB17]. odd [Mas16c, Mas16a]. odd-order [Mas16c, Mas16a]. off [Lac05]. oktober [dB89]. Old [Uhl55b, Uhl63a, Uhl62a]. One [KUH63]. One-Dimensional [KUH63]. op [vC16, van35]. Operator [App06, Roc92, BP07, FH66, Mit87, Üst95]. Opmerking [GU25]. Oppenheimer [Lau56]. Optimal [BS15, Igl05, WZ11, BSS13, KS08]. options [ZGO12]. Order [KU35b, Lau56, BMMR15, MD05, Mas16c, Mas16a, PS09, PSLL15]. Oriel [Ano38b]. Origin [PKU72, Jac96]. Ornstein [Ma10, RS12, Sza13, WC97, Abd15, AIK18, AM14, AM15a, Ai16, ÁLBRM16, AHR17, AC87, App06, App15, AKS96, ABI99, APvZ01, AF02, AM15b, Bak82, BH12, BPvZ03, BSS13, BS15, BN01, BNBO11, BOGP12, BB15, Bee75, Ben05, BR02, BCS11, BPS14, BR15, BR17, BW94, BMW^{+}06, BRF17, Bis01, BB01, Bis06, Bis08a, Bis08b, BWY11, BY12, BG91, Bor85, BN08, BD13, BDY07, BPS15, BCNP15, CM97, CF17, CS08, CS09, CGXM96, CGXM97, Cha90, CMN18, CKM03, CK19, CMG02, CW93, CM18, CP07, CCS92, CKS99, CL90, CLS93, Dan91, DMR97, DR98, Deu89, DG13, Dio16, DY10, Dir75, Dit07, DMLF^{+}17, DMRYZ04, DM72, DMM62, DNO18, EK05, ELY99, EH82, EM08, EM05, ES13, FZ99, FK02, Fan04, Fas10, Fas13]. Ornstein [Fat09, Fer90, Fer92, FK11, FLP99, FP03, GJ09a, GJ09b, GO00, Gas11, Gil13, GNdC12, Glo01, GLST16, GP11, GS06, GS10, Gri68, HW97, Hec99, HP14, HA13, HS81, HLT16, HN10, HJM^{+}16, HKST17, IT99, IM89, IMM^{+}90, Jac96, Jeg09, JP99, JM14, Jia12, JD15, JVV05, Jør78, KM11, KM12, Kaw13, KL06, KL08, KS08, KP04, KR94, KP94, Lac05, Lef97, LP00, Li00, LiS09, LS19, Liu08, Liu12, LZ16, LS00, Lon09, MY03, Mai14, MM05, Mas04, MY13, MS09a, Maz17, Mia13, MS92, Mit87, NRS85, Nor98, Nov04, OV04, Pat07, Ped02, PS05, PS09, QH95, Ren87, RE13, RS87, RS88, RD05, RPD04, Roc92, Sal84, Sch90, Sch93, Sch14, SK13, Sha93, Shi90, Shi12, Sim11, Sim91, SWW15, Str58, SG15, Sza12, TM08]. Ornstein [TL09, Tay89, Tho75, Tho11, TLZ12, Tra17, Üst95, Vak12, Vee15, Vri16, Wan11a, Wan11b, Wan12, WX15, WTZ15, WSW15, Wan16, Wen90, Wil10, WP11, Wit84, WZ11, XZW09, Xin12, XXY12, Yur17, ZZ16, ZZ19, ZZ09, Zha11, ZGO12, ZZ13, ZW17, ZWB17, dRdC17]. orthogonality [APvZ01]. Oscillator [MM05, And14a, AGGM14, And14b, And15, ACVV17, BCG13, BMMR15, DMTUV09, KL15, KYH16, MD05, Mas15, Mas16b, Mas16c, Mas16a, Mos10, Pav16, PG13, PSLL15, Smi09, oT06]. oscillators [DS06, KLS16, LP00]. Other [Ano38b]. Oude [Uhl55b]. outline [Uhl68]. output [CS95]. Oxford [Uhl40, Ano38b].

P [van35]. Pairs [RU35]. Pais [Pav16, And14a, AGGM14, And14b, And15, ACVV17, Bag11, BCG13, BMMR15, DS06, DMTUV09, JDD13, KL15, KLS16, KYH16, MD05, Mar81, Mas15, Mas16b, Mas16c, Mas16a, Mos10, PG13, PSLL15, SSMS17, Smi09, oT06]. Pais-Uhlenbeck [Pav16]. Paper [Cer73, Gre50]. papers [Rom91, Uhl60a]. Parabolic [Mas07, Roc92]. Paradox [EU27b]. Parahydrogen [HM75]. Parameter [Glo01, HN10, JD15, AHR17, AM15b, DY10, KP94, ZZ16]. Parameters [Jia12, WTZ15, Abd15, AKS96, BS15, GJ09a, KS08, OV04]. Parseval [LP09]. Particle [AM14, CM97, KL06, KL08, RE13, AM15a, Cha90, PSV18]. Particles [PU49, UG35]. parts [SG15]. Passage [BWY11, NRS85, Nov04, RS88, Tho75, Tho11, Dit07, HP14, Vee15]. Passive [Ben05, Fan04]. path [Sch90]. Paul [Uhl56, MUG^{+}33, Uhl60a]. Pauli [Bel82, FW60]. Peierls [Uhl59b]. Percentages [DF79, MYZ67]. performance [CS09]. Pergamon [Uhl59b]. Periodic [Ped02, PS05, BCNP15]. Personal [Uhl76]. Perspective [Seg85]. Perturbation [Wil10, ACVV17, Cic05]. Phase [AHR17, Dys69, Dys71, EU26a, Uhl78]. Phasenwellen [EU26a]. Phenomena [UU33, CU70c, UH65]. Phenomenon [LG78, Rai85]. Philosophy [Ano38b]. photon [Blo44]. Phylogenetic [JM14]. Physical [Bus60, KU59]. physicist [Meh73]. Physicists [Ano54b, Ano64]. Physics [ESU68, Lau56, Lau58, Meh73, Uhl59b, U^{+}63, Uhl73c, Uhl80a, FW60, KHFA67, Uhl55b, Uhl60b, WH72, Uhl59b, Rom91]. piecewise [DG13]. Pieter [van35]. Pioneers [Ano54a]. place [Lef97]. Planar [Vak12, BW94]. Planck [KL06, KL08]. plasma [PSV18]. plug [ÁLBRM16]. plug-in [ÁLBRM16]. point [CMN18, Kro60]. Poisson [AT90, GZ10, Tra17, Xin12]. Pol [Bus60]. Polyatomic [CUdB64]. Polynomial [DM72, Mit87]. Population [CW93]. Portrait [Pai00]. position [Uhl35]. Positive [DG13, PS09]. Positron [RU35]. positroni [FU33b]. Positrons [FU33a, FU33b]. possibilities [GU26a]. Possible [PKU72]. Post [Ano38a, Ano38b]. potential [Xin12]. pour [Blo44]. Pp [McL62, Uhl59b, Bri62, Uhl60a, Uhl63a]. practical [BRF17]. Prediction [DM72, BSS13]. predictor [ÁLBRM16]. Presentations [Rei63]. Press [Uhl40, Uhl59b]. pressure [RU50]. pricing [ZGO12]. principe [Uhl36]. Principle [Deu89, Uhl40, Uhl36]. Principles [Hec99]. Probability [Bus60, Dir75, KU59, RS87, Uhl35]. Problem [BS98, DU32, UG29, BDT11, CU70a, CMN18, DMTUV09, Min92a, Mor94, NLU40, Pat07, Pol90, Rai76, SU42, Sus86]. Problems [Coh68, GMS93, Uhl73c, FU56a, FNU56, FU56b, FU57, Uhl50, Uhl60b, Uhl74]. procedures [WZ11]. Proceedings [KU59, Mei65, WH72, Kac60]. Process [Bee75, BR02, BCS11, Bis08b, CK19, CW93, CKS99, Dan91, DMR97, DR98, Deu89, Dir75, DM72, DMM62, Gri68, Hec99, Jia12, KL06, KL08, KR94, LS19, LS00, Pat07, Ren87, RE13, RS87, RS88, Sim91, Str58, TM08, Tay89, Tho75, Tho11, Wen90, Wil10, Yur17, dRdC17, Ai16, ÁLBRM16, AF02, AM15b, Bak82, BH12, BB15, BR15, BR17, BW94, Bis01, BB01, Bis06, Bor85, BPS15, BCNP15, CF17, CM18, DY10, Dit07, DMLF^{+}17, DNO18, GJ09a, GJ09b, GO00, Gas11, Gil13, GNdC12, Glo01, IT99, Jac96, JD15, Jør78, KS08, KP94, Lef97, Mas04, MY13, Mia13, Mit87, RD05, SG15, Tra17, Vee15, Wit84, WZ11, XXY12, ZGO12, ZW17]. Processes [App06, BN01, BOGP12, BWY11, BD13, BDY07, CCS92, Eng10, EM05, Fat09, FP03, GLST16, Hec99, IM89, IMM^{+}90, JP99, KM12, Kaw13, Lac05, Li00, Liu12, LZ16, LP09, MY03, MS09a, Nov04, Ped02, PS05, RPD04, Sal84, Sza12, Sza13, Uhl73a, Vak12, Wan12, WX15, WTZ15, WP11, XZW09, ZZ19, ZZ09, Abd15, AIK18, AC87, App15, AKS96, ABI99, BPvZ03, BNS11, BNBO11, BB15, BPS14, BMW^{+}06, BRF17, BY12, BG91, BN08, CS09, Cha90, CKM03, CL90, CLS93, DG13, DY10, DMRYZ04, ELY99, EH82, EM08, ES13, FZ99, Fas10, Fas13, FK11, GP11, GS06, GS10, HP14, HS81, HN10, HJM^{+}16, JVV05, KM11, KMiS06, LiS09, Liu08, Lon09, Ma10, Mai14, Mas04, MS92, Mor94, Nor98, OV04, PS09, RS12, Sch90, Sch93]. processes [Sha93, Shi90, Shi12, Sim11, SWW15, TL09, TLZ12, Vri16, WC97, Wan11a, Wan11b, WSW15, Wan16, Xin12, ZZ16, Zha11, ZZ13, ZWB17]. processus [Bak82, Jac85]. Production [WX15]. Professor [Ano38b, Cer73, Uhl56]. Professors [Ano38a]. Progress [Ano74, Uhl60d]. Propagation [FU67, CU70b, FUF72]. Proper [BOGP12]. Properties [Mas07, S.64, KMiS06, LiS09, Row86, Sch90, UF62]. Property [Kaw13, Liu12, Tra17, WC97]. proportion [Woo80]. Proposal [Fra05]. propose [Ano54a]. Proton [WU34a]. Protons [WU34b]. Provost [Ano38b]. Publications [Bus60, Rei63]. Publishing [Uhl59b]. put [SK13].

Quadratic [Wen90, DG13]. Quanta [Uhl27]. Quantenvektoren [GU26a]. quantization [BMMR15, DS06, MD05]. Quantized [PKU72]. Quantum [PU59, Uhl63a, Uhl71b, And15, AL97, CS08, GU26a, KP04, KHFA67, Mar81, Mos10, UB36, UB37, Uhl71a, Yan17, dBvW78, Uhl63a]. quark [BBD^{+}14]. Quasi [BNBO11, LS00]. Quasi-Stationary [LS00]. Quasiclassics [Kol01]. Quest [Lau56]. Questions [Lau56, Uhl62a, Uhl71a]. Queue [Sim91].

R [Ano38b, Bus60, Uhl59b]. Radial [CW93, dRdC17, ELY99, GJ09b]. Radiation [RU35, KU36]. Radiations [FU50b]. Radioactivity [KU35a, KU41]. Random [BWY11, Bis01, Fer90, Fer92, OV04]. Randomization [CLS93]. Range [FK02, IT99]. Rarefied [Ano74]. rarified [CU70c]. Rate [Igl05, Sim91, WX15, AHR17]. Rate-Optimal [Igl05]. Rates [BB01, Bis06, Bis08b, CS95]. Ratio [EH82]. Ray [SU43, NLU40, SU42]. Rayleigh [CU70a, HM75]. Reading [Uhl59b]. ready [Ano54a]. realization [KLS16]. Receive [Ano64]. Recombination [FU33a, FU33b]. recurrence [DG13, Shi90, WC97]. recurrent [Shi12]. recursion [Vee15]. recursive [BB15]. rede [Uhl64]. Reflected [BWY11, XZW09, ZZ19, BY12, GNdC12, XXY12, ZZ16]. Reflecting [RS87]. reflection [BDT11]. Refugee [Seg85]. Refutation [Fra05]. Region [HKU64]. regression [JM14]. Regularité [Fer92, Fer90]. Regularity [Liu12, LZ16, Fer90, Fer92, MS09b]. regularization [Mar81]. Regularizing [Mas07]. Related [Bus60, KU59]. Relation [Uhl80a, Blo44]. Relations [UL31]. Relativistic [DMR97, DR98, PSV18, RD05]. Remarks [Cic05, DS06, Gre50, Uhl78, Bak82]. Remarques [Bak82]. Reminiscences [Jul78, Uhl56, Uhl76]. reminiscenses [Uhl80b]. Repairman [GMS93]. Replacement [UG25]. report [KHFA67]. Representations [GP11]. requirement [UG25]. Research [Ano54a, Ano54a, Ano54b]. resolution [SK99]. resolving [SSMS17]. Restricted [BCNP15]. result [Dit07]. results [Fas10, OV04]. Retarded [Liu12, Liu08]. reversion [DNO18]. Review [Bri62, Buf63, Koz66, Lau58, McL62, S.64, Uhl40, Uhl41, Uhl59b, Uhl60a, Uhl63a, Uhl73a]. revisited [BW94, HW97, Uhl64, ZUK77]. Richard [Uhl40]. ricombinazione [FU33b]. Riddell [Gre50]. Riemannian [Jør78]. risk [CS09]. Rivière [MS09b]. Rods [vLU31]. Ross [Ano38b]. rotating [GU26b, PU69]. roteerende [GU26b]. Rounded [MYZ67]. Rounding [DF79]. ruimten [Uhl25]. Russian [Uhl59b, UF65].

S [Uhl41]. Sam [vC16, vCH18]. Sample [Lac05, Sch90, Fas10]. Sampling [Kaw13, Glo01, KM11]. Samuel [Sul78]. Sargent [UK37]. says [Ano54a]. Scalar [Fan04]. scheme [Jør78, PSV18]. Science [Ano64, Lau58, Pai00]. Sciences [KU59, WH72, Bus60]. Scientific [Uhl60a]. Scientist [Ano54b, Lau56]. Scientists [Seg85]. second [AM15b, PS09]. seems [Mor07]. Selected [Uhl63b]. selection [CM18]. Self [App06, Eng10]. Self-Decomposability [App06]. Self-Interaction [Eng10]. semidefinite [PS09]. semigroup [CS08]. Semigroups [Mas07, CMG02, HW97, KP04, QH95]. Semilinear [Mas07]. Seminar [KU59, Kac60]. sense [SK99]. September [Meh73]. Sequential [BY12, LS19]. Series [Igl05]. Service [GMS93]. Service-Adaptive [GMS93]. Set [Lau56, BH12, BDT11]. set-indexed [BH12]. sets [BS15]. setting [CMN18]. Sharp [BR02, BCS11]. Shear [Ben05]. sheet [CP07]. sheets [APvZ01, BPvZ03, BSS13, BS15]. shift [BR15]. shifted [BS15]. shock [Pie92]. shock-wave [Pie92]. Short [Coh90b]. Showers [SU43, NLU40, SU42]. sided [Pat07, ZWB17]. signal [CKX09]. Signals [LU50, LU65, BCNP15]. Significant [GF44]. Simple [CW93]. Simulation [Gri68, TL09, Zha11]. Sin [Uhl63a]. Sin-Itiro [Uhl63a]. Singular [EMV08]. size [Mia13]. Skew [WSW15]. skewed [SWW15]. Small [HP14, Jeg09, Kol01, Sim95, FZ99, Gas11, Lon09, Ma10, UG35]. Small-Noise [Sim95]. smooth [BDT11]. Sobolev [Üst95]. Solution [Ani17, Che02, Pol90, YU30, EU26b, EMV07, Mor94]. Solutions [HM75, Min92b, EMV08, Liu08]. Solvability [Sus86, Min92a]. solvable [CM18]. Some [BOGP12, KMiS06, Tho75, Uhl50, Uhl60b, Uhl60d, Uhl62b, Uhl71a, Uhl74, Uhl78, Uhl80a, Uhl80b, Woo80, Fer90, LiS09, OV04, Row86]. Sons [McL62]. Sound [FU67, UU32, CU70b, FUF72]. Sources [KHFA67]. Space [App06, BS98, CKS99, GZ10, Roc92, BDT11, Sch93]. Spaces [FP03, Liu12, LZ16, Mas07, ÁLBRM16, EU26b, Jeg09, Liu08, Uhl25]. Spatial [Eng10, vB09, APvZ01]. Spectra [KU41, UG26, UG67, UGB84, GU25, GU26b]. Spectrally [Pat07].

Spectrum [HM75]. speed [DNO18]. Spherical [WP11]. Spin [Ano54a, Mor07, PU49, Sul78, Sul88, Gou71, Uhl76]. Spinning [Tho26, UG26, UG67, UGB84]. Spinor [LU31]. spins [Ive10]. Spontaneous [WU34a]. Square [Dan91]. Squared [CW93, dRdC17, CL90, CLS93, GJ09b]. squares [DMRYZ04, ES13, Lon09, Ma10, ZZ13]. Stability [HU38, ACVV17, KL15, PG13]. Stable [FMW06, KM12, MY03, Pat07, ZZ09, BRF17, HP14, KM11, Mos10, Zha11, ZZ13]. Staff [Ano38a]. State [UG32, RU53, ZW17]. statement [Uhl25]. states [Bag11, PG13]. stationarity [EM08]. Stationary [Gri68, LS19, Liu08, LS00, Sim91, XZW09, Ai16, BPvZ03, JD15, LiS09, PS09, ZW17]. Statistical [Bri62, Buf63, CK19, Coh68, Coh71, Fas13, GF44, Koz66, McL62, Mei65, Rei63, Uhl27, UG35, Uhl40, UF60, UF63b, U^{+}63, Uhl63b, Uhl71b, Uhl73c, Uhl80a, UF86, dBU62, dBU64, dBU65, AF02, Coh90a, Uhl50, Uhl60c, Uhl60b, Uhl60d, Uhl62a, Uhl66, Uhl68, Uhl71a, UF65, VU37, Uhl59b]. statisticeskoj [UF65]. Statistics [AM14, EU27a, Sus89]. Statistik [EU27a]. Statistiken [EU27a]. statistische [Uhl27]. statistischen [Uhl62a]. stelling [Uhl25]. Stochastic [EK05, RPD04, Wil10, BDT11, DMLF^{+}17, FK11, GS06, GS10, HLT16, RS12]. Stochastically [RE13]. Stokes [AL97]. strangeness [Woo80]. Stretching [CGXM96, CGXM97]. strictly [ELY99]. Strings [vLU31]. structural [BCG13]. Structure [UG26, Uhl66, UG67, UGB84, AT90, Bel82, GU26b, Pie92, PS09]. structuur [GU26b]. Studies [dBU62, dBU64, dBU65, Bri62, Buf63, Koz66, McL62, Rei63]. Successive [FU50b, Uhl60c]. Sulla [FU33b]. Summer [KU59, Kac60]. Sums [MYZ67]. Super [FMW06, Gil13]. Super-Brownian [FMW06]. Superdiffusive [Ben05]. superfluids [PU69]. Superposition [BN01]. Superpositions [GLST16, GS10]. Superprocesses [Kol01]. supersymmetric [Mas15, Mas16a]. supOU [BNS11]. Suprathreshold [Tho11]. sure [Xin12]. Surface [CGXM96, CGXM97]. swaps [CF17]. switching [SK13, ZW17]. Symmetric [CMG02, ZZ13]. Symmetries [ASA89, And14b]. symmetry [AGGM14]. Symposium [Meh73, Mei65, Woo80]. System [AM14, KL06, KL08, RE13, AM15a, Del03, Mos10]. Systematic [FK89, FK90]. Systems [FP03, Cha90, DMLF^{+}17, HS81, SK99].

T [Ano38b, Uhl41]. T3 [PU49]. Tables [GF44]. Tails [BD13, BPS15]. Teacher [Lau56]. Teachers [Ano38a]. Teaching [ESU68]. tears [BR17]. temperature [DNO18, SK99]. temperatures [UB37]. Tempered [BRF17, KM12, ZZ09, KM11, Zha11]. Test [HM75]. Testing [LS19, Bis08a]. tests [AIK18]. their [CMG02, KLS16, PSLL15, WSW15]. theorem [RD05]. theorems [EH82, Sus89, Sus93]. Theoretical [FW60, Meh73, Uhl59b]. Theorie [Uhl27]. Theories [PU50]. Theory [Ano54b, CUFF70, CT73, FUF72, Fra05, GCF66, HKU64, KUH63, KU35a, KU35b, KU41, PU59, S.64, SU43, Uhl27, UO30, Uhl41, Uhl73a, WU45b, WU45c, Wil10, WU34a, ACVV17, CU70a, FU56a, FNU56, FU56b, FU57, FU70b, FU76, HA13, KU37, KU38, Mar81, NLU40, RU53, SU42, UB36, UB37, Uhl62b, UF62, UF63a, UHK63, Uhl74, Uhl78, WU45a, FU70a, Uhl63a]. There [UC52, SK99]. Thermodynamic [PU69]. thermodynamics [FU70a, FU70b, FU76, Mei65]. Third [PU59, LLAT16]. thought [Uhl80c]. Threshold [LU50, LU65]. tick [Mia13]. Ties [Sul78]. Time [CS95, Fat09, FP03, Gri68, KYH16, LZ16, NRS85, RS88, Tho75, Tho11, Uhl80c, Dit07, DNO18, FZ99, Jeg09, LP00, Sch14, TLZ12, Vee15, Vri16]. Time-dependent [CS95, KYH16]. time-inhomogeneous [Vri16]. Times [BWY11, Lac05, Nov04, Sim95, Vak12, BN08, HP14, SWW15, ZWB17]. Tolman [Uhl40]. Tomonaga [Uhl63a]. Topics [Bus60, KU59, Uhl63b]. Topologically [oT06]. Tracer [Ben05]. Trajectory [ZZ19]. transformation [And14a, And15]. Transition [Mas07, ZZ09, AHR17, XXY12]. Transitions [Dys69, Dys71, Uhl78]. Translated [Uhl59b, Uhl63a]. Transport [UU33, Uhl73a, CU70c]. traveling [Min92b]. treatment [PSV18]. tree [AHR17, BB15, HA13]. trees [HU53]. Trend [Str58, Gas11]. Trends [NRS85]. Trieste [Meh73]. Trigonometric [Igl05]. Turbulent [Ben05]. Turn [Coh71]. turning [Kro60]. Twentieth [FW60]. Two [DU32, LS19, Pat07, XZW09, PSV18, ZW17, ZWB17]. Two-Minima [DU32]. two-particle [PSV18]. Two-sided [Pat07, ZWB17]. two-state [ZW17]. Type [BN01, BD13, GLST16, JP99, Li00, WTZ15, BB15, Cha90, CM18, CCS92, HLT16, PS09, Sch14, Shi90, TLZ12].

U.S. [Ano64, ESU68]. Uehling [AL97, BP07, EMV07, EMV08, Min92b, Min92a, Pie91, Pie92, Pol90, PSV18, Sus93, Yan17, dBvW78]. Uhlenbeck [Ano38b, Bri62, Bus60, GO00, Ma10, McL62, S.64, dB89, Abd15, AIK18, AM14, AM15a, Ai16, ÁLBRM16, And14a, AGGM14, And14b, And15, AHR17, Ani17, Ano89, AC87, App06, App15, AKS96, ABI99, APvZ01, AF02, AL97, ASA89, AT90, ACVV17, AM15b, Bag11, BCG13, Bak82, BH12, BS85, BPvZ03, BSS13, BS15, BN01, BNBO11, BOGP12, BB15, Bee75, Ben05, BP07, BR02, BCS11, BPS14, BR15, BR17, BMMR15, BW94, BMW^{+}06, BRF17, BS98, Bis01, BB01, Bis06, Bis08a, Bis08b, BBD^{+}14, Blo44, BWY11, BY12, BG91, Bor85, BN08, BD13, BDY07, BPS15, Buf63, BCNP15, CM97, CF17, CS08, CS09, CGXM96, CGXM97, Cas88, Cer73, Cha90, CMN18, Che02, CKM03, CK19, CMG02, Cic05, CW93, Coh90a, Coh90b]. Uhlenbeck [CM18, CP07, CCS92, CKS99, CL90, CLS93, DS06, Dan91, DMR97, DR98, DMTUV09, Del03, Deu89, DG13, Dio16, DY10, Dir75, Dit07, DMLF^{+}17, DMRYZ04, DM72, DMM62, Dre89, DNO18, EK05, ELY99, EH82, EM08, EM05, ES13, EMV07, EMV08, FZ99, FK02, Fan04, Fas10, Fas13, Fat09, Fer90, Fer92, FK11, FH66, FLP99, FK89, FK90, Fra05, FP03, GJ09a, GJ09b, GCF66, Gas11, Gil13, GNdC12, Glo01, GLST16, GP11, Gre50, GS06, GS10, Gri68, HW97, Hec99, HP14, HA13, HS81, HLT16, HN10, HJM^{+}16, HM75, HKST17, IT99, IM89, IMM^{+}90, Ive10, Jac96, Jac85, Jeg09, JP99, JM14, Jia12, JD15, JDD13, JVV05, Jør78, KL15, KM11, KM12, Kaw13, KL06, KL08, KS08, KP04, Koz66, KLS16, KR94, KP94, KYH16]. Uhlenbeck [Lac05, LLAT16, Lef97, LP00, Li00, LiS09, LS19, Liu08, Liu12, LZ16, LS00, Lon09, MY03, Mai14, MM05, MD05, Mar81, Mas15, Mas16b, Mas16c, Mas16a, Mas04, MY13, MS09a, Maz17, Mia13, MS92, Min92b, Min92a, Mit87, Mor94, Mos10, MS09b, NRS85, Nor98, Nov04, OR01, OV04, Pat07, Pav16, Ped02, PS05, Pie91, Pie92, PS09, Pol90, PSV18, PG13, PSLL15, QH95, RS12, Rei63, Ren87, RE13, RS87, RS88, RD05, RPD04, Roc92, Row86, Sal84, SSMS17, Sch90, Sch93, Sch14, SK99, SK13, Sha93, Shi90, Shi12, Sim11, Sim91, Smi09, SWW15, Sté03, Str58, Sul88, SG15, Sus86, Sus89, Sus93, Sza12, Sza13, TM08, TL09, Tay89, Tho75, Tho11, TLZ12, Tra17, Üst95, Vak12, Vee15, Vri16, WC97, Wan11a, Wan11b]. Uhlenbeck [Wan12, WX15, WTZ15, WSW15, Wan16, Wen90, Wil10, WP11, Wit84, WZ11, XZW09, Xin12, XXY12, Yan17, Yur17, ZZ16, ZZ19, ZZ09, Zha11, ZGO12, ZZ13, ZW17, ZWB17, dBvW78, dRdC17, oT06, vB13]. Ühling [Sus86, Sus89]. uitbreiding [Uhl25]. Unbounded [FP03, FZ99]. uniform [Uhl41]. uniqueness [Sus89]. unitary [Mar81, Mos10]. unitons [ASA89]. Universe [Lau56]. Universities [ESU68]. University [Uhl40]. unknown [CU64, CU70d, Lor36, Uhl58, Uhl59a, WU64]. unmechanical [UG25]. unmechanischen [UG25]. Urbanik [GP11]. using [Bis01, GS10, Yan17]. Utrecht [Ano38b].

Valeurs [Fer92]. Validity [Uhl73b]. valuation [Sch14]. Value [UY30]. Valued [App06, IM89, IMM^{+}90, BG91, CCS92, Sch90]. values [Fer92]. Vapor [HKU64, KUH63, UHK63]. variance [CF17]. Variation [KM12, KM11]. Varying [Sim91, DNO18]. vectors [GU26a]. Velocity [BS98, CGXM96, RE13, UU32, Bor85, CGXM97, Pie91, Pie92]. Veranschaulichung [EU26a]. Verhaltens [UG25]. verhandelingen [van35]. verification [Abd15]. versus [DM72]. via [BW94, CP07, MS09b, Sch93, Vak12]. view [BRF17]. Virial [PU59, S.64, RU53, UF62]. Viscosity [CUdB64]. Visiting [Ano38a, Ano38b]. visits [Uhl80b]. Vol [Koz66, McL62, Bri62, Uhl63a]. Volatility [RPD04, GS06, GS10, RS12, Sch14]. Volume [Buf63, Rei63, Uhl59b, FW60]. vom [UG25]. voor [Uhl25]. Vortices [PKU72].

W [Ano38b, S.64]. Waals [HKU64, KUH63, UHK63, Uhl64, UH65]. wagen [Uhl55b]. Wang [Ani17, BS98, Che02, Del03, FH66, HM75, Mor94, Sté03]. was [Ano54a]. waterstof [GU25]. Wave [EU27a, YU30, EU26b, Min92b, Pie92]. Waves [EU26a]. ways [Uhl55b]. Weak [Bis08b, Bis06]. wellenmechanische [EU27a]. Welt [EU26a]. Wentzel [YU30]. Wesley [Uhl59b]. who [Ano54a]. Wiener [Sza12, CKS99, Roc92, Sza13]. Wiley [Bri62, McL62, Uhl63a]. Winding [Vak12]. windings [BW94]. winner [Ano54a]. without [BR17]. Wolfgang [FW60]. Work [Ano54b, BS85]. workload [BPS15]. World [EU26a, Ano54a]. Wright [Pat07]. x [Bri62]. xvi [Uhl63a].

Year [Ano38b]. years [Uhl76]. yields [Mos10]. York [Bri62, McL62, Uhl60a, Uhl63a].

Zeeman [van35, van35]. Zum [EU27b]. Zwang [UG25].

References

Abdelrazeq:2015:MVL

[Abd15] Ibrahim Abdelrazeq. Model verification for Lévy-driven Ornstein– Uhlenbeck processes with estimated parameters. Statistics & Probability Letters, 104(??):26–35, September 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715215001248.

Arato:1999:FCO

[ABI99] M. Arató, S. Baran, and M. Ispány. Functionals of complex Ornstein–Uhlenbeck processes. Computers and Mathematics and Applications, 37(1):1–13, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0898122198002375.

Antoniadis:1987:EEI

[AC87] Anestis Antoniadis and Rene Carmona. Eigenfunction expansions for infinite dimensional Ornstein–Uhlenbeck processes. Probability Theory and Related Fields, 74(1):31–54, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF01845638.

Avendano-Camacho:2017:PTA

[ACVV17] M. Avendaño-Camacho, J. A. Vallejo, and Yu. Vorobiev. A perturbation theory approach to the stability of the Pais–Uhlenbeck oscillator. Journal of Mathematical Physics, 58(9):093501, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arato:2002:NSI

[AF02] M. Arató and S. Fegyverneki. New statistical investigations of the Ornstein–Uhlenbeck process. Computers and Mathematics and Applications, 44(5–6):677–692, September 2002. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0898122102001827.

Andrzejewski:2014:CNHb

[AGGM14] K. Andrzejewski, A. Galajinsky, J. Gonera, and I. Masterov. Conformal Newton–Hooke symmetry of Pais–Uhlenbeck oscillator. Nuclear Physics B, 885:150–162, 2014. CODEN NUPBBO. ISSN 0550-3213 (print), 1873-1562 (electronic).

Ane:2017:PTC

[AHR17] Cécile Ané, Lam Si Tung Ho, and Sebastien Roch. Phase transition on the convergence rate of parameter estimation under an Ornstein–Uhlenbeck diffusion on a tree. Journal of Mathematical Biology, 74(1–2):355–385, January 2017. CO-DEN JMBLAJ. ISSN 0303-6812 (print), 1432-1416 (electronic).

URL http://link.springer.com/article/10.1007/s00285-
016-1029-x;   http://link.springer.com/content/pdf/10.
1007/s00285-016-1029-x.pdf.

Ai:2016:NKL

[Ai16] Xiaohui Ai. A note on Karhunen–Loève expansions for the demeaned stationary Ornstein–Uhlenbeck process. Statistics & Probability Letters, 117(??):113–117, October 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715215300419.

Abdelrazeq:2018:GFT

[AIK18] Ibrahim Abdelrazeq, B. Gail Ivanoff, and Rafal Kulik. Goodnessof-fit tests for Lévy-driven Ornstein–Uhlenbeck processes. Canadian Journal of Statistics = Revue canadienne de statistique, 46 (2):355–376, June 2018. CODEN ???? ISSN 0319-5724 (print),

1708-945X (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1002/cjs.11352.

Arato:1996:EDE

[AKS96] M. Arató, A. Kuki, and A. Szabó. Exact distribution of estimators of parameters in Ornstein–Uhlenbeck processes. Computers and Mathematics and Applications, 31(11):45–54, June 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0898122196000600.

Arlotti:1997:ENSb

[AL97] Luisa Arlotti and Miroslaw Lachowicz. Euler and Navier–Stokes limits of the Uehling–Uhlenbeck quantum kinetic equations. Journal of Mathematical Physics, 38(7):3571–3588, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alvarez-Liebana:2016:CPF

[ÁLBRM16] Javier Álvarez-Liébana, Denis Bosq, and María D. Ruiz-Medina. Consistency of the plug-in functional predictor of the Ornstein– Uhlenbeck process in Hilbert and Banach spaces. Statistics & Probability Letters, 117(??):12–22, October 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771521630044X.

Adamczak:2014:SOU

[AM14] Radoslaw Adamczak and Piotr Miloś. U-statistics of Ornstein– Uhlenbeck branching particle system. Journal of Theoretical Probability, 27(4):1071–1111, December 2014. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10959-013-0503-
2.pdf.

Adamczak:2015:COU

[AM15a] Radosław Adamczak and Piotr Miłoś. CLT for Ornstein– Uhlenbeck branching particle system. Electronic Journal of Probability, 20(??):42:1–42:35, ???? 2015. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/4233.

Azmoodeh:2015:DPE

[AM15b] Ehsan Azmoodeh and José Igor Morlanes. Drift parameter estimation for fractional Ornstein–Uhlenbeck process of the second kind. Statistics: a Journal of Theoretical and Applied Statistics, 49(1):1–18, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Andrzejewski:2014:CNHa

[And14a] Krzysztof Andrzejewski. Conformal Newton–Hooke algebras, Niederer’s transformation and Pais–Uhlenbeck oscillator. Physics Letters B, 738:405–411, 2014. CODEN PYLBAJ. ISSN 0370-2693 (print), 1873-2445 (electronic).

Andrzejewski:2014:HFS

[And14b] Krzysztof Andrzejewski. Hamiltonian formalisms and symmetries of the Pais–Uhlenbeck oscillator. Nuclear Physics B, 889:333–350, 2014. CODEN NUPBBO. ISSN 0550-3213 (print), 1873-1562 (electronic).

Andrzejewski:2015:GNT

[And15] Krzysztof Andrzejewski. Generalized Niederer’s transformation for quantum Pais–Uhlenbeck oscillator. Nuclear Physics B, 901: 216–228, 2015. CODEN NUPBBO. ISSN 0550-3213 (print), 1873-1562 (electronic).

Anikin:2017:SWC

[Ani17] Yu. A. Anikin. Solution of the Wang Chang–Uhlenbeck equation for molecular hydrogen. Computational Mathematics and Mathematical Physics, 57(6):1048–1065, 2017. ISSN 0965-5425 (print), 1555-6662 (electronic).

Anonymous:1938:CNP

[Ano38a] Anonymous. Columbia names 55 as professors: H. D. Gideonse gets economics post at Barnard —- E. A. Lee on teachers college staff. 2 named as associates. 27 assistants, 12 clinical and 12 visiting professors complete list. New York Times, ??(??):20, October 20, 1938. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/102387271/.

Anonymous:1938:DWD

[Ano38b] Anonymous. Dr. W. D. Ross named to Columbia post: Oriel College Provost, Oxford, to be Visiting Professor of Philosophy next year 8. other appointments: Dr. G. E. Uhlenbeck of Utrecht, Holland and Dr. R. T. Ely added to faculty. New York Times, ??(??):47, May 15, 1938. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL https://search.proquest.

com/hnpnewyorktimes/docview/102557946/.

Anonymous:1954:RES

[Ano54a] Anonymous. Research: Electron spin pioneers honored: “discoverer” of a concept may be the man who happened to propose the idea when the world was ready to accept it, says Research Corp. award winner Goudsmit. Chemical & Engineering News, 32(6):

480–486, February 1954. CODEN CENEAR. ISSN 0009-2347 (print), 1520-605X (electronic).

Anonymous:1954:SWH

[Ano54b] Anonymous. Scientists’ work of 1925 honored: Research Corporation gives award to 2 Dutch physicists for key to atomic theory. New York Times, ??(??):13, January 23, 1954. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/112966315/.

Anonymous:1964:UPR

[Ano64] Anonymous. 2 U.S. physicists to receive international science medal. New York Times, ??(??):12, June 17, 1964. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/115824554/.

Anonymous:1974:RGD

[Ano74] Anonymous, editor. Rarefied Gas Dynamics [Progress in Astronautics and Aeronautics], volume 12. Academic Press, New York, USA, 1974. CODEN PRGDAJ. ISSN 0539-0613.

Anonymous:1989:MGU

[Ano89] Anonymous. In memoriam G. E. Uhlenbeck (1900–1988). Nederlands Tijdschrift voor Natuurkunde, 55(2):??, January 1989. CO-DEN NTINEL. ISSN 0926-4264. URL https://www.ntvn.nl/

archief/pdf/431/10/.

Applebaum:2006:MVM

[App06] David Applebaum. Martingale-valued measures, Ornstein– Uhlenbeck processes with jumps and operator self-decomposability in Hilbert space. Lecture Notes in Mathematics, 1874:171–196, 2006. CODEN LNMAA2. ISBN 3-540-30994-2 (print), 3-540-35513-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-
3-540-35513-7_14.pdf.

Applebaum:2015:IDO

[App15] David Applebaum. Infinite dimensional Ornstein–Uhlenbeck processes driven by Lévy processes. Probability Surveys, 12(??):33– 54, ???? 2015. CODEN ???? ISSN 1549-5787. URL http://

projecteuclid.org/euclid.ps/1440075825.

Arato:2001:AIS

[APvZ01] M. Arató, G. Pap, and M. C. A. van Zuijlen. Asymptotic inference for spatial autoregression and orthogonality of Ornstein–Uhlenbeck sheets. Computers and Mathematics and Applications, 42(1–2):219–229, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0898122101001468.

Arsenault:1989:SEC

[ASA89] Guy Arsenault and Yvan Saint-Aubin. Symmetries of the Euclidean U(3) chiral model and Uhlenbeck’s unitons. In XVIIth International Colloquium on Group Theoretical Methods in Physics (Sainte-Adèle, PQ, 1988), pages 468–471. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1989.

Avan:1990:PSI

[AT90] J. Avan and M. Talon. Poisson structure and integrability of the Neumann–Moser–Uhlenbeck model. International Journal of Modern Physics A. Particles and Fields. Gravitation. Cosmology, 5(23):4477–4488, 1990. ISSN 0217-751X.

Bagarello:2011:NPU

[Bag11] F. Bagarello. A note on the Pais–Uhlenbeck model and its coherent states. International Journal of Theoretical Physics, 50(10):3241– 3250, 2011. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic).

Bakry:1982:RPO

[Bak82] D. Bakry. Remarques sur le processus d’Ornstein–Uhlenbeck en dimension infinie. (French) [Remarks on the Ornstein–Uhlenbeck process in infinite dimension]. Lecture Notes in Mathematics, 920: 134–137, 1982. CODEN LNMAA2. ISBN 3-540-11485-8 (print), 3-540-39158-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0092774/.

Bishwal:2001:RCA

[BB01] J. P. N. Bishwal and A. Bose. Rates of convergence of approximate maximum likelihood estimators in the Ornstein–Uhlenbeck process. Computers and Mathematics and Applications, 42(1–2): 23–38, July 2001. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0898122101001274.

Baur:2015:FPI

[BB15] Erich Baur and Jean Bertoin. The fragmentation process of an infinite recursive tree and Ornstein–Uhlenbeck type processes. Electronic Journal of Probability, 20(??):98:1–98:20, ???? 2015. CO-DEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/3866.

Blaschke:2014:GBU

[BBD^{+}14] D. Blaschke, M. Buballa, A. Dubinin, G. Röpke, and D. Zablocki. Generalized Beth–Uhlenbeck approach to mesons and diquarks in hot, dense quark matter. Annals of Physics, 348:228–255, 2014. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035x (electronic).

Bagchi:2013:CSF

[BCG13] B. Bagchi, A. Ghose Choudhury, and Partha Guha. Comments on the structural features of the Pais–Uhlenbeck oscillator. Modern Phys. Lett. A, 28(14):1375001, 9, 2013. ISSN 0217-7323 (print), 1793-6632 (electronic).

Buonocore:2015:ROU

[BCNP15] A. Buonocore, L. Caputo, A. G. Nobile, and E. Pirozzi. Restricted Ornstein–Uhlenbeck process and applications in neuronal models with periodic input signals. Journal of Computational and Applied Mathematics, 285(??):59–71, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042715000643.

Bercu:2011:SLD

[BCS11] B. Bercu, L. Coutin, and N. Savy. Sharp large deviations for the fractional Ornstein–Uhlenbeck process. Theory of Probability and its Applications, 55(4):575–610, ???? 2011. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://

epubs.siam.org/tvp/resource/1/tprbau/v55/i4/p575_s1.

Borovkov:2013:OUT

[BD13] K. Borovkov and G. Decrouez. Ornstein–Uhlenbeck type processes with heavy distribution tails. Theory of Probability and its Applications, 57(3):396–418, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Barbu:2011:KEA

[BDT11] Viorel Barbu, Giuseppe Da Prato, and Luciano Tubaro. Kolmogorov equation associated to the stochastic reflection problem on a smooth convex set of a Hilbert space II. Annales de l’Institut Henri Poincaré, Probabilités et Statistiques, 47(3):699– 724, August 2011. CODEN AHPBAR. ISSN 0246-0203 (print),

1778-7017 (electronic).  URL http://projecteuclid.org/
euclid.aihp/1308834856;   http://www.numdam.org/item?
id=AIHPB_2011__47_3_699_0.

Brockwell:2007:ENL

[BDY07] Peter J. Brockwell, Richard A. Davis, and Yu Yang. Estimation for nonnegative Lévy–driven Ornstein–Uhlenbeck processes. Journal of Applied Probability, 44(4):977–989, December 2007. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595901.

Beekman:1975:ADOb

[Bee75] John A. Beekman. Asymptotic distributions for the Ornstein– Uhlenbeck process. Journal of Applied Probability, 12(1):107– 114, March 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212412.

Belloni:1982:PNH

[Bel82] Lanfranco Belloni. Pauli’s 1924 note on hyperfine structure. American Journal of Physics, 50(5):461–464, May 1982. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Benford:1938:LAN

[Ben38] Frank Benford. The law of anomalous numbers. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 78(4):551–572, March 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL http://links.jstor.org/sici?sici=0003-049X%
2819380331%2978:4%3C551:TLOAN%3E2.0.CO%3B2-G. See com-

ments about Benford’s biased rounding practices [DF79].

Benabou:2005:SBP

[Ben05] Gaël Benabou. Superdiffusive behaviour of a passive Ornstein– Uhlenbeck tracer in a turbulent shear flow. Journal of Statistical Physics, 121(3–4):319–341, November 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-005-7003-4.

Bojdecki:1991:GNG

[BG91] Tomasz Bojdecki and Luis G. Gorostiza. Gaussian and non-Gaussian distribution-valued Ornstein–Uhlenbeck processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(??):1136–1149, ???? 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Balanca:2012:SIO

[BH12] Paul Balança and Erick Herbin. A set-indexed Ornstein– Uhlenbeck process. Electronic Communications in Probability, 17: 39:1–39:14, 2012. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1903.

Bishwal:2001:ANA

[Bis01] J. P. N. Bishwal. Accuracy of normal approximation for the maximum likelihood estimator and Bayes estimators in the Ornstein–Uhlenbeck process using random normings. Statistics & Probability Letters, 52(4):427–439, May 1, 2001. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715201000268.

Bishwal:2006:RWC

[Bis06] Jaya P. N. Bishwal. Rates of weak convergence of approximate minimum contrast estimators for the discretely observed Ornstein–Uhlenbeck process. Statistics & Probability Letters, 76(13):1397–1409, July 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715206000472.

Bishwal:2008:LDT

[Bis08a] Jaya P. N. Bishwal. Large deviations in testing fractional Ornstein–Uhlenbeck models. Statistics & Probability Letters, 78(8):953–962, June 1, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207003471.

Bishwal:2008:RWC

[Bis08b] Jaya P. N. Bishwal. Rates of weak convergence of estimators in the Ornstein–Uhlenbeck process. Lecture Notes in Mathematics, 1923:159–200, 2008. CODEN LNMAA2. ISBN 3-540-74447-9 (print), 3-540-74448-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-540-74448-1_8.pdf.

Bloch:1944:RUL

[Blo44] Léon Bloch. Sur la relation de Uhlenbeck et Laporte pour le photon. Comptes rendus de l’Académie des sciences, Paris, 219:674– 675, 1944.

Berra-Montiel:2015:DQP

[BMMR15] Jasel Berra-Montiel, Alberto Molgado, and Efra´ ın Rojas. Deformation quantization of the Pais–Uhlenbeck fourth order oscillator. Annals of Physics, 362:298–310, 2015. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035x (electronic).

Bezuglyy:2006:GOU

[BMW^{+}06] V. Bezuglyy, B. Mehlig, M. Wilkinson, K. Nakamura, and E. Arvedson. Generalized Ornstein–Uhlenbeck processes. Journal of Mathematical Physics, 47(7):073301, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i7/
p073301_s1.

Barndorff-Nielsen:2001:SOU

[BN01] O. E. Barndorff-Nielsen. Superposition of Ornstein–Uhlenbeck type processes. Theory of Probability and its Applications, 45(2): 175–194, June 2001. CODEN TPRBAU. ISSN 0040-585X (print),

1095-7219 (electronic). URL http://epubs.siam.org/sam-bin/
dbq/article/97816.

Borovkov:2008:ETL

[BN08] Konstantin Borovkov and Alexander Novikov. On exit times of Lévy-driven Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 78(12):1517–1525, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715208000060.

Barndorff-Nielsen:2011:QOU

[BNBO11] Ole E. Barndorff-Nielsen and Andreas Basse-O’Connor. Quasi Ornstein–Uhlenbeck processes. Bernoulli: official journal of the Bernoulli Society for Mathematical Statistics and Probability, 17 (3):916–941, August 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1310042850.

Barndorff-Nielsen:2011:MSP

[BNS11] Ole Eiler Barndorff-Nielsen and Robert Stelzer. Multivariate supOU processes. Annals of Applied Probability, 21(1):140–182, February 2011. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1292598030.

Basse-OConnor:2012:SCP

[BOGP12] Andreas Basse-O’Connor, Svend-Erik Graversen, and Jan Pedersen. Some classes of proper integrals and generalized Ornstein– Uhlenbeck processes. Lecture Notes in Mathematics, 2046:61– 74, 2012. CODEN LNMAA2. ISBN 3-642-27460-9 (print), 3-642-27461-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-27461-9_3/.

Borell:1985:GBO

[Bor85] Christer Borell. Geometric bounds on the Ornstein–Uhlenbeck velocity process. Zeitschrift für Wahrscheinlichkeitstheorie und verwandte Gebiete, 70(1):1–13, ???? 1985. CODEN ZWVGAA.

ISSN 0044-3719. URL http://link.springer.com/article/
10.1007/BF00532234.

Benedetto:2007:CLU

[BP07] D. Benedetto and M. Pulvirenti. The classical limit for the Uehling–Uhlenbeck operator. Bull. Inst. Math. Acad. Sin. (N.S.), 2(4):907–920, 2007. ISSN 2304-7909.

Bercu:2014:OUD

[BPS14] Bernard Bercu, Frédéric Proïa, and Nicolas Savy. On Ornstein– Uhlenbeck driven by Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 85(??):36–44, February 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715213003635.

Budhiraja:2015:ACM

[BPS15] Amarjit Budhiraja, Vladas Pipiras, and Xiaoming Song. Admission control for multidimensional workload input with heavy tails and fractional Ornstein–Uhlenbeck process. Advances in Applied Probability, 47(2):476–505, June 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1435236984.

Baran:2003:EMS

[BPvZ03] S. Baran, G. Pap, and M. C. A. van Zuijlen. Estimation of the mean of stationary and nonstationary Ornstein–Uhlenbeck processes and sheets. Computers and Mathematics and Applications, 45(4–5):563–579, February/March 2003. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0898122103000178.

Bercu:2002:SLD

[BR02] B. Bercu and A. Rouault. Sharp large deviations for the Ornstein– Uhlenbeck process. Theory of Probability and its Applications, 46(1):1–19, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http://epubs.siam.org/

sam-bin/dbq/article/97873.

Bercu:2015:LDO

[BR15] Bernard Bercu and Adrien Richou. Large deviations for the Ornstein–Uhlenbeck process with shift. Advances in Applied Probability, 47(3):880–901, September 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1444308886.

Bercu:2017:LDO

[BR17] Bernard Bercu and Adrien Richou. Large deviations for the Ornstein–Uhlenbeck process without tears. Statistics & Probability Letters, 123(??):45–55, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715216302838.

Bianchi:2017:TSO

[BRF17] Michele Leonardo Bianchi, Svetlozar T. Rachev, and Frank J. Fabozzi. Tempered stable Ornstein–Uhlenbeck processes: a practical view. Communications in Statistics: Simulation and Computation, 46(1):423–445, 2017. CODEN CSSCDB. ISSN 0361-0918.

Brittin:1962:BRS

[Bri62] W. E. Brittin. Book review: Studies in Statistical Mechanics. vol. 1. J. De Boer and G. E. Uhlenbeck, Eds. North-Holland, Amsterdam; Interscience (Wiley), New York, 1962. x + 350 pp. $13.75. Science, 138(3544):965, November 1962. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Baracca:1985:BCI

[BS85] Angelo Baracca and Edvige Schettino. Birth of the concept of intrinsic magnetic moment of electrons in chemistry and magnetism before the work of Uhlenbeck and Goudsmit (1915–1921). Rendiconti della Accademia Nazionale delle Scienze detta dei XL. Memorie di Scienze Fisiche e Naturali. Serie V. Parte II, 9:263–267, 1985. ISSN 0392-4130. Proceedings of the fifth national congress on the history of physics (Italian) (Rome, 1984).

Bicout:1998:WUP

[BS98] D. J. Bicout and Attila Szabo. On the Wang–Uhlenbeck problem in discrete velocity space. Journal of Statistical Physics, 91 (5–6):1047–1054, June 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/A:1023088118307.

Baran:2015:ODP

[BS15] S. Baran and M. Stehlík. Optimal designs for parameters of shifted Ornstein–Uhlenbeck sheets measured on monotonic sets. Statistics & Probability Letters, 99(??):114–124, April 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715215000127.

Baran:2013:ODP

[BSS13] Sándor Baran, Kinga Sikolya, and Milan Stehlík. On the optimal designs for the prediction of Ornstein–Uhlenbeck sheets. Statistics & Probability Letters, 83(6):1580–1587, June 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771521300076X.

Buff:1963:BRS

[Buf63] Frank P. Buff. Book review: Studies in Statistical Mechanics, Volume I by J. De Boer and G. E. Uhlenbeck. American Scientist, 51(1):102A–103A, March 1963. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL https://www.jstor.

org/stable/27838664.

Bush:1960:RPP

[Bus60] Kenneth A. Bush. Recent publications: Probability and Related Topics in Physical Sciences, by Marc Kac, G. E. Uhlenbeck, A. R. Hibbs, and B. van der Pol. American Mathematical Monthly, 67(6):606–607, June/July 1960. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.

org/stable/2309209.

Bertoin:1994:AWP

[BW94] Jean Bertoin and Wendelin Werner. Asymptotic windings of planar Brownian motion revisited via the Ornstein–Uhlenbeck process. Lecture Notes in Mathematics, 1583:138–152, 1994. CODEN LNMAA2. ISBN 3-540-58331-9 (print), 3-540-48656-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073842/.

Bo:2011:FPT

[BWY11] Lijun Bo, Yongjin Wang, and Xuewei Yang. First passage times of (reflected) Ornstein–Uhlenbeck processes over random jump boundaries. Journal of Applied Probability, 48(3):723–732, September 2011. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
23065932.

Bo:2012:SML

[BY12] Lijun Bo and Xuewei Yang. Sequential maximum likelihood estimation for reflected generalized Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 82(7):1374–1382, July 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715212001113.

Casimir:1988:GEU

[Cas88] H. B. G. Casimir. George Eugene Uhlenbeck (1900–1988). Nature, 336(6201):717, December 29, 1988. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Csaki:1992:FTI

[CCS92] Endre Csáki, Miklós Csörgő, and Qi-Man Shao. Fernique type inequalities and moduli of continuity for l^{2}-valued Ornstein– Uhlenbeck processes. Annales de l’Institut Henri Poincaré, Probabilités et Statistiques, 28(4):479–517, ???? 1992. CODEN AHP-BAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//www.numdam.org/item?id=AIHPB_1992__28_4_479_0.

Cercignani:1973:CPU

[Cer73] C. Cercignani. Comment to Professor Uhlenbeck’s paper. In Cohen and Thirring [CT73], pages 121–122. ISBN 3-7091-8338-3 (print), 3-7091-8336-7 (e-book). ISSN 0177-8811. LCCN

QC173.96-174.52. URL https://link.springer.com/book/10.
1007/978-3-7091-8336-6. See [Uhl73b].

Cao:2017:AAV

[CF17] Jian-Peng Cao and Yan-Bing Fang. An analytical approach for variance swaps with an Ornstein–Uhlenbeck process. The ANZIAM Journal, 59(1):83–102, July 2017. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic). URL https://

www.cambridge.org/core/journals/anziam-journal/article/
an-analytical-approach-for-variance-swaps-with-an-ornsteinuhlenbeck-
process/EE61B8C62B4240CD4A91B1E1DF29F2EF.

Carmona:1996:SSO

[CGXM96] Rene Carmona, Stanislav Grishin, Lin Xu, and Stanislav Molchanov. Surface stretching for Ornstein–Uhlenbeck velocity fields. Electronic Communications in Probability, 2:1:1–1:11, 1996. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/980.

Carmona:1997:SSO

[CGXM97] Rene A. Carmona, Stanislav Grishin, Lin Xu, and Stanislav Molchanov. Surface stretching for Ornstein Uhlenbeck velocity fields. Electronic Communications in Probability, 2:1:1–1:11, 1997. CODEN ???? ISSN 1083-589X.

Chari:1990:GJT

[Cha90] Ravi Chari. Generalized jump-type Ornstein–Uhlenbeck processes as limits of infinite-particle systems. Journal of Theoretical Probability, 3(1):9–29, January 1990. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.

springer.com/article/10.1007/BF01063326.

Cheremisin:2002:SWC

[Che02] F. G. Cheremisin. Solution of the Wang Chang–Uhlenbeck kinetic equation. Doklady Akademii nauk SSSR, 387(4):487–490, 2002. CODEN DANKAS. ISSN 0869-5652.

Cicortas:2005:RUP

[Cic05] Graţiela Cicortaş. Remarks on Uhlenbeck’s perturbation method. Mathematica, 47(70)(2):137–143, 2005. ISSN 1222-9016.

Chigansky:2019:SAM

[CK19] P. Chigansky and M. Kleptsyna. Statistical analysis of the mixed fractional Ornstein–Uhlenbeck process. Theory of Probability and its Applications, 63(3):408–425, ???? 2019. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cheridito:2003:FOU

[CKM03] Patrick Cheridito, Hideyuki Kawaguchi, and Makoto Maejima. Fractional Ornstein–Uhlenbeck processes. Electronic Journal of Probability, 8:3:1–3:14, 2003. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/125.

Csaki:1999:CEB

[CKS99] Endre Csáki, Davar Khoshnevisan, and Zhan Shi. Capacity estimates, boundary crossings and the Ornstein–Uhlenbeck process in Wiener space. Electronic Communications in Probability, 4: 13:103–13:109, 1999. CODEN ???? ISSN 1083-589X. URL http:

//ecp.ejpecp.org/article/view/1011.

Crisan:2009:NFS

[CKX09] Dan Crisan, Michael Kouritzin, and Jie Xiong. Nonlinear filtering with signal dependent observation noise. Electronic Journal of Probability, 14:63:1863–63:1883, 2009. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/687.

Csorgo:1990:MCG

[CL90] Miklós Csörgö and Zheng Yan Lin. On moduli of continuity for Gaussian and l^{2}-norm squared processes generated by Ornstein– Uhlenbeck processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 42(??):141–158, ???? 1990. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Csorgo:1993:RMC

[CLS93] M. Csörgö, Z.-Y. Lin, and Q.-M. Shao. Randomization moduli of continuity for l^{2}-norm squared Ornstein–Uhlenbeck processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):269–283, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Calzolari:1997:LMO

[CM97] Antonella Calzolari and Federico Marchetti. Limit motion of an Ornstein–Uhlenbeck particle on the equilibrium manifold of a force field. Journal of Applied Probability, 34(4):924–938, December 1997. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3215007.

Cortines:2018:GES

[CM18] Aser Cortines and Bastien Mallein. The genealogy of an exactly solvable Ornstein–Uhlenbeck type branching process with selection. Electronic Communications in Probability, 23(??):98:1–98:13, ???? 2018. CODEN ???? ISSN 1083-589X. URL https://

projecteuclid.org/euclid.ecp/1545102494.

Chojnowska-Michalik:2002:SOU

[CMG02] Anna Chojnowska-Michalik and Beniamin Goldys. Symmetric Ornstein–Uhlenbeck semigroups and their generators. Probability Theory and Related Fields, 124(4):459–486, December 2002. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400200222.

Chen:2018:ICP

[CMN18] Fuqi Chen, Rogemar Mamon, and Sévérien Nkurunziza. Inference for a change-point problem under a generalised Ornstein– Uhlenbeck setting. Annals of the Institute of Statistical Mathematics (Tokyo), 70(4):807–853, August 2018. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://

link.springer.com/article/10.1007/s10463-017-0610-4.

Cohen:1968:FPS

[Coh68] E. G. D. Cohen, editor. Fundamental Problems in Statistical Mechanics. North-Holland Publishing Co., Amsterdam, The Netherlands, 1968.

Cohen:1971:SMT

[Coh71] E. G. D. Cohen, editor. Statistical Mechanics at the Turn of the Decade. Marcel Dekker, Inc., New York, NY, USA, 1971. ISBN 0-8247-1111-4. LCCN QC175 .S7.

Cohen:1990:GUS

[Coh90a] E. G. D. Cohen. George E. Uhlenbeck and statistical mechanics. American Journal of Physics, 58(7):619–625, 1990. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Cohen:1990:SBG

[Coh90b] E. G. D. Cohen. Short biography of G. E. Uhlenbeck. American Journal of Physics, 58(7):618, July 1990. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Coutin:2007:AFB

[CP07] Laure Coutin and Monique Pontier. Approximation of the fractional Brownian sheet via Ornstein–Uhlenbeck sheet. ESAIM: Probability and Statistics, 11(??):115–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Chen:1995:TDB

[CS95] Thomas M. Chen and Vijay K. Samalam. Time-dependent behavior of fluid buffer models with Markov input and constant output rates. SIAM Journal on Applied Mathematics, 55(3):784–799, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-

712X (electronic). URL http://epubs.siam.org/sam-bin/dbq/
article/24010.

Carbone:2008:HQO

[CS08] Raffaella Carbone and Emanuela Sasso. Hypercontractivity for a quantum Ornstein–Uhlenbeck semigroup. Probability Theory and Related Fields, 140(3–4):505–522, March 2008. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:

//link.springer.com/article/10.1007/s00440-007-0073-2.

Cariboni:2009:JIM

[CS09] Jessica Cariboni and Wim Schoutens. Jumps in intensity models: investigating the performance of Ornstein–Uhlenbeck processes in credit risk modeling. Metrika. International Journal for Theoretical and Applied Statistics, 69(2–3):173–198, March 2009.

CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00184-008-0213-4.

Cohen:1973:BET

[CT73] E. G. D. Cohen and W. Thirring, editors. The Boltzmann Equation: Theory and Applications, volume 10/1973 of Acta Physica Austriaca, Supplementum X Proceedings of the International Symposium “100 Years Boltzmann Equation” in Vienna 4th–8th September 1972. Springer Vienna, Vienna, Austria, 1973. ISBN 3-7091-8338-3 (print), 3-7091-8336-7 (e-book). ISSN 0177-8811.

LCCN QC173.96-174.52. URL https://link.springer.com/
book/10.1007/978-3-7091-8336-6.

Chang:1964:Ua

[CU64] W. Chang and G. E. Uhlenbeck. [unknown]. In The Kinetic Theory of Gases [CUFF70], pages 17–?? ISBN 0-444-10008-3. LCCN QC175 S933 v.5 1970.

Chang:1970:KTGa

[CU70a] C. S. Wang Chang and G. E. Uhlenbeck. The kinetic theory of a gas in alternating outside force fields: A generalization of the Rayleigh problem. In The Kinetic Theory of Gases [CUFF70], pages 76–98. ISBN 0-444-10008-3. LCCN QC175 S933 v.5 1970.

Chang:1970:PSM

[CU70b] C. S. Wang Chang and G. E. Uhlenbeck. On the propagation of sound in monatomic gases. In The Kinetic Theory of Gases [CUFF70], pages 43–68. ISBN 0-444-10008-3. LCCN QC175 S933 v.5 1970.

Chang:1970:TPR

[CU70c] C. S. Wang Chang and G. E. Uhlenbeck. On the transport phenomena in rarified gases. In The Kinetic Theory of Gases [CUFF70], pages 116–?? ISBN 0-444-10008-3. LCCN QC175 S933 v.5 1970.

Chang:1970:Ub

[CU70d] W. Chang and G. E. Uhlenbeck. [unknown]. In The Kinetic Theory of Gases [CUFF70], pages 27–?? ISBN 0-444-10008-3. LCCN QC175 S933 v.5 1970.

Chang:1964:HCV

[CUdB64] C. S. Wang Chang, G. E. Uhlenbeck, and J. de Boer. The heat conductivity and viscosity of polyatomic gases. In Studies in Statistical Mechanics, volume 2, pages 241–268. North-Holland Publishing Co., Amsterdam, The Netherlands, 1964.

Chang:1970:KTGb

[CUFF70] C. S. Wang Chang, G. E. Uhlenbeck, J. Foch, and G. W. Ford. The Kinetic Theory of Gases, volume 5 of Studies in statistical mechanics. North-Holland Publishing Co., Amsterdam, The Netherlands, 1970. ISBN 0-444-10008-3. xiv + 239 pp. LCCN QC175 S933 v.5 1970.

Clifford:1993:ECP

[CW93] Peter Clifford and Gang Wei. The equivalence of the Cox process with squared radial Ornstein–Uhlenbeck intensity and the death process in a simple population model. Annals of Applied Probability, 3(3):863–873, August 1993. CODEN ???? ISSN 1050-5164 (print), 2168-8737 (electronic). URL http://projecteuclid.

org/euclid.aoap/1177005368.

Dankel:1991:DIS

[Dan91] Thad Dankel, Jr. On the distribution of the integrated square of the Ornstein–Uhlenbeck process. SIAM Journal on Applied Mathematics, 51(2):568–574, April 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

deBoer:1989:GEU

[dB89] J. de Boer. George Eugène Uhlenbeck: 6 december 1900– 31 oktober 1988. Jaarboek. Proceedings of the Royal Netherlands Academy of Arts and Sciences (KNAW), ??(??):182–

190, ???? 1989. URL http://www.dwc.knaw.nl/biografie/
pmknaw/?pagetype=authorDetail&aId=PE00003410;   http:/
/www.dwc.knaw.nl/DL/levensberichten/PE00003410.pdf.

deBoer:1962:SSM

[dBU62] Jan Hendrik de Boer and George Eugene Uhlenbeck. Studies in Statistical Mechanics, volume I of Series in Physics. North-Holland Publishing Co., Amsterdam, The Netherlands, 1962. ISBN 0-7204-1105-X. x + 350 pp. LCCN QC175 S77.

deBoer:1964:SSM

[dBU64] Jan Hendrik de Boer and George Eugene Uhlenbeck. Studies in Statistical Mechanics, volume II. North-Holland Publishing Co., Amsterdam, The Netherlands, 1964. vii + 272 pp. LCCN ????

deBoer:1965:SSM

[dBU65] Jan Hendrik de Boer and George Eugene Uhlenbeck. Studies in Statistical Mechanics, volume III. North-Holland Publishing Co., Amsterdam, The Netherlands, 1965. x + 388 pp. LCCN ????

deBoer:1978:DQK

[dBvW78] W. P. H. de Boer and Ch. G. van Weert. On the derivation of quantum kinetic equations. II. Nonlocal Uehling–Uhlenbeck equation. Journal of Statistical Physics, 18(3):281–292, March 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01018094.

Dellacherie:2003:CWCb

[Del03] Stéphane Dellacherie. Coupling of the Wang Chang–Uhlenbeck equations with the multispecies Euler system. Journal of Computational Physics, 189(1):239–276, July 20, 2003. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999103002109.

Deuschel:1989:IPE

[Deu89] Jean-Dominique Deuschel. Invariance principle and empirical mean large deviations of the critical Ornstein–Uhlenbeck process. Annals of Probability, 17(1):74–90, January 1989. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1176991495.

Diaconis:1979:RP

[DF79] Persi Diaconis and David Freedman. On rounding percentages. Journal of the American Statistical Association, 74(366):359–364, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284288. The

authors extend prior work on correctness of sums of rounded percentages [MYZ67], and criticize biased rounding practices in [Ben38].

Dieker:2013:PRP

[DG13] A. B. Dieker and Xuefeng Gao. Positive recurrence of piecewise Ornstein–Uhlenbeck processes and common quadratic Lyapunov functions. Annals of Applied Probability, 23(4):1291–1317, August 2013. CODEN ???? ISSN 1050-5164 (print), 2168-8737

(electronic). URL http://projecteuclid.org/euclid.aoap/
1371834029.

Dion:2016:NEM

[Dio16] Charlotte Dion. Nonparametric estimation in a mixed-effect Ornstein–Uhlenbeck model. Metrika. International Journal for Theoretical and Applied Statistics, 79(8):919–951, November 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s00184-016-0583-y.

Dirkse:1975:APO

[Dir75] John P. Dirkse. An absorption probability for the Ornstein– Uhlenbeck process. Journal of Applied Probability, 12(3):595–599, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212876.

Ditlevsen:2007:RFP

[Dit07] Susanne Ditlevsen. A result on the first-passage time of an Ornstein–Uhlenbeck process. Statistics & Probability Letters, 77(18):1744–1749, December 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715207001605.

Donelson:1972:CLV

[DM72] John Donelson III and Fred Maltz. A comparison of linear versus non-linear prediction for polynomial functions of the Ornstein– Uhlenbeck process. Journal of Applied Probability, 9(4):725–744, December 1972. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3212611.

Donado:2017:BMN

[DMLF^{+}17] F. Donado, R. E. Moctezuma, L. López-Flores, M. Medina-Noyola, and J. L. Arauz-Lara. Brownian motion in nonequilibrium systems and the Ornstein–Uhlenbeck stochastic process. Scientific Reports, 7(1), October 3, 2017. CODEN SRCEC3. ISSN 2045-2322.

Doyle:1962:DCN

[DMM62] Worthie Doyle, J. A. McFadden, and Imanuel Marx. The distribution of a certain nonlinear functional of an Ornstein–Uhlenbeck process. Journal of the Society for Industrial and Applied Mathematics, 10(2):381–393, June 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic).

Debbasch:1997:ROU

[DMR97] F. Debbasch, K. Mallick, and J. P. Rivet. Relativistic Ornstein– Uhlenbeck process. Journal of Statistical Physics, 88(3–4): 945–966, August 1997. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1023/B:JOSS.0000015180.16261.53.

Donati-Martin:2004:LDS

[DMRYZ04] C. Donati-Martin, A. Rouault, M. Yor, and M. Zani. Large deviations for squares of Bessel and Ornstein–Uhlenbeck processes. Probability Theory and Related Fields, 129(2):261–289, June 2004. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00440-004-0338-y.

Dector:2009:ACA

[DMTUV09] A. Déctor, H. A. Morales-Técotl, L. F. Urrutia, and J. D. Vergara. An alternative canonical approach to the ghost problem in a complexified extension of the Pais–Uhlenbeck oscillator. SIGMA Symmetry Integrability Geom. Methods Appl., 5:Paper 053, 22, 2009. ISSN 1815-0659.

Dzupire:2018:LPB

[DNO18] Nelson Christopher Dzupire, Philip Ngare, and Leo Odongo. Lévy process based Ornstein–Uhlenbeck temperature model with time varying speed of mean reversion. Advances and Applications in Statistics, 53(3):199–224, September 2018. CODEN ???? ISSN

0972-3617. URL http://www.pphmj.com/abstract/12012.htm.

Debbasch:1998:DER

[DR98] F. Debbasch and J. P. Rivet. A diffusion equation from the relativistic Ornstein–Uhlenbeck process. Journal of Statistical Physics, 90(5–6):1179–1199, March 1998. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1023/A:1023275210656.

deChaumaray:2017:LDS

[dRdC17] M. du Roy de Chaumaray. Large deviations for the squared radial Ornstein–Uhlenbeck process. Theory of Probability and its Applications, 61(3):408–441, ???? 2017. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dresden:1989:OGU

[Dre89] Max Dresden. Obituary: George E. Uhlenbeck. Physics Today, 42 (12):91–94, December 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Damaskinsky:2006:RQP

[DS06] E. V. Damaskinsky and M. A. Sokolov. Remarks on quantization of Pais–Uhlenbeck oscillators. Journal of Physics A (Mathematical and General), 39(33):10499–10507, 2006. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Dennison:1932:TMP

[DU32] David M. Dennison and G. E. Uhlenbeck. The two-minima problem and the ammonia molecule. Physical Review (2), 41(3):313– 321, August 1932. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Diop:2010:MDP

[DY10] Aliou Diop and Armel Fabrice Yode. Minimum distance parameter estimation for Ornstein–Uhlenbeck processes driven by Lévy process. Statistics & Probability Letters, 80(2):122–127, January 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209003678.

Dyson:1969:PTF

[Dys69] Freeman J. Dyson. Phase transitions in ferromagnets. Talk given at Evanston Statistical Mechanics Conference in honor of G. Uhlenbeck. Published in [Dys71]., October 31, 1969.

Dyson:1971:PTF

[Dys71] Freeman J. Dyson. Phase transitions in ferromagnets. In Cohen [Coh71], page ?? ISBN 0-8247-1111-4. LCCN QC175 .S7.

Enderle:1982:RLT

[EH82] K. Enderle and H. Hering. Ratio limit theorems for branching Ornstein–Uhlenbeck processes. Stochastic Processes and Their Applications, 13(1):75–85, 1982. CODEN STOPB7. ISSN 0304-4149 (print), 1879-209x (electronic).

Eliazar:2005:SOU

[EK05] Iddo Eliazar and Joseph Klafter. Stochastic Ornstein–Uhlenbeck capacitors. Journal of Statistical Physics, 118(1–2):177–198, January 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-004-8782-8.

Elworthy:1999:ISL

[ELY99] K. D. Elworthy, Xue-Mei Li, and M. Yor. The importance of strictly local martingales; applications to radial Ornstein– Uhlenbeck processes. Probability Theory and Related Fields, 115 (3):325–355, October 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/s004400050240.

Erickson:2005:GOU

[EM05] K. Bruce Erickson and Ross A. Maller. Generalised Ornstein– Uhlenbeck processes and the convergence of Lévy integrals. Lecture Notes in Mathematics, 1857:70–94, 2005. CODEN LN-MAA2. ISBN 3-540-23973-1 (print), 3-540-31449-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/content/pdf/10.1007/978-3-540-31449-
3_6.pdf.

Endo:2008:SMG

[EM08] Kotaro Endo and Muneya Matsui. The stationarity of multidimensional generalized Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 78(14):2265–2272, October 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771520800120X.

Escobedo:2007:FSL

[EMV07] M. Escobedo, S. Mischler, and J. J. L. Vélazquez. On the fundamental solution of a linearized Uehling–Uhlenbeck equation. Archive for Rational Mechanics and Analysis, 186(2):309–349, 2007. CODEN AVRMAW. ISSN 0003-9527 (print), 1432-0673 (electronic).

Escobedo:2008:SSU

[EMV08] M. Escobedo, S. Mischler, and J. J. L. Velázquez. Singular solutions for the Uehling–Uhlenbeck equation. Proceedings of the Royal Society of Edinburgh. Section A, Mathematical and Physical Sciences, 138(1):67–107, 2008. CODEN PEAMDU. ISSN 0308-2105 (print), 1473-7124 (electronic).

Englander:2010:CMS

[Eng10] Janos Englander. The center of mass for spatial branching processes and an application for self-interaction. Electronic Journal of Probability, 15:63:1938–63:1970, 2010. CODEN ???? ISSN

1083-6489. URL http://ejp.ejpecp.org/article/view/822.

Es-Sebaiy:2013:BEB

[ES13] Khalifa Es-Sebaiy. Berry–Esséen bounds for the least squares estimator for discretely observed fractional Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 83(10):2372–2385, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002447.

Emrich:1968:TPF

[ESU68] Raymond J. Emrich, Howard A. Snyder, and George E. Uhlenbeck. Teaching of physics of fluids in U.S. colleges and universities. American Journal of Physics, 36(10):886–895, October 1968. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Ehrenfest:1926:GVB

[EU26a] P. Ehrenfest and G. E. Uhlenbeck. Graphische Veranschaulichung der De Broglieschen Phasenwellen in der fünfdimensionalen Welt von O. Klein. (German) [Graphical illustration of de Broglie’s phase waves in the five-dimensional world by O. Klein]. Zeitschrift für Physik, 39(7–8):495–498, July 1926. CODEN ZEPYAA. ISSN 0044-3328.

Ehrenfest:1926:CDM

[EU26b] P. Ehrenfest and G. E. Uhlenbeck. On the connection of different methods of solution of the wave equation in multi-dimensional spaces. Verhandelingen der Koninklijke Akademie van Wetenschappen te Amsterdam. Wis- en natuurkundige vakken, 29(??): 1280–1285, ???? 1926.

Ehrenfest:1927:WIB

[EU27a] P. Ehrenfest and G. E. Uhlenbeck. Die wellenmechanische Interpretation der Boltzmannschen Statistik neben der der neueren Statistiken. (German) [The wave mechanical interpretation of Boltzmann statistics according to recent statistics]. Zeitschrift für Physik, 41(1):24–26, January 1927. CODEN ZEPYAA. ISSN 0044-3328.

Ehrenfest:1927:EMG

[EU27b] P. Ehrenfest and G. E. Uhlenbeck. Zum Einsteinschen ,,Mischungsparadoxon”. (German) [On Einstein’s “mixing paradox”]. Zeitschrift für Physik, 41(8–9):576–582, November 1927. CODEN ZEPYAA. ISSN 0044-3328.

Fannjiang:2004:CPS

[Fan04] Albert C. Fannjiang. Convergence of passive scalar fields in Ornstein–Uhlenbeck flows to Kraichnan’s model. Journal of Statistical Physics, 114(1–2):115–135, January 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1023/B:JOSS.
0000003106.02538.ae.

Fasen:2010:ARS

[Fas10] Vicky Fasen. Asymptotic results for sample autocovariance functions and extremes of integrated generalized Ornstein–Uhlenbeck processes. Bernoulli: official journal of the Bernoulli Society for Mathematical Statistics and Probability, 16(1):51–79, February 2010. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.bj/
1265984704.

Fasen:2013:SEM

[Fas13] Vicky Fasen. Statistical estimation of multivariate Ornstein– Uhlenbeck processes and applications to co-integration. Journal of Econometrics, 172(2):325–337, February 2013. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407612002035.

Fatalov:2009:OTE

[Fat09] V. R. Fatalov. Occupation time and exact asymptotics of distributions of L^{p}-functionals of the Ornstein–Uhlenbeck processes, p > 0. Theory of Probability and its Applications, 53(1):13–36, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fernique:1990:RCF

[Fer90] X. Fernique. Sur la régularité de certaines fonctions aléatoires d’Ornstein–Uhlenbeck. (French) [On the regularity of some Ornstein–Uhlenbeck random functions]. Annales de l’Institut Henri Poincaré, Probabilités et Statistiques, 26(3):399–417, ???? 1990. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017

(electronic). URL http://www.numdam.org/item?id=AIHPB_
1990__26_3_399_0.

Fernique:1992:RFA

[Fer92] X. Fernique. Sur la regularité des fonctions aléatoires d’Ornstein– Uhlenbeck à valeurs dans l_{p}, p ∈ [1, ∞[. (French) [On the regularity of random Ornstein–Uhlenbeck functions with values in l_{p}, p ∈ [1, ∞[]. Annals of Probability, 20(3):1441–1449, July 1992. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176989699.

Finkelstein:1966:KLW

[FH66] Leib Finkelstein and S. Harris. Kernel of the linearized Wang Chang—Uhlenbeck collision operator. Physics of Fluids, 9:8–11, 1966. CODEN PHFLE6. ISSN 0031-9171.

Foidl:1989:SGL

[FK89] Ch Foidl and P. Kasperkovitz. Systematic generation of linear graphs — check and extension of the list of Uhlenbeck and Ford. Journal of Computational Physics, 85(2):504, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999189901770.

Foidl:1990:SGLb

[FK90] Ch. Foidl and P. Kasperkovitz. Systematic generation of linear graphs — check and extension of the list of Uhlenbeck and Ford. Journal of Computational Physics, 89(1):246–250, July 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919090125K.

Fannjiang:2002:DLR

[FK02] Albert Fannjiang and Tomasz Komorowski. Diffusion in longrange correlated Ornstein–Uhlenbeck flows. Electronic Journal of Probability, 7:20:1–20:22, 2002. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/119.

Fink:2011:FLD

[FK11] Holger Fink and Claudia Klüppelberg. Fractional Lévy-driven Ornstein–Uhlenbeck processes and stochastic differential equations. Bernoulli: official journal of the Bernoulli Society for Mathematical Statistics and Probability, 17(1):484–506, February 2011. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1297173852.

Florens-Landais:1999:LDE

[FLP99] Danielle Florens-Landais and Huyên Pham. Large deviations in estimation of an Ornstein–Uhlenbeck model. Journal of Applied Probability, 36(1):60–77, March 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3215402.

Fleischmann:2006:HLF

[FMW06] Klaus Fleischmann, Peter Mörters, and Vitali Wachtel. Hydrodynamic limit fluctuations of super-Brownian motion with a stable catalyst. Electronic Journal of Probability, 11:29:723–29:767, 2006. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/348.

Ford:1956:CPTb

[FNU56] G. W. Ford, R. Z. Norman, and G. E. Uhlenbeck. Combinatorial problems in the theory of graphs. II. Proceedings of the National Academy of Sciences of the United States of America, 42:203– 208, 1956. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Fuhrman:2003:LCS

[FP03] Marco Fuhrman and Anna Maria Paganoni. Linear control systems on unbounded time intervals and invariant measures of Ornstein–Uhlenbeck processes in Hilbert spaces. SIAM Journal on Control and Optimization, 42(5):1776–1794, September 2003. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

URL http://epubs.siam.org/sam-bin/dbq/article/41465.

Franklin:2005:KUT

[Fra05] Allan Franklin. The Konopinski–Uhlenbeck theory of β decay: Its proposal and refutation. In Jed Z. Buchwald and Allan Franklin, editors, Wrong for the Right Reasons, volume 11 of Archimedes: New Studies in the History and Philosophy of Science and Technology, chapter 10, pages 209–228. Springer, Berlin, Heidelberg, New York, Tokyo, 2005. ISBN 1-4020-3048-7. ISSN 1385-0180 (print), 2215-0064 (electronic). URL https://link.springer.

com/chapter/10.1007/1-4020-3048-7_10.

Fermi:1933:REP

[FU33a] Enrico Fermi and George E. Uhlenbeck. On the recombination of electrons and positrons. Physical Review (2), 44(6):510–511, September 15, 1933. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v44/i6/p510_2.

Fermi:1933:SRE

[FU33b] Enrico Fermi and George E. Uhlenbeck. Sulla ricombinazione di elettroni e positroni. (Italian) [On the recombination of electrons and positrons]. Ricerca Scientifica ed il Progresso Tecnico, 4(2): 157–160, 1933. CODEN RSPTB6.

Falkoff:1950:BGA

[FU50a] David L. Falkoff and G. E. Uhlenbeck. On the beta–gammaangular correlation. Physical Review (2), 79(2):334–340, July 1950. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Falkoff:1950:DCS

[FU50b] David L. Falkoff and G. E. Uhlenbeck. On the directional correlation of successive nuclear radiations. Physical Review (2), 79(2): 323–333, July 1950. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Ford:1956:CPTa

[FU56a] G. W. Ford and G. E. Uhlenbeck. Combinatorial problems in the theory of graphs. I. Proceedings of the National Academy of Sciences of the United States of America, 42:122–128, 1956. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Ford:1956:CPTc

[FU56b] G. W. Ford and G. E. Uhlenbeck. Combinatorial problems in the theory of graphs. III. Proceedings of the National Academy of Sciences of the United States of America, 42:529–535, 1956. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Ford:1957:CPT

[FU57] G. W. Ford and G. E. Uhlenbeck. Combinatorial problems in the theory of graphs. IV. Proceedings of the National Academy of Sciences of the United States of America, 43:163–167, 1957. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Foch:1967:PSM

[FU67] J. Foch and G. E. Uhlenbeck. Propagation of sound in monatomic gases. Physical Review Letters, 19(18):1025–1027, October 1967. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Fox:1970:CNE

[FU70a] Ronald Forrest Fox and George E. Uhlenbeck. Contributions to non-equilibrium thermodynamics. I. Theory of hydrodynamical fluctuations. Physics of Fluids, 13(8):1893–1902, August 1970. CODEN PHFLE6. ISSN 1070-6631.

Fox:1970:CNT

[FU70b] Ronald Forrest Fox and George E. Uhlenbeck. Contributions to nonequilibrium thermodynamics. II. Fluctuation theory for the Boltzmann equation. Physics of Fluids, 13(12):2881–2890, December 1970. CODEN PHFLE6. ISSN 1070-6631.

Fox:1976:CNTb

[FU76] R. F. Fox and G. E. Uhlenbeck. Contributions to nonequilibrium thermodynamics II. fluctuation theory for the Boltzmann equation. Physics of Fluids, 13(??):2881–2890, ???? 1976. CODEN PHFLE6. ISSN 1070-6631.

Foch:1972:TSP

[FUF72] James D. Foch, Jr., George E. Uhlenbeck, and Manuel Fuentes Losa. Theory of sound propagation in mixtures of monatomic gases. Physics of Fluids, 15(7):1224–1232, July 1972. CODEN PHFLE6. ISSN 1070-6631.

Fierz:1960:TPT

[FW60] Markus Fierz and Victor Frederick Weisskopf. Theoretical physics in the Twentieth Century: a memorial volume to Wolfgang Pauli. Interscience Publishers, New York, NY, USA, 1960. x + 328 pp. LCCN QC3 .F52. See [?] for comments on some inaccuracies in this book, and on the history of the discovery of the spin of the electron and the nucleus.

Fang:1999:STB

[FZ99] S. Fang and T. S. Zhang. On the small time behavior of Ornstein–Uhlenbeck processes with unbounded linear drifts. Probability Theory and Related Fields, 114(4):487–504, July 1999. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s004400050232.

Gasanenko:2011:SDO

[Gas11] V. A. Gasanenko. The small deviation of Ornstein–Uhlenbeck process with large trend. Adv. Appl. Math. Sci., 10(2):159–167, 2011. ISSN 0974-6803.

Garcia-Colin:1966:GCU

[GCF66] Leopoldo S. García-Colín and Asdrúbal Flores. The generalization of Choh–Uhlenbeck’s method in the kinetic theory of dense gases. Journal of Mathematical Physics, 7(2):254–259, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v7/i2/p254_s1.

Goudsmit:1944:SFN

[GF44] S. A. Goudsmit and W. H. Furry. Significant figures of numbers in statistical tables. Nature, 154(3921):800–801, December 23, 1944. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v154/

n3921/pdf/154800a0.pdf. See later work [LG78], and severe criticism [Rai76, Rai85].

Gill:2013:SOU

[Gil13] Hardeep Gill. A super Ornstein–Uhlenbeck process interacting with its center of mass. Annals of Probability, 41(2):989–1029, March 2013. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1362750948.

Gao:2009:DIM

[GJ09a] Fuqing Gao and Hui Jiang. Deviation inequalities and moderate deviations for estimators of parameters in an Ornstein–Uhlenbeck process with linear drift. Electronic Communications in Probability, 14:21:210–21:223, 2009. CODEN ???? ISSN 1083-589X. URL

http://ecp.ejpecp.org/article/view/1466.

Gao:2009:MDS

[GJ09b] Fuqing Gao and Hui Jiang. Moderate deviations for squared radial Ornstein–Uhlenbeck process. Statistics & Probability Letters, 79(11):1378–1386, June 1, 2009. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209000765.

Gloter:2001:PED

[Glo01] Arnaud Gloter. Parameter estimation for a discrete sampling of an integrated Ornstein–Uhlenbeck process. Statistics: a Journal of Theoretical and Applied Statistics, 35(3):225–243, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Grahovac:2016:ISO

[GLST16] Danijel Grahovac, Nikolai N. Leonenko, Alla Sikorskii, and Irena Tesnjak. Intermittency of superpositions of Ornstein–Uhlenbeck type processes. Journal of Statistical Physics, 165(2):390–408, October 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-016-1616-7.

Gaver:1993:SAM

[GMS93] Donald P. Gaver, John A. Morrison, and Rogerio Silveira. Serviceadaptive multitype repairman problems. SIAM Journal on Applied Mathematics, 53(2):459–470, April 1993. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Giorno:2012:ROU

[GNdC12] V. Giorno, A. G. Nobile, and R. di Cesare. On the reflected Ornstein–Uhlenbeck process with catastrophes. Applied Mathematics and Computation, 218(23):11570–11582, August 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300312005498.

Garbaczewski:2000:OUC

[GO00] Piotr Garbaczewski and Robert Olkiewicz. Ornstein–Uhlenbeck– Cauchy process. Journal of Mathematical Physics, 41(10):6843– 6860, October 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goudsmit:1971:DES

[Gou71] Samuel A. Goudsmit. The discovery of the electron spin. Web

document, April 1971. URL http://rowdysites.msudenver.
edu/~dovej/courses/am/supplemental/goudsmit-spin.pdf;
 http://www.lorentz.leidenuniv.nl/history/spin/goudsmit.

html. English translation by J. H. van der Waals of Goudsmit’s lecture in Dutch for the golden jubilee of the Dutch Physical Society in April 1971.

Graversen:2011:RUC

[GP11] Svend-Erik Graversen and Jan Pedersen. Representations of Urbanik’s classes and multiparameter Ornstein–Uhlenbeck processes. Electronic Communications in Probability, 16:20:200– 20:212, 2011. CODEN ???? ISSN 1083-589X. URL http://

ecp.ejpecp.org/article/view/1621.

Green:1950:RPR

[Gre50] H. S. Green. Remarks on a paper by Riddell and Uhlenbeck. Journal of Chemical Physics, 18:1123–1124, 1950. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Gringorten:1968:EFT

[Gri68] Irving I. Gringorten. Estimating finite-time maxima and minima of a stationary Gaussian Ornstein–Uhlenbeck process by Monte Carlo simulation. Journal of the American Statistical Association, 63(324):1517–1521, December 1968. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2285900.

Griffin:2006:ING

[GS06] J. E. Griffin and M. F. J. Steel. Inference with non-Gaussian Ornstein–Uhlenbeck processes for stochastic volatility. Journal of Econometrics, 134(2):605–644, October 2006. CO-DEN JECMB6. ISSN 0304-4076 (print), 1872-6895 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304407605001570.

Griffin:2010:BIS

[GS10] J. E. Griffin and M. F. J. Steel. Bayesian inference with stochastic volatility models using continuous superpositions of non-Gaussian Ornstein–Uhlenbeck processes. Computational Statistics & Data Analysis, 54(11):2594–2608, November 1, 2010. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167947309002382.

Goudsmit:1925:OSV

[GU25] S. Goudsmit and G. E. Uhlenbeck. Opmerking over de spectra van waterstof en helium. (Dutch) [Note on the spectra of hydrogen and helium]. Physica (The Hague), 5(??):266–270, ???? 1925. CODEN PYSIA7. ISSN 0370-2707.

Goudsmit:1926:KQI

[GU26a] S. Goudsmit and G. E. Uhlenbeck. Die Kopplungsmöglichkeiten der Quantenvektoren im Atom. (German) [The coupling possibilities of quantum vectors in the atom]. Zeitschrift für Physik, 35(8–9):618–625, August 1926. CODEN ZEPYAA.

ISSN 0044-3328. URL http://link.springer.com/article/
10.1007/BF01379808.

Goudsmit:1926:RES

[GU26b] S. Goudsmit and G. E. Uhlenbeck. Over het roteerende electron en de structuur der spectra. (Dutch) [On the rotating electron and the structure of the spectra]. Physica (The Hague), 6(??):273–290, ???? 1926. CODEN PYSIA7. ISSN 0370-2707.

Giovanni:2010:MDG

[GZ10] Peccati Giovanni and Cengbo Zheng. Multi-dimensional Gaussian fluctuations on the Poisson space. Electronic Journal of Probability, 15:48:1487–48:1527, 2010. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/813.

Ho:2013:ATH

[HA13] Lam Si Tung Ho and Cécile Ané. Asymptotic theory with hierarchical autocorrelation: Ornstein–Uhlenbeck tree models. Annals of Statistics, 41(2):957–981, April 2013. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1369836966.

Harge:2005:CEC

[Har05] Gilles Hargé. Characterization of equality in the correlation inequality for convex functions, the U-conjecture. Annales de l’Institut Henri Poincaré, Probabilités et Statistiques, 41 (4):753–765, July/August 2005. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

www.numdam.org/item?id=AIHPB_2005__41_4_753_0;   http:
//www.sciencedirect.com/science/article/pii/S0246020304000639.

Heck:1999:PLD

[Hec99] Matthias K. Heck. Principles of large deviations for the empirical processes of the Ornstein–Uhlenbeck process. Journal of Theoretical Probability, 12(1):147–179, January 1999. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://

link.springer.com/article/10.1023/A:1021700811752.

Huang:2016:MMO

[HJM^{+}16] G. Huang, H. M. Jansen, M. Mandjes, P. Spreij, and K. De Turck. Markov-modulated Ornstein–Uhlenbeck processes. Advances in Applied Probability, 48(1):235–254, March 2016. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1457466164.

Hughes:2017:ELM

[HKST17] Rachael A. Hughes, Michael G. Kenward, Jonathan A. C. Sterne, and Kate Tilling. Estimation of the linear mixed integrated Ornstein–Uhlenbeck model. Journal of Statistical Computation and Simulation, 87(8):1541–1558, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Hemmer:1964:VWTa

[HKU64] P. C. Hemmer, M. Kac, and G. E. Uhlenbeck. On the van der Waals theory of the vapor–liquid equilibrium. III. Discussion of the critical region. Journal of Mathematical Physics, 5(1):60–74, January 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i1/p60_s1.

Hopfner:2016:ESH

[HLT16] R. Höpfner, E. Löcherbach, and M. Thieullen. Ergodicity for a stochastic Hodgkin–Huxley model driven by Ornstein–Uhlenbeck type input. Annales de l’Institut Henri Poincaré, Probabilités et Statistiques, 52(1):483–501, February 2016. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http://

projecteuclid.org/euclid.aihp/1452089277.

Hubert:1975:RBS

[HM75] M. Hubert and A. D. May. The Rayleigh–Brillouin spectrum of normal and parahydrogen: A test of model solutions of the Wang– Chang–Uhlenbeck equation. Canadian Journal of Physics = Journal canadien de physique, 53(4):343–350, February 1975. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (electronic).

Hu:2010:PEF

[HN10] Yaozhong Hu and David Nualart. Parameter estimation for fractional Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 80(11–12):1030–1038, June 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715210000659.

Hintze:2014:SNA

[HP14] Robert Hintze and Ilya Pavlyukevich. Small noise asymptotics and first passage times of integrated Ornstein–Uhlenbeck processes driven by α-stable Lévy processes. Bernoulli: official journal of the Bernoulli Society for Mathematical Statistics and Probability, 20(1):265–281, February 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1390407289.

Holley:1981:GOU

[HS81] R. Holley and D. Stroock. Generalized Ornstein–Uhlenbeck processes as limits of interacting systems. Lecture Notes in Mathematics, 851:152–168, 1981. CODEN LNMAA2. ISBN 3-540-10690-1 (print), 3-540-38613-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0088727/.

Hebb:1938:SNI

[HU38] Malcolm H. Hebb and G. E. Uhlenbeck. Stability of nuclear isomers. Physica, 5(7):605–608, July 1938. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Harary:1953:NHT

[HU53] Frank Harary and George E. Uhlenbeck. On the number of Husimi trees. I. Proceedings of the National Academy of Sciences of the United States of America, 39:315–322, 1953. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL https://

www.jstor.org/stable/88580.

He:1997:HOU

[HW97] Sheng-Wu He and Jia-Gang Wang. The hypercontractivity of Ornstein–Uhlenbeck semigroups with drift, revisited. Lecture Notes in Mathematics, 1655:80–84, 1997. CODEN LN-MAA2. ISBN 3-540-62634-4 (print), 3-540-68352-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0119295/.

Igloi:2005:ROT

[Igl05] Endre Iglói. A rate-optimal trigonometric series expansion of the fractional Brownian motion. Electronic Journal of Probability, 10: 41:1381–41:1397, 2005. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/287.

Iscoe:1989:LDV

[IM89] I. Iscoe and D. McDonald. Large deviations for l^{2}-valued Ornstein–Uhlenbeck processes. Annals of Probability, 17(1):58–73, January 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.
aop/1176991494.

Iscoe:1990:CVO

[IMM^{+}90] I. Iscoe, M. B. Marcus, D. McDonald, M. Talagrand, and J. Zinn. Continuity of l^{2}-valued Ornstein–Uhlenbeck processes. Annals of Probability, 18(1):68–84, January 1990. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176990938.

Igloi:1999:LRD

[IT99] E. Iglói and G. Terdik. Long-range dependence through gammamixed Ornstein–Uhlenbeck process. Electronic Journal of Probability, 4:16:1–16:33, 1999. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/53.

Ivezic:2010:GUG

[Ive10] Tomislav Ivezić. The generalized Uhlenbeck–Goudsmit hypothesis: ‘magnetic’ S^{a} and ‘electric’ Z^{a} spins. Physica Scripta, 81 (2):025001:1–025001:8, January 2010. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic).

Jacod:1985:GFP

[Jac85] Jean Jacod. Grossissement de filtration et processus d’Ornstein– Uhlenbeck généralisé. (French) []. Lecture Notes in Mathematics, 1118:36–44, 1985. CODEN LNMAA2. ISBN 3-540-15210-5 (print), 3-540-39339-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0075769/.

Jacobsen:1996:LOO

[Jac96] Martin Jacobsen. Laplace and the origin of the Ornstein– Uhlenbeck process. Bernoulli: official journal of the Bernoulli Society for Mathematical Statistics and Probability, 2(3):271–286, September 1996. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.bj/
1178291723.

Jiang:2015:PEN

[JD15] Hui Jiang and Xing Dong. Parameter estimation for the nonstationary Ornstein–Uhlenbeck process with linear drift. Statistical Papers, 56(1):257–268, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic). URL http://

link.springer.com/article/10.1007/s00362-014-0580-z.

Jimenez:2013:LGD

[JDD13] Jose Beltrán Jiménez, Enea Di Dio, and Ruth Durrer. A longitudinal gauge degree of freedom and the Pais–Uhlenbeck field. Journal of High Energy Physics, (4):030, front matter + 22, 2013. CODEN JHEPAB. ISSN 1126-6708.

Jegaraj:2009:STA

[Jeg09] Terence Jegaraj. Small time asymptotics of Ornstein–Uhlenbeck densities in Hilbert spaces. Electronic Communications in Probability, 14:53:552–53:559, 2009. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1510.

Jiang:2012:BEB

[Jia12] Hui Jiang. Berry–Esseen bounds and the law of the iterated logarithm for estimators of parameters in an Ornstein–Uhlenbeck process with linear drift. Journal of Applied Probability, 49(4): 978–989, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/

stable/41713823.

Jhwueng:2014:POU

[JM14] Dwueng-Chwuan Jhwueng and Vasileios Maroulas. Phylogenetic Ornstein–Uhlenbeck regression curves. Statistics & Probability Letters, 89(??):110–117, June 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715214000881.

Jorgensen:1978:CBM

[Jør78] E. Jørgensen. Construction of the Brownian motion and the Ornstein–Uhlenbeck process in a Riemannian manifold on basis of the Gangolli–McKean injection scheme. Zeitschrift für Wahrscheinlichkeitstheorie und verwandte Gebiete, 44(1):71–87, ???? 1978. CODEN ZWVGAA. ISSN 0044-3719. URL http:

//link.springer.com/article/10.1007/BF00534142.

Jensen:1999:OUT

[JP99] Jens Ledet Jensen and Jan Pedersen. Ornstein–Uhlenbeck type processes with non-normal distribution. Journal of Applied Probability, 36(2):389–402, June 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215463.

Julenbek:1978:R

[Jul78] Džordž Julenbek. Reminiscences. Bulgarska Akademiya na Naukite Fizicheski Institut. Matematicheski Institut. Fiziko-Matematichesko Spisanie, 21(54)(2):112–121, 1978. ISSN 0015-3265. Translated by N. Ahababjan.

Jongbloed:2005:NIL

[JVV05] G. Jongbloed, F. H. Van Der Meulen, and A. W. Van Der Vaart. Nonparametric inference for Lévy-driven Ornstein–Uhlenbeck processes. Bernoulli: official journal of the Bernoulli Society for Mathematical Statistics and Probability, 11(5):759–791, October 2005. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.bj/
1130077593.

Kac:1960:PSS

[Kac60] Mark Kac, editor. Proceedings of the Summer Seminar, Boulder, Colorado, 1960. American Mathematical Society, Providence, RI, USA, 1960. LCCN QC20 .S84 1960. Four volumes.

Kawai:2013:LAN

[Kaw13] Reiichiro Kawai. Local asymptotic normality property for Ornstein–Uhlenbeck processes with jumps under discrete sampling. Journal of Theoretical Probability, 26(4):932–967, December 2013. CODEN JTPREO. ISSN 0894-9840 (print), 1572-

9230 (electronic). URL http://link.springer.com/article/
10.1007/s10959-012-0455-y.

Kuhn:1967:SHQ

[KHFA67] Thomas S. Kuhn, John L. Heilbron, Paul Forman, and Lini Allen. Sources for history of quantum physics: an inventory and report, volume 68 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1967. ix + 176 pp. LCCN QC174.1 .S66.

URL http://www.amphilsoc.org/guides/ahqp/;   http://
www.amphilsoc.org/guides/ahqp/s-t.htm#schrodinger.

Knopova:2011:EAD

[KK11] Victoria Knopova and Alexei Kulik. Exact asymptotic for distribution densities of Lévy functionals. Electronic Journal of Probability, 16:52:1394–52:1433, 2011. CODEN ???? ISSN 1083-6489.

URL http://ejp.ejpecp.org/article/view/909.

Kiessling:2006:LFP

[KL06] Michael Kiessling and Carlo Lancellotti. The linear Fokker– Planck equation for the Ornstein–Uhlenbeck process as an (almost) nonlinear kinetic equation for an isolated N-particle system. Journal of Statistical Physics, 123(3):525–546, May 2006. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-006-9057-3.

Kiessling:2008:LFP

[KL08] Michael Kiessling and Carlo Lancellotti. The linear Fokker– Planck equation for the Ornstein–Uhlenbeck process as an (almost) nonlinear kinetic equation for an isolated N-particle system. Journal of Statistical Physics, 130(4):827, February 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-007-9447-1.

Kaparulin:2015:ESP

[KL15] D. S. Kaparulin and S. L. Lyakhovich. Energy and stability of the Pais–Uhlenbeck oscillator. In Geometric methods in physics, Trends Math., pages 127–134. Birkhäuser/Springer, Cham, Switzerland, 2015.

Krivonos:2016:MRC

[KLS16] Sergey Krivonos, Olaf Lechtenfeld, and Alexander Sorin. Minimal realization of ℓ-conformal Galilei algebra, Pais–Uhlenbeck oscillators and their deformation. Journal of High Energy Physics, (10): 078, front matter+16, 2016. CODEN JHEPAB. ISSN 1126-6708.

Kawai:2011:EDS

[KM11] Reiichiro Kawai and Hiroki Masuda. Exact discrete sampling of finite variation tempered stable Ornstein–Uhlenbeck processes. Monte Carlo Methods and Applications, 17(3):279–300, September 2011. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-3961

(electronic). URL http://www.degruyter.com/view/j/mcma.
2011.17.issue-3/mcma.2011.012/mcma.2011.012.xml.

Kawai:2012:IVT

[KM12] Reiichiro Kawai and Hiroki Masuda. Infinite variation tempered stable Ornstein–Uhlenbeck processes with discrete observations. Communications in Statistics: Simulation and Computation, 41 (1):125–139, 2012. CODEN CSSCDB. ISSN 0361-0918.

Kondo:2006:SPE

[KMiS06] Hitoshi Kondo, Makoto Maejima, and Ken iti Sato. Some properties of exponential integrals of Lévy processes and examples. Electronic Communications in Probability, 11:30:291–30:303, 2006. CODEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/1232.

Kolokoltsov:2001:SDF

[Kol01] Vassili Kolokoltsov. Small diffusion and fast dying out asymptotics for superprocesses as non-Hamiltonian quasiclassics for evolution equations. Electronic Journal of Probability, 6:21:1–21:16, 2001. CODEN ???? ISSN 1083-6489. URL http://ejp.ejpecp.org/

article/view/94.

Kozak:1966:BRS

[Koz66] John J. Kozak. Book review: Studies in Statistical Mechanics, Vol. II by J. de Boer and G. E. Uhlenbeck. American Scientist, 54(1):

143A–144A, March 1966. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL https://www.jstor.org/

stable/27836374.

Kutoyants:1994:MNE

[KP94] Y. Kutoyants and P. Pilibossian. On minimum L_{1}-norm estimate of the parameter of the Ornstein–Uhlenbeck process. Statistics & Probability Letters, 20(2):117–123, May 27, 1994. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0167715294900264.

Ko:2004:CFQ

[KP04] Chul Ki Ko and Yong Moon Park. Construction of a family of quantum Ornstein–Uhlenbeck semigroups. Journal of Mathematical Physics, 45(2):609–627, February 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i2/p609_s1.

Kulkarni:1994:FMD

[KR94] Vidyadhar Kulkarni and Tomasz Rolski. Fluid model driven by an Ornstein–Uhlenbeck process. Probability in the Engineering and Informational Sciences, 8(3):403–417, July 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4ED62752F39F5B6B5D20C63ABE4E152B.

Kronig:1960:TP

[Kro60] Ralph Kronig. The turning point. In Theoretical physics in the Twentieth Century: a memorial volume to Wolfgang Pauli [FW60], page ?? LCCN QC3 .F52. See [?] for comments on some inaccuracies in this book, and on the history of the discovery of the spin of the electron and the nucleus.

Kiselak:2008:EOD

[KS08] Jozef Kiselák and Milan Stehlík. Equidistant and D-optimal designs for parameters of Ornstein–Uhlenbeck process. Statistics & Probability Letters, 78(12):1388–1396, September 1, 2008. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715207004208.

Konopinski:1935:FTR

[KU35a] E. J. Konopinski and G. E. Uhlenbeck. On the Fermi theory of βradioactivity. Physical Review (2), 48(1):7–12, July 1935. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Konopinski:1935:HOD

[KU35b] E. J. Konopinski and George E. Uhlenbeck. Higher order derivatives in the interaction “Ansatz” of the Fermi theory. Physical Review (2), 48(1):107–108, July 1935. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Knipp:1936:EGR

[KU36] J. K. Knipp and G. E. Uhlenbeck. Emission of gamma radiation during the beta decay of nuclei. Physica, 3(6):425–439, June 1936. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Kahn:1937:TC

[KU37] B. Kahn and G. E. Uhlenbeck. On the theory of condensation. Physica, 4(10):1155–1156, October 1937. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Kahn:1938:TC

[KU38] B. Kahn and G. E. Uhlenbeck. On the theory of condensation. Physica, 5(5):399–416, May 1938. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Konopinski:1941:FTR

[KU41] E. J. Konopinski and G. E. Uhlenbeck. On the Fermi theory of β-radioactivity. II. The “forbidden” spectra. Physical Review (2), 60(4):308–320, August 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Kac:1959:PRT

[KU59] Mark Kac and George E. Uhlenbeck, editors. Probability and Related Topics in Physical Sciences: Proceedings of the Summer Seminar, Boulder, Colorado, 1957, volume 1 of Lectures in applied mathematics. American Mathematical Society, Providence, RI, USA, 1959. ISBN 0-8218-0047-7. LCCN QA273 .K11 197.

Kac:1963:VWT

[KUH63] M. Kac, G. E. Uhlenbeck, and P. C. Hemmer. On the van der Waals theory of the vapor–liquid equilibrium. I. Discussion of a one-dimensional model. Journal of Mathematical Physics, 4(2): 216–228, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i2/p216_s1.

Kuwabara:2016:TDP

[KYH16] Hirosuke Kuwabara, Tsukasa Yumibayashi, and Hiromitsu Harada. Time-dependent Pais–Uhlenbeck oscillator and its decomposition. In Geometric methods in physics, Trends Math., pages 255–260. Birkhäuser/Springer, Cham, Switzerland, 2016.

Lachaud:2005:CHT

[Lac05] B. Lachaud. Cut-off and hitting times of a sample of Ornstein– Uhlenbeck processes and its average. Journal of Applied Probability, 42(4):1069–1080, December 2005. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/30040829.

Laurence:1956:PGS

[Lau56] William L. Laurence. Physics goal set by Oppenheimer: Atomic scientist envisions quest for explaining new order in the universe questions to be answered obligations for the teacher. New York Times, ??(??):13, February 3, 1956. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/113908716/.

Laurence:1958:SRD

[Lau58] William L. Laurence. Science in review: Discoveries on the frontiers of physics are explained at a Harvard conference. New York Times, ??(??):E9, March 30, 1958. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/114447661/.

Lefebvre:1997:FHP

[Lef97] Mario Lefebvre. First hitting place distributions for the Ornstein– Uhlenbeck process. Statistics & Probability Letters, 34(3):309– 312, June 16, 1997. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715296001952.

Logan:1978:FDP

[LG78] Jonothan L. Logan and Samuel A. Goudsmit. The first digit phenomenon. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 122(4):193–197, August 18, 1978. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://links.jstor.org/sici?
sici=0003-049X(19780818)122:4%3C193:TFDP%3E2.0.CO%3B2-
C;   http://www.jstor.org/stable/986530. This paper con-

tains derivations of both Stigler’s Law and Benford’s Law, and receives strong criticism in [Rai85]. This paper contains an important historical note that is recorded in entry [Ben38].

Li:2000:OUT

[Li00] Zeng-Hu Li. Ornstein–Uhlenbeck type processes and branching processes with immigration. Journal of Applied Probability, 37 (3):627–634, September 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215600.

Lindner:2009:CPI

[LiS09] Alexander Lindner and Ken iti Sato. Continuity properties and infinite divisibility of stationary distributions of some generalized Ornstein–Uhlenbeck processes. Annals of Probability, 37(1): 250–274, January 2009. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1234881690.

Liu:2008:SSR

[Liu08] Kai Liu. Stationary solutions of retarded Ornstein–Uhlenbeck processes in Hilbert spaces. Statistics & Probability Letters, 78 (13):1775–1783, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715208000394.

Liu:2012:RPR

[Liu12] Kai Liu. On regularity property of retarded Ornstein–Uhlenbeck processes in Hilbert spaces. Journal of Theoretical Probability, 25(2):565–593, June 2012. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic). URL http://link.springer.

com/article/10.1007/s10959-011-0374-3.

Larsen:2016:GUB

[LLAT16] Sigurd Yves Larsen, Monique Lassaut, and Alejandro Amaya-Tapia. A generalized Uhlenbeck and Beth formula for the third cluster coefficient. Annals of Physics, 374:291–313, 2016. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035x (electronic).

Long:2009:LSE

[Lon09] Hongwei Long. Least squares estimator for discretely observed Ornstein–Uhlenbeck processes with small Lévy noises. Statistics & Probability Letters, 79(19):2076–2085, October 1, 2009. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715209002429. See note [Ma10].

Lorentz:1936:U

[Lor36] H. A. Lorentz. [unknown]. In H. A. Lorentz: Collected Works, volume 7, pages 179–?? Nyhoff, The Hague, The Netherlands, 1936.

Leon:2000:AOU

[LP00] José R. León and Gonzalo Perera. Approximation of the Ornstein– Uhlenbeck local time by harmonic oscillators. Bernoulli: official journal of the Bernoulli Society for Mathematical Statistics and Probability, 6(2):357–379, April 2000. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://

projecteuclid.org/euclid.bj/1081788033.

Luschgy:2009:EGP

[LP09] Harald Luschgy and Gilles Pagès. Expansions for Gaussian processes and Parseval frames. Electronic Journal of Probability, 14: 42:1198–42:1221, 2009. CODEN ???? ISSN 1083-6489. URL

http://ejp.ejpecp.org/article/view/649.

Lladser:2000:DAQ

[LS00] Manuel Lladser and Jaime San Martín. Domain of attraction of the quasi-stationary distributions for the Ornstein–Uhlenbeck process. Journal of Applied Probability, 37(2):511–520, June 2000. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3215725.

Lisovskii:2019:STT

[LS19] D. I. Lisovskii and A. N. Shiryaev. Sequential testing of two hypotheses for a stationary Ornstein–Uhlenbeck process. Theory of Probability and its Applications, 63(4):580–593, ???? 2019. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Laporte:1931:ASA

[LU31] Otto Laporte and George E. Uhlenbeck. Application of spinor analysis to the Maxwell and Dirac equations. Physical Review (2), 37(11):1380–1397, June 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Lawson:1950:TS

[LU50] James L. Lawson and George E. Uhlenbeck, editors. Threshold Signals, volume 24 of Radiation laboratory series. McGraw-Hill, New York, NY, USA, 1950. xii + 388 pp.

Lawson:1965:TS

[LU65] James L. Lawson and George E. Uhlenbeck, editors. Threshold Signals, volume 24 of Radiation laboratory series. Dover, New York, NY, USA, 1965. xii + 388 pp.

Liu:2016:TRG

[LZ16] Yong Liu and Jianliang Zhai. Time regularity of generalized Ornstein–Uhlenbeck processes with Lévy noises in Hilbert spaces. Journal of Theoretical Probability, 29(3):843–866, September 2016. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-015-0594-z.

Ma:2010:NLS

[Ma10] Chunhua Ma. A note on “Least squares estimator for discretely observed Ornstein–Uhlenbeck processes with small Lévy noises”. Statistics & Probability Letters, 80(19–20):1528–1531, October 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001677. See [Lon09].

Mai:2014:EML

[Mai14] Hilmar Mai. Efficient maximum likelihood estimation for Lévydriven Ornstein–Uhlenbeck processes. Bernoulli: official journal of the Bernoulli Society for Mathematical Statistics and Probability, 20(2):919–957, May 2014. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1393594010.

Martellini:1981:PUR

[Mar81] M. Martellini. Pais–Uhlenbeck regularization of free quantum gravity: toward a unitary theory. Il Nuovo Cimento B (11), 64 (1):33–40, 1981. CODEN NCIBAW. ISSN 0369-4100.

Masuda:2004:MOU

[Mas04] Hiroki Masuda. On multidimensional Ornstein–Uhlenbeck processes driven by a general Lévy process. Bernoulli: official journal of the Bernoulli Society for Mathematical Statistics and Probability, 10(1):97–120, February 2004. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1077544605.

Masiero:2007:RPT

[Mas07] Federica Masiero. Regularizing properties for transition semigroups and semilinear parabolic equations in Banach spaces. Electronic Journal of Probability, 12:13:387–13:419, 2007. CODEN

???? ISSN 1083-6489. URL http://ejp.ejpecp.org/article/
view/401.

Masterov:2015:SPU

[Mas15] Ivan Masterov. N = 2 supersymmetric Pais–Uhlenbeck oscillator. Modern Phys. Lett. A, 30(21):1550107, 12, 2015. ISSN 0217-7323 (print), 1793-6632 (electronic).

Masterov:2016:SOO

[Mas16a] Ivan Masterov. N = 2 supersymmetric odd-order Pais–Uhlenbeck oscillator. Nuclear Physics B, 910:40–54, 2016. CODEN NUPBBO. ISSN 0550-3213 (print), 1873-1562 (electronic).

Masterov:2016:AHF

[Mas16b] Ivan Masterov. An alternative Hamiltonian formulation for the Pais–Uhlenbeck oscillator. Nuclear Physics B, 902:95–114, 2016. CODEN NUPBBO. ISSN 0550-3213 (print), 1873-1562 (electronic).

Masterov:2016:OOP

[Mas16c] Ivan Masterov. The odd-order Pais–Uhlenbeck oscillator. Nuclear Physics B, 907:495–508, 2016. CODEN NUPBBO. ISSN 0550-3213 (print), 1873-1562 (electronic).

Mazzolo:2017:COU

[Maz17] Alain Mazzolo. Constraint Ornstein–Uhlenbeck bridges. Journal of Mathematical Physics, 58(9):093302, September 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McLennan:1962:BRS

[McL62] J. A. McLennan. Book review: Studies in Statistical Mechanics. Vol. 1, J. de Boer and G. E. Uhlenbeck, Editors. Pp. 350, Vol. 1. Interscience Division, John Wiley & Sons, Inc., New York, 1962. American Journal of Physics, 30(11):848, November 1962. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Mannheim:2005:DQP

[MD05] Philip D. Mannheim and Aharon Davidson. Dirac quantization of the Pais–Uhlenbeck fourth order oscillator. Physical Review A (3), 71(4):042110, 9, 2005. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Mehra:1973:PCN

[Meh73] Jagdish Mehra, editor. The physicist’s conception of nature: Symposium on the Development of the Physicist’s Conception of Nature in the 20th century. Held at the International Centre for Theoretical Physics, Miramare, Trieste, Italy, 18–25 September 1972. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1973. ISBN 90-277-0345-0, 90-277-2536-5. LCCN QC173.96 .S95 1972. URL http://www.springer.

com/us/book/9789027703453.

Meixner:1965:SME

[Mei65] Josef Meixner, editor. Statistical Mechanics of equilibrium and non-equilibrium; Proceedings of the International Symposium on Statistical Mechanics and Thermodynamics. Held at Aachen, Germany, 15–20 June 1964. North-Holland Publishing Co., Amsterdam, The Netherlands, 1965.

Miao:2013:ADO

[Mia13] Daniel Wei-Chung Miao. Analysis of the discrete Ornstein– Uhlenbeck process caused by the tick size effect. Journal of Applied Probability, 50(4):1102–1116, December 2013. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1389370102.

Mingalev:1992:SLC

[Min92a] I. V. Mingalev. On the solvability in the large of the Cauchy problem for discrete models of the Uehling–Uhlenbeck equation. Doklady Akademii nauk SSSR, 324(1):42–45, 1992. CODEN DANKAS. ISSN 0869-5652.

Mingalev:1992:SFT

[Min92b] O. V. Mingalev. Solutions in the form of a traveling wave in a discrete model of the Uehling–Uhlenbeck equation. Doklady Akademii nauk SSSR, 323(6):1029–1033, 1992. CODEN DANKAS. ISSN 0869-5652.

Mitoma:1987:GOU

[Mit87] Itaru Mitoma. Generalized Ornstein–Uhlenbeck process having a characteristic operator with polynomial coefficients. Probability Theory and Related Fields, 76(4):533–555, ???? 1987. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL

http://link.springer.com/article/10.1007/BF00960073.

Mallick:2005:AOD

[MM05] Kirone Mallick and Philippe Marcq. Anharmonic oscillator driven by additive Ornstein–Uhlenbeck noise. Journal of Statistical Physics, 119(1–2):1–33, April 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-004-2135-5.

Morita:1994:DPIb

[Mor94] Akio Morita. Diffusion processes with inertial effects and with boundary conditions — a solution to the Wang and Uhlenbeck problem. Journal of Mathematical Chemistry, 16(1):49–60, December 1994. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-

8897 (electronic). URL http://link.springer.com/article/
10.1007/BF01169195.

Morrison:2007:SAW

[Mor07] Margaret Morrison. Spin: All is not what it seems. Studies in History and Philosophy of Modern Physics, 38(3):529–557, September 2007. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000931.

Mostafazadeh:2010:HFP

[Mos10] Ali Mostafazadeh. A Hamiltonian formulation of the Pais– Uhlenbeck oscillator that yields a stable and unitary quantum system. Physics Letters A, 375(2):93–98, 2010. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

Millet:1992:COU

[MS92] A. Millet and W. Smoleński. On the continuity of Ornstein– Uhlenbeck processes in infinite dimensions. Probability Theory and Related Fields, 92(4):529–547, ???? 1992. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:

//link.springer.com/article/10.1007/BF01274267.

Matsui:2009:EFO

[MS09a] Muneya Matsui and Narn-Rueih Shieh. On the exponentials of fractional Ornstein–Uhlenbeck processes. Electronic Journal of Probability, 14:23:594–23:611, 2009. CODEN ???? ISSN 1083-

6489. URL http://ejp.ejpecp.org/article/view/628.

Muller:2009:BRU

[MS09b] Frank Müller and Armin Schikorra. Boundary regularity via Uhlenbeck–Rivière decomposition. Analysis (Munich), 29(2):199– 220, 2009. ISSN 0174-4747.

Merritt:1933:OPE

[MUG^{+}33] E. Merritt, G. E. Uhlenbeck, S. Goudsmit, G. H. Dieke, and H. W. Stunkard. Obituary: Paul Ehrenfest. Science, 78(2026):377–378, October 27, 1933. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Maejima:2003:LMS

[MY03] Makoto Maejima and Kenji Yamamoto. Long-memory stable Ornstein–Uhlenbeck processes. Electronic Journal of Probability, 8:19:1–19:18, 2003. CODEN ???? ISSN 1083-6489. URL http:/

/ejp.ejpecp.org/article/view/168.

Masuda:2013:EEI

[MY13] Hiroki Masuda and Nakahiro Yoshida. Edgeworth expansion for the integrated Lévy driven Ornstein–Uhlenbeck process. Electronic Communications in Probability, 18:94:1–94:10, 2013. CO-DEN ???? ISSN 1083-589X. URL http://ecp.ejpecp.org/

article/view/2726.

Mosteller:1967:DSR

[MYZ67] Frederick Mosteller, Cleo Youtz, and Douglas Zahn. The distribution of sums of rounded percentages. Demography, 4(2):850– 858, June 1967. CODEN ???? ISSN 0070-3370 (print), 1533-

7790 (electronic). URL http://muse.jhu.edu/journals/dem/
;   http://www.biomedsearch.com/nih/Distribution-Sums-
Rounded-Percentages/21318695.html;   http://www.jstor.

org/stable/2060324. See [DF79] for further work.

Nordsieck:1940:TCR

[NLU40] A. Nordsieck, W. E. Lamb, Jr., and G. E. Uhlenbeck. On the theory of cosmic-ray showers. I. The Furry model and the fluctuation problem. Physica, 7(4):344–360, April 1940. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Norris:1998:OUP

[Nor98] J. R. Norris. Ornstein–Uhlenbeck processes indexed by the circle. Annals of Probability, 26(2):465–478, April 1998. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1022855640.

Novikov:2004:MFP

[Nov04] A. Novikov. Martingales and first-passage times for Ornstein– Uhlenbeck processes with a jump component. Theory of Probability and its Applications, 48(2):288–303, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL

http://epubs.siam.org/sam-bin/dbq/article/98040.

Nobile:1985:ETO

[NRS85] A. G. Nobile, L. M. Ricciardi, and L. Sacerdote. Exponential trends of Ornstein–Uhlenbeck first-passage-time densities. Journal of Applied Probability, 22(2):360–369, June 1985. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3213779.

OConnor:2001:GEU

[OR01] J. J. O’Connor and E. F. Robertson. George Eugene Uhlenbeck. Web site at the School of Mathematics and Statistics, University of St Andrews, Scotland., May 2001. URL http://www-history.

mcs.st-andrews.ac.uk/Biographies/Uhlenbeck.html.

oglu:2006:TMG

[oT06] Özgür Sarıoğlu and Bayram Tekin. Topologically massive gravity as a Pais–Uhlenbeck oscillator. Classical and quantum gravity, 23(24):7541–7549, 2006. CODEN CQGRDG. ISSN 0264-9381 (print), 1361-6382 (electronic).

Oppenheim:2004:ARP

[OV04] Georges Oppenheim and Marie-Claude Viano. Aggregation of random parameters Ornstein–Uhlenbeck or AR processes: some convergence results. Journal of Time Series Analysis, 25(3):335–350, May 2004. CODEN JTSADL. ISSN 0143-9782 (print), 1467-9892 (electronic).

Pais:2000:GSP

[Pai00] Abraham Pais. The Genius of Science: a Portrait Gallery. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850614-7 (hardcover). 356 pp. LCCN Q141 .P29 2000.

URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
;   http://www.loc.gov/catdir/toc/fy02/99046603.html.

Patie:2007:TSE

[Pat07] Pierre Patie. Two-sided exit problem for a spectrally negative αstable Ornstein–Uhlenbeck process and the Wright’s generalized hypergeometric functions. Electronic Communications in Probability, 12:16:146–16:160, 2007. CODEN ???? ISSN 1083-589X.

URL http://ecp.ejpecp.org/article/view/1265.

Pavsic:2016:PUO

[Pav16] Matej Pavšič. Pais-Uhlenbeck oscillator and negative energies. Int. J. Geom. Methods Mod. Phys., 13(9):1630015, 35, 2016. ISSN 0219-8878.

Pedersen:2002:POU

[Ped02] Jan Pedersen. Periodic Ornstein–Uhlenbeck processes driven by Lévy processes. Journal of Applied Probability, 39(4):748–763, December 2002. CODEN JPRBAM. ISSN 0021-9002 (print),

1475-6072 (electronic). URL http://www.jstor.org/stable/
3216000.

Pramanik:2013:SCS

[PG13] Souvik Pramanik and Subir Ghosh. On the stability of coherent states for Pais–Uhlenbeck oscillator. Modern Phys. Lett. A, 28(12): 1350038, 9, 2013. ISSN 0217-7323 (print), 1793-6632 (electronic).

Piechor:1991:CFF

[Pie91] Kazimierz Piechór. The Couette flow by a four-velocity model of the Uehling–Uhlenbeck equation. In Discrete models of fluid dynamics (Figueira da Foz, 1990), volume 2 of Ser. Adv. Math. Appl. Sci., pages 176–185. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1991.

Piechor:1992:SWS

[Pie92] Kazimierz Piechór. The shock-wave structure by a four-velocity model of the Uehling–Uhlenbeck equations. In Thermodynamics and kinetic theory (Ma̧dralin, 1990), volume 12 of Ser. Adv. Math. Appl. Sci., pages 146–155. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1992.

Putterman:1972:POQ

[PKU72] Seth J. Putterman, Mark Kac, and George E. Uhlenbeck. Possible origin of the quantized vortices in He II. Physical Review Letters, 29(9):546–549, August 1972. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Polityukov:1990:SCP

[Pol90] V. P. Polityukov. Solution of the Cauchy problem for the nonlinear Uehling–Uhlenbeck problem. Doklady Akademii nauk SSSR, 312 (6):1357–1360, 1990. CODEN DANKAS. ISSN 0002-3264.

Pedersen:2005:CDP

[PS05] Jan Pedersen and Ken-Iti Sato. The class of distributions of periodic Ornstein–Uhlenbeck processes driven by Lévy processes. Journal of Theoretical Probability, 18(1):209–235, January 2005. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10959-004-2595-1.

Pigorsch:2009:DSD

[PS09] Christian Pigorsch and Robert Stelzer. On the definition, stationary distribution and second order structure of positive semidefinite Ornstein–Uhlenbeck type processes. Bernoulli: official journal of the Bernoulli Society for Mathematical Statistics and Probability, 15(3):754–773, August 2009. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic). URL http://projecteuclid.

org/euclid.bj/1251463280.

Pulgar:2015:HOL

[PSLL15] Gustavo Pulgar, Joel Saavedra, Genly Leon, and Yoelsy Leyva. Higher order Lagrangians inspired by the Pais–Uhlenbeck oscillator and their cosmological applications. Journal of Cosmology and Astroparticle Physics, (5):046, front matter + 28, 2015. ISSN 1475-7516.

Prakapenia:2018:NSTb

[PSV18] M. A. Prakapenia, I. A. Siutsou, and G. V. Vereshchagin. Numerical scheme for treatment of Uehling–Uhlenbeck equation for two-particle interactions in relativistic plasma. Journal of Computational Physics, 373(??):533–544, November 15, 2018. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999118304650.

Pais:1949:TIC

[PU49] A. Pais and G. E. Uhlenbeck. T3. On the interaction of charged spin 0 particles with the electromagnetic field. Physical Review (2), 75(8):1321, April 15, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL https://journals.

aps.org/pr/pdf/10.1103/PhysRev.75.1279. Brief contribution

to the Minutes of the 1948 Annual Meeting at New York, January 26–29, 1949.

Pais:1950:FTN

[PU50] A. Pais and G. E. Uhlenbeck. On field theories with non-localized action. Physical Review (2), 79(1):145–165, July 1950. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Pais:1959:QTT

[PU59] A. Pais and G. E. Uhlenbeck. On the quantum theory of the third virial coefficient. Physical Review (2), 116(2):250–269, October 1959. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Putterman:1969:TER

[PU69] Seth Putterman and George E. Uhlenbeck. Thermodynamic equilibrium of rotating superfluids. Physics of Fluids, 12(11):2299– 2236, November 1969. CODEN PHFLE6. ISSN 1070-6631.

Qian:1995:HOU

[QH95] Zhongmin Qian and Sheng-Wu He. On the hypercontractivity of Ornstein–Uhlenbeck semigroups with drift. Lecture Notes in Mathematics, 1613:202–217, 1995. CODEN LNMAA2. ISBN 3-540-60219-4 (print), 3-540-44744-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0094212/.

Raimi:1976:FDP

[Rai76] Ralph A. Raimi. The first digit problem. American Mathematical Monthly, 83(7):521–538, August/September 1976. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2319349.

Raimi:1985:FDP

[Rai85] Ralph A. Raimi. The first digit phenomenon again. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 129(2):211–219, June 1985. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL http://www.jstor.org/stable/986989. This paper con-

tains strong criticism of a derivation of Benford’s Law [LG78].

Rigotti:2005:TGR

[RD05] M. Rigotti and F. Debbasch. An H-theorem for the general relativistic Ornstein–Uhlenbeck process. Journal of Mathematical Physics, 46(10):103303, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v46/i10/p103303_s1.

Ribeiro:2013:OUL

[RE13] Bruno V. Ribeiro and Yves Elskens. Ornstein–Uhlenbeck limit for the velocity process of an N-particle system interacting stochastically. Journal of Statistical Physics, 153(4):626–640, November 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-013-0835-4.

Reich:1963:RPP

[Rei63] Edgar Reich. Recent publications and presentations: Studies in Statistical Mechanics, Volume I, by J. De Boer and G. E. Uhlenbeck. American Mathematical Monthly, 70(6):684–685, June/July 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Renshaw:1987:DUO

[Ren87] Eric Renshaw. The discrete Uhlenbeck–Ornstein process. Journal of Applied Probability, 24(4):908–917, December 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214215.

Rockner:1992:PMB

[Roc92] Michael Rockner. On the parabolic Martin boundary of the Ornstein–Uhlenbeck operator on Wiener space. Annals of Probability, 20(2):1063–1085, April 1992. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://

projecteuclid.org/euclid.aop/1176989818.

Romer:1991:EMP

[Rom91] Robert H. Romer. Editorial: Memorable papers from the American Journal of Physics, 1933–1990. American Journal of Physics, 59(3):201–207, March 1991. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Rowlinson:1986:PCG

[Row86] J. S. Rowlinson. The properties of connected graphs and some corrections to the list of Uhlenbeck and Ford. Journal of Statistical Physics, 42(5–6):931–933, March 1986. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/BF01010453.

Roberts:2004:BIN

[RPD04] Gareth O. Roberts, Omiros Papaspiliopoulos, and Petros Dellaportas. Bayesian inference for non-Gaussian Ornstein–Uhlenbeck stochastic volatility processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(2):369–393, May 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/stable/3647531.

Ricciardi:1987:PDO

[RS87] L. M. Ricciardi and L. Sacerdote. On the probability densities of an Ornstein–Uhlenbeck process with a reflecting boundary. Journal of Applied Probability, 24(2):355–369, June 1987. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/3214260.

Ricciardi:1988:FPT

[RS88] Luigi M. Ricciardi and Shunsuke Sato. First-passage-time density and moments of the Ornstein–Uhlenbeck process. Journal of Applied Probability, 25(1):43–57, March 1988. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214232.

Raknerud:2012:IIM

[RS12] Arvid Raknerud and Øivind Skare. Indirect inference methods for stochastic volatility models based on non-Gaussian Ornstein–Uhlenbeck processes. Computational Statistics & Data Analysis, 56(11):3260–3275, November 2012. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167947311000302.

Rose:1935:FEP

[RU35] Morris E. Rose and George E. Uhlenbeck. The formation of electron–positron pairs by internal conversion of γ-radiation. Physical Review (2), 48(3):211–223, August 1935. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Riddell:1950:NPC

[RU50] R. J. Riddell, Jr. and G. E. Uhlenbeck. On the notion of pressure in a canonical ensemble. Journal of Chemical Physics, 18:1066– 1069, 1950. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Riddell:1953:TVD

[RU53] R. J. Riddell, Jr. and G. E. Uhlenbeck. On the theory of the virial development of the equation of state of mono-atomic gases. Journal of Chemical Physics, 21(11):2056–2064, November 1, 1953. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

S:1964:BRT

[S.64] D. S. Book review: The Theory of Linear Graphs with Applications to the Theory of the Virial Development of the Properties of Gases by G. E. Uhlenbeck, G. W. Ford. Mathematics of Computation, 18 (86):338, April 1964. CODEN MCMPAF. ISSN 0025-5718 (print),

1088-6842 (electronic). URL https://www.jstor.org/stable/
2003333.

Salminen:1984:COU

[Sal84] P. Salminen. On conditional Ornstein–Uhlenbeck processes. Advances in Applied Probability, 16(4):920–922, December 1984. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427347.

Schmuland:1990:SPP

[Sch90] B. Schmuland. Sample path properties of l^{p}-valued Ornstein– Uhlenbeck processes. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 33(3):358–368, September 1990. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Schmuland:1993:NOU

[Sch93] Byron Schmuland. Nonsymmetric Ornstein–Uhlenbeck processes in Banach space via Dirichlet forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(6):1324–1338, December 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schroder:2014:CCV

[Sch14] Michael Schröder. On contingent-claim valuation in continuoustime for volatility models of Ornstein–Uhlenbeck type. Journal of Computational and Applied Mathematics, 260(??):36–53, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042713004937.

Segre:1985:HPR

[Seg85] Emilio Segrè. Historical perspective: Refugee scientists and nuclear energy. Annals of the New York Academy of Sciences, 452 (1):xv–xix, 1985. CODEN ANYAA9. ISBN 0-89766-298-9, 0-89766-299-7 (paperback). ISSN 0077-8923 (print), 1749-6632 (electronic). Sixth International Conference on Collective Phenomena: reports from the Moscow Refusnik Seminar / edited by Inga Fischer-Hjalmars and Joel L. Lebowitz. Contributions from the Moscow Refusnik Seminar and from two International Conferences on Collective Phenomena, one held in Stockholm, Sweden, 1–2 December 1983, and the other in Tel Aviv, Israel, 31 May–1 June

Sun:2015:IPF

[SG15] Xiaoxia Sun and Feng Guo. On integration by parts formula and characterization of fractional Ornstein–Uhlenbeck process. Statistics & Probability Letters, 107:170–177, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003144.

Shao:1993:LIL

[Sha93] Qi Man Shao. On the law of the iterated logarithm for infinitedimensional Ornstein–Uhlenbeck processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(??):159– 175, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shiga:1990:RCM

[Shi90] Tokuzo Shiga. A recurrence criterion for Markov processes of Ornstein–Uhlenbeck type. Probability Theory and Related Fields, 85(4):425–447, ???? 1990. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.springer.

com/article/10.1007/BF01203163.

Shimizu:2012:LAM

[Shi12] Yasutaka Shimizu. Local asymptotic mixed normality for discretely observed non-recurrent Ornstein–Uhlenbeck processes. Annals of the Institute of Statistical Mathematics (Tokyo), 64(1): 193–211, February 2012. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.springer.

com/article/10.1007/s10463-010-0307-4.

Simonian:1991:SAF

[Sim91] Alain Simonian. Stationary analysis of a fluid queue with input rate varying as an Ornstein–Uhlenbeck process. SIAM Journal on Applied Mathematics, 51(3):828–842, June 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Simonian:1995:ADE

[Sim95] Alain Simonian. Asymptotic distribution of exit times for smallnoise diffusions. SIAM Journal on Applied Mathematics, 55(3): 809–826, June 1995. CODEN SMJMAP. ISSN 0036-1399 (print),

1095-712X (electronic). URL http://epubs.siam.org/sam-bin/
dbq/article/24460.

Simon:2011:ACM

[Sim11] Thomas Simon. On the absolute continuity of multidimensional Ornstein–Uhlenbeck processes. Probability Theory and Related Fields, 151(1–2):173–190, October 2011. CODEN PTRFEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http://link.

springer.com/article/10.1007/s00440-010-0296-5.

Scully:1999:DCT

[SK99] Marlan O. Scully and Kishor T. Kapale. Does the critical temperature concept make sense for Bose Einstein condensation in mesoscopic systems? or Is there an easy resolution to the Uhlenbeck dilemma? In Relativity, particle physics and cosmology (College Station, TX, 1998), pages 181–189. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1999.

Seaquist:2013:APO

[SK13] Thomas Seaquist and Andrzej Korzeniowski. The American put with an Ornstein–Uhlenbeck model under Markovian switching. Communications in Applied Analysis, 17(3-4):379–394, 2013. ISSN 1083-2564.

Smilga:2009:CDP

[Smi09] Andrei V. Smilga. Comments on the dynamics of the Pais– Uhlenbeck oscillator. SIGMA Symmetry Integrability Geom. Methods Appl., 5:Paper 017, 13, 2009. ISSN 1815-0659.

Sarkar:2017:CFP

[SSMS17] Kaushik Sarkar, Nayem Sk, Ranajit Mandal, and Abhik Kumar Sanyal. Canonical formulation of Pais–Uhlenbeck action and resolving the issue of branched Hamiltonian. Int. J. Geom. Methods Mod. Phys., 14(3):1750038, 23, 2017. ISSN 0219-8878.

Stephane:2003:WCU

[Sté03] Dellacherie Stéphane. On the Wang Chang–Uhlenbeck equations. Discrete Contin. Dyn. Syst. Ser. B, 3(2):229–253, 2003. ISSN 1531-3492.

Striebel:1958:EET

[Str58] Charlotte T. Striebel. On the efficiency of estimates of trend in the Ornstein–Uhlenbeck process. Annals of Mathematical Statistics, 29(1):192–200, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706716.

Scott:1942:TCR

[SU42] W. T. Scott and G. E. Uhlenbeck. On the theory of cosmicray showers. II. Further contributions to the fluctuation problem. Physical Review (2), 62(11-12):497–508, December 1942. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Scott:1943:TCR

[SU43] William T. Scott and George E. Uhlenbeck. On the Theory of Cosmic Ray Showers. Lancaster Press, Lancaster, PA, USA, 1943. ???? pp.

Sullivan:1978:SGC

[Sul78] Walter Sullivan. Samuel Goudsmit, codiscoverer of electron’s spin, is dead at 76: Many European ties. Discovery at age 11. New York Times, ??(??):B6, December 6, 1978. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/123572175/.

Sullivan:1988:GUD

[Sul88] Walter Sullivan. George E. Uhlenbeck dies at 87; co-discoverer of electron’s spin. New York Times, ??(??):D27, November 2, 1988. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL https://search.proquest.com/hnpnewyorktimes/
docview/110452184/.

Suslin:1986:SCP

[Sus86] V. M. Suslin. Solvability of the Cauchy problem for a homogeneous kinetic equation with the Ühling–Uhlenbeck collision integral. Akademiya Nauk SSSR. Institut Prikladnoĭ Matematiki. Preprint, (38):13, 1986.

Suslin:1989:EUT

[Sus89] V. M. Suslin. Existence and uniqueness theorems for an inhomogeneous kinetic equation with the Ühling–Uhlenbeck collision integral (the case of Fermi–Dirac statistics). Akad. Nauk SSSR Inst. Prikl. Mat. Preprint, (56):17, 1989.

Suslin:1993:ETK

[Sus93] V. M. Suslin. Existence theorems for kinetic equations with the Uehling–Uhlenbeck collision integral. Doklady Akademii nauk SSSR, 328(5):567–569, 1993. CODEN DANKAS. ISSN 0869-5652.

Song:2015:FHT

[SWW15] Shiyu Song, Suxin Wang, and Yongjin Wang. First hitting times for doubly skewed Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 96(??):212–222, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S016771521400340X.

Szablowski:2012:WOU

[Sza12] P. J. Szabłowski. q-wiener and (α, q)-Ornstein–Uhlenbeck processes. A generalization of known processes. Theory of Probability and its Applications, 56(4):634–659, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Szablowski:2013:WOU

[Sza13] P. J. Szabłowski. q-Wiener and (α, q)-Ornstein–Uhlenbeck processes. A generalization of known processes. Theory of Probability and its Applications, 56(4):634–659, ???? 2013. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Taylor:1989:MEC

[Tay89] J. C. Taylor. The minimal eigenfunctions characterize the Ornstein–Uhlenbeck process. Annals of Probability, 17(3): 1055–1062, July 1989. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL http://projecteuclid.

org/euclid.aop/1176991256.

Thomas:1926:MSE

[Tho26] L. H. (Llewellyn Hilleth) Thomas. The motion of the spinning electron. Nature, 117(2945):514, April 10, 1926. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Thomas:1975:SMF

[Tho75] Marlin U. Thomas. Some mean first-passage time approximations for the Ornstein–Uhlenbeck process. Journal of Applied Probability, 12(3):600–604, September 1975. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3212877.

Thomas:2011:LBF

[Tho11] Peter J. Thomas. A lower bound for the first passage time density of the suprathreshold Ornstein–Uhlenbeck process. Journal of Applied Probability, 48(2):420–434, June 2011. CODEN JPRBAM.

ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/23065894.

Taufer:2009:SLD

[TL09] Emanuele Taufer and Nikolai Leonenko. Simulation of Lévydriven Ornstein–Uhlenbeck processes with given marginal distribution. Computational Statistics & Data Analysis, 53(6):2427– 2437, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print),

1872-7352 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167947308001461.

Tong:2012:LTM

[TLZ12] Changqing Tong, Zhengyan Lin, and Jing Zheng. The local time of the Markov processes of Ornstein–Uhlenbeck type. Statistics & Probability Letters, 82(7):1229–1234, July 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715212000752.

Taillefumier:2008:HLC

[TM08] Thibaud Taillefumier and Marcelo O. Magnasco. A Haar-like construction for the Ornstein Uhlenbeck process. Journal of Statistical Physics, 132(2):397–415, July 2008. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-008-9545-8.

Tran:2017:LPE

[Tra17] Ngoc Khue Tran. LAN property for an ergodic Ornstein– Uhlenbeck process with Poisson jumps. Communications in Statistics: Theory and Methods, 46(16):7942–7968, 2017. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1167908.

Uhlenbeck:1963:SP

[U^{+}63] G. E. Uhlenbeck et al. Statistical Physics, volume 3 of Brandeis University summer institute. Lectures in theoretical physics 1962. Benjamin, New York, NY, USA, 1963. vii + 252 pp. Notes by N. Ranganthan and others.

Uhlenbeck:1936:QTN

[UB36] G. E. Uhlenbeck and E. Beth. The quantum theory of the nonideal gas I. Deviations from the classical theory. Physica, 3(8):729– 745, August 1936. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Uhlenbeck:1937:QTN

[UB37] G. E. Uhlenbeck and E. Beth. The quantum theory of the nonideal gas II. Behavior at low temperatures. Physica, 4(10):915– 924, October 1937. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Uhlenbeck:1952:TNM

[UC52] G. E. Uhlenbeck and C. S. Wang Chang. Is there a neutral μmeson? Physical Review (2), 85(4):684, February 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Uhlenbeck:1960:LSM

[UF60] G. E. Uhlenbeck and G. W. Ford. Lectures in statistical mechanics. In Kac [Kac60], page ?? LCCN QC20 .S84 1960. With an appendix on quantum statistics of interacting particles, by E. W. Montroll.

Uhlenbeck:1962:TLG

[UF62] G. E. Uhlenbeck and G. W. Ford. The theory of linear graphs with applications to the theory of the virial development of the properties of gases. In Studies in Statistical Mechanics [dBU62], pages 119–211. ISBN 0-7204-1105-X. LCCN QC175 S77.

Uhlenbeck:1963:KTD

[UF63a] G. E. Uhlenbeck and G. W. Ford. The kinetic theory of dense gases. In ????, editor, Lectures in Statistical Mechanics, pages 118–142. American Mathematical Society, Providence, RI, USA, 1963.

Uhlenbeck:1963:LSM

[UF63b] G. E. Uhlenbeck and G. W. Ford. Lectures in Statistical Mechanics, volume 1 of Lectures in Applied Mathematics: proceedings of the Summer Seminar, Boulder, Colorado, 1960. American Mathematical Society, Providence, RI, USA, 1963. ISBN 0-8218-1101-0. x + 181 pp. LCCN QC20 .S84 1960, v.1; QC175 .U31 1974. With an appendix on quantum statistics of interacting particles by E. W. Montroll.

Ulenbek:1965:LPS

[UF65] Dž. Ulenbek and Dž. Ford. Lekcii po statističeskoj mechanik. (Russian) [Lectures on statistical mechanics]. Izdatelsvo Mir, Moscow, USSR, 1965. 307 pp.

Uhlenbeck:1986:LSM

[UF86] G. E. Uhlenbeck and G. W. Ford. Lectures in Statistical Mechanics, volume 1 of Lectures in Applied Mathematics: proceedings of the Summer Seminar, Boulder, Colorado, 1960. American Mathematical Society, Providence, RI, USA, 1986. ISBN 0-8218-1101-0. x + 181 pp. LCCN QC20 .S84 1960, v.1; QC175 .U31 1974. With an appendix by E. W. Montroll.

Uhlenbeck:1925:EHU

[UG25] George E. Uhlenbeck and Samuel Goudsmit. Ersetzung der Hypothese vom unmechanischen Zwang durch eine Forderung bezüglich des inneren Verhaltens jedes einzelnen Elektrons. (German) [Replacement of the hypothesis of unmechanical force with a requirement for the inner behavior of each electron]. Die Naturwissenschaften, 13(47):953–954, November 1925. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://link.springer.com/article/10.1007/BF01558878;
http://www.springerlink.com/content/u20q85248gp8166q/.

See [Lor36] for a computation in 1927 of the radius of a rotating charged particle with the spin of the electron.

Uhlenbeck:1926:SES

[UG26] G. E. Uhlenbeck and S. Goudsmit. Spinning electrons and the structure of spectra. Nature, 117(2938):264–265, February 20, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). The paper is followed by a half-column supportive letter from Niels Bohr. See [Tho26] for a relativistic correction to the Zeeman splitting of spectral lines.

Uhlenbeck:1929:PBM

[UG29] G. E. Uhlenbeck and S. Goudsmit. A problem in Brownian motion. Physical Review (2), 34(1):145–151, July 1929. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Uhlenbeck:1932:ESN

[UG32] G. E. Uhlenbeck and L. Gropper. The equation of state of a nonideal Einstein–Bose or Fermi–Dirac gas. Physical Review (2), 41 (1):79–90, July 1932. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Uhlenbeck:1935:SED

[UG35] G. E. Uhlenbeck and S. Goudsmit. Statistical energy distribution for a small number of particles. In ????, editor, Zeeman Verhandelingen, pages 201–211. Martinus Nijhoff Publishers, Dordrecht, The Netherlands, 1935.

Uhlenbeck:1967:SES

[UG67] G. E. Uhlenbeck and S. Goudsmit. Spinning electrons and the structure of spectra. In W. R. (William Russell) Hindmarsh, editor, Atomic Spectra, The Commonwealth and International Library: Selected Readings in Physics, pages 253–258. Pergamon Press, Oxford, UK and New York, NY, USA, 1967. ISBN 0-08-012103-9, 1-4831-5580-3 (e-book). LCCN QC454 .H626

1967. URL http://www.sciencedirect.com/science/article/

pii/B9780080121031500220. Published in Nature 117, 264 (1926).

Uhlenbeck:1984:SES

[UGB84] G. E. Uhlenbeck, S. Goudsmit, and Niels Bohr. Spinning electrons and the structure of spectra. In Klaus Stolzenburg, editor, The Emergence of Quantum Mechanics (Mainly 1924–1926), volume 5 of Niels Bohr Collected Works, chapter VI, pages 287– 289. Elsevier, Amsterdam, The Netherlands, 1984. ISSN 1876-

0503. URL http://www.sciencedirect.com/science/article/
pii/S1876050308705358.

Uhlenbeck:1965:CPV

[UH65] G. E. Uhlenbeck and P. C. Hemmer. On critical phenomena in a van der Waals gas. In Meixner [Mei65], pages 241–252.

Uhlenbeck:1963:VWT

[UHK63] G. E. Uhlenbeck, P. C. Hemmer, and M. Kac. On the van der Waals theory of the vapor–liquid equilibrium. II. Discussion of the distribution functions. Journal of Mathematical Physics, 4(2):229– 247, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v4/i2/p229_s1.

Uhlenbeck:1924:JHD

[Uhl24] George E. Uhlenbeck. Over Johannes Heckius. (Dutch) [On Johannes Heckius]. Mededeelingen van het Nederlandsch Historisch Instituut te Rome, 4:217–228, 1924. URL http:

//resources.huygens.knaw.nl/retroboeken/knir/#page=11&
accessor=toc1&view=imagePane&size=686&source=4.

Uhlenbeck:1925:SVL

[Uhl25] G. E. Uhlenbeck. Over een stelling van Lorentz en haar uitbreiding voor meerdimensionale ruimten. (Dutch) [On a statement by Lorentz and its extension for multi-dimensional spaces]. Physica (The Hague), 5(??):423–428, ???? 1925. CODEN PYSIA7. ISSN 0370-2707.

Uhlenbeck:1927:SMT

[Uhl27] George Eugene Uhlenbeck. Over statistische Methoden in de Theorie der Quanta. (Dutch) [On Statistical Methods in the Theory of Quanta]. Ph.D. thesis, Universiteit Leiden, Leiden, The Nether-

lands, 1927. viii + 97 pp. URL http://ilorentz.org/history/
proefschriften/sources/Uhlenbeck_1927.pdf;   https://
genealogy.math.ndsu.nodak.edu/id.php?id=58901.

Uhlenbeck:1935:PPC

[Uhl35] G. E. Uhlenbeck. The probability of position in a canonical ensemble. Journal of Mathematics and Physics / Massachusetts Institute of Technology, 14(1–4):10–16, ???? 1935. CODEN JMPHA9. ISSN 0097-1421.

Uhlenbeck:1936:PVB

[Uhl36] George E. Uhlenbeck. Het principe van behoud van energie. (Dutch) [The principle of energy conservation]. Nijhoff, ’s-Gravenhage, The Netherlands, 1936. 16 pp. Inaugurele rede Rijksuniversiteit Utrecht. (Dutch) [Inaugural lecture at the State University of Utrecht].

Uhlenbeck:1940:BRP

[Uhl40] George E. Uhlenbeck. Book review: The Principle of Statistical Mechanics by Richard C. Tolman, Oxford University Press, 1939. Science, 92(2387):287–288, September 27, 1940. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL https://

www.jstor.org/stable/1667342.

Uhlenbeck:1941:BRM

[Uhl41] George E. Uhlenbeck. Book review: The Mathematical Theory of Non-uniform Gases. By S. Chapman and T. G. Cowling. Journal of Physical Chemistry, 45(5):876–877, May 1941. CODEN JPCHAX. ISSN 0022-3654 (print), 1541-5740 (electronic).

Uhlenbeck:1950:SBP

[Uhl50] George E. Uhlenbeck. Some basic problems of statistical mechanics. American Mathematical Society, Gibbs Lecture, December 1950.

Uhlenbeck:1955:BE

[Uhl55a] G. E. Uhlenbeck. The Boltzmann equation. Report, Ft. Belvoir Defense Technical Information Center, Fort Belvoir, VA, USA, January 1955. 23 pp.

Uhlenbeck:1955:ONW

[Uhl55b] G. E. Uhlenbeck. Oude en nieuwe wagen der natuurkunde (Dutch) [Old and new ways of physics]. North-Holland Publishing Co., Amsterdam, The Netherlands, 1955. ???? pp.

Uhlenbeck:1956:RPP

[Uhl56] G. E. Uhlenbeck. Reminiscences of Professor Paul Ehrenfest. American Journal of Physics, 24(6):431–433, September 1956. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Uhlenbeck:1958:U

[Uhl58] G. E. Uhlenbeck. [unknown]. In ????, editor, Proceedings of the International Congress of Mathematics, pages 256–?? ????, ????, 1958.

Uhlenbeck:1959:U

[Uhl59a] G. E. Uhlenbeck. [unknown]. In Kac and Uhlenbeck [KU59], pages 183–203. ISBN 0-8218-0047-7. LCCN QA273 .K11 197.

Uhlenbeck:1959:BRS

[Uhl59b] George E. Uhlenbeck. Book review: Statistical Physics, Volume 5 of Course of Theoretical Physics. L. D. Landau and E. M. Lifshitz. Translated from the Russian by E. Peierls and R. F. Peierls. Pp.

484. Pergamon Press, Inc., London. Addison-Wesley Publishing Company, Inc., Reading, 1958. American Journal of Physics, 27 (5):371–372, May 1959. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Uhlenbeck:1960:BRP

[Uhl60a] G. E. Uhlenbeck. Book review: Paul Ehrenfest, Collected Scientific Papers. Martin J. Klein, Ed. Introduction by H. B. G. Casimir. North-Holland, Amsterdam; Interscience, New York, 1959. 632 pp. $13.75. Science, 131(3408):1204, April 1960. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL https://www.jstor.org/stable/1706196.

Uhlenbeck:1960:SFP

[Uhl60b] G. E. Uhlenbeck. Some fundamental problems in statistical physics. In Proceedings of the International Congress of Mathematicians. 1958, pages 256–267. Cambridge University Press, Cambridge, UK, 1960.

Uhlenbeck:1960:SAM

[Uhl60c] G. E. Uhlenbeck. Successive approximation methods in classical statistical mechanics. Physica, 26(Supplement 1):S17–S29, 1960. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Uhlenbeck:1960:SRP

[Uhl60d] George E. Uhlenbeck. Some recent progress in statistical mechanics. Physics Today, 13(7):16–21, July 1960. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Uhlenbeck:1962:ANF

[Uhl62a] G. E. Uhlenbeck. Alte und neue Fragen der statistischen mechanik. (German) [Old and new questions of statistical mechanics]. In K. Hecht, editor, Physikertagung Dusseldorf (1964). (German) [Physics days Dusseldorf (1964)], pages 1–19. Deutsche Physikalische Gesellschaft, Physik Verlag, Mosbach/Baden, West Germany, 1962.

Uhlenbeck:1962:SAK

[Uhl62b] G. E. Uhlenbeck. Some aspects of kinetic theory. In Fluid Dynamics and Applied Mathematics (Proceedings of the Symposium, University of Maryland, 1961), pages 197–207. Gordon and Breach, New York, NY, USA, 1962.

Uhlenbeck:1963:BRQ

[Uhl63a] G. E. Uhlenbeck. Book review: Quantum Mechanics. vol. 1, Old Quantum Theory. Sin-Itiro Tomonaga. Translated from the Japanese by Koshiba. North-Holland, Amsterdam; Interscience (Wiley), New York, 1962. xvi + 313 pp. Illus. $12.50. Science, 140(3569):886, May 24, 1963. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL https://www.jstor.org/

stable/1710612.

Uhlenbeck:1963:STS

[Uhl63b] G. E. Uhlenbeck. Selected topics in statistical mechanics. In Statistical Physics (Brandeis Summer Institute, 1962, Vol. 3), pages 1–90. W. A. Benjamin, Inc., New York, NY, USA, 1963.

Uhlenbeck:1964:VWR

[Uhl64] G. E. (George Eugène) Uhlenbeck. Van der Waals revisited: rede. Noord-Hollandsche Uitgevers Maatschappij, Amsterdam, The Netherlands, 1964. 16 pp. LCCN ????

Uhlenbeck:1966:SSM

[Uhl66] G. E. Uhlenbeck. Structure of statistical mechanics. In ????, editor, Proceedings 1966 Midwest Conference on Theoretical Physics (Indiana University, Bloomington, IN, 1966), pages 1–12. ????, ????, 1966.

Uhlenbeck:1968:OSM

[Uhl68] G. E. Uhlenbeck. An outline of statistical mechanics. In Cohen [Coh68], pages 1–29.

Uhlenbeck:1971:SQA

[Uhl71a] G. E. Uhlenbeck. Some questions about quantum statistical mechanics. Physica Norvegica, 5(??):139–144, ???? 1971. CODEN PHNOAH. ISSN 0031-8930.

Uhlenbeck:1971:SMQ

[Uhl71b] G. E. Uhlenbeck. Statistical mechanics and quantum mechanics. Nature, 232(5311):449–450, August 1, 1971. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Uhlenbeck:1973:BRM

[Uhl73a] G. E. Uhlenbeck. Book review: Mathematical Theory of Transport Processes in Gases (J. H. Ferziger and H. G. Kaper). SIAM Review, 15(4):804–805, ???? 1973. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Uhlenbeck:1973:VLB

[Uhl73b] G. E. Uhlenbeck. The validity and the limitations of the Boltzmann equation. In Cohen and Thirring [CT73], pages 107–119. ISBN 3-7091-8338-3 (print), 3-7091-8336-7 (e-book). ISSN 0177-8811. LCCN QC173.96-174.52. URL https://link.springer.

com/book/10.1007/978-3-7091-8336-6. See comment [Cer73].

Uhlenbeck:1973:PSP

[Uhl73c] George E. Uhlenbeck. Problems of statistical physics. In Mehra [Meh73], pages 501–513. ISBN 90-277-0345-0, 90-277-2536-5.

LCCN QC173.96 .S95 1972. URL http://www.springer.com/
us/book/9789027703453.

Uhlenbeck:1974:SPK

[Uhl74] G. E. Uhlenbeck. Some problems in the kinetic theory of dilute gases. In Anonymous [Ano74], pages 3–18. CODEN PRGDAJ. ISSN 0539-0613.

Uhlenbeck:1976:FYS

[Uhl76] George E. Uhlenbeck. Fifty years of spin: Personal reminiscences. Physics Today, 29(6):43–48, June 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v29/i6/p40_s1.

Uhlenbeck:1978:SHC

[Uhl78] George E. Uhlenbeck. Some historical and critical remarks about the theory of phase transitions. In The Ta-You Wu Festschrift: Science of matter, pages 99–107. Gordon & Breach, New York, NY, USA; Paris, France; London, UK, 1978.

Uhlenbeck:1980:SNR

[Uhl80a] G. E. Uhlenbeck. Some notes on the relation between fluid mechanics and statistical physics. In Annual Review of Fluid Mechanics, volume 12, pages 1–9. Annual Reviews, Palo Alto, CA, USA, 1980.

Uhlenbeck:1980:SRA

[Uhl80b] George E. Uhlenbeck. Some reminiscenses about Einstein’s visits to Leiden. In Woolf [Woo80], pages 524–525. ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Uhlenbeck:1980:TT

[Uhl80c] George E. Uhlenbeck. Time for thought. Nature, 288(5788):305– 306, November 20, 1980. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Uhlenbeck:1937:NLS

[UK37] G. E. Uhlenbeck and H. Kuiper. Note on the law of Sargent. Physica, 4(7):601–605, July 1937. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Uhlenbeck:1931:NCR

[UL31] G. E. Uhlenbeck and Otto Laporte. New covariant relations following from the Dirac equations. Physical Review (2), 37(11):1552– 1554, June 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Uhlenbeck:1930:TBM

[UO30] G. E. Uhlenbeck and L. S. Ornstein. On the theory of the Brownian motion. Physical Review (2), 36(5):823–841, September 1930. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Ustunel:1995:GSD

[Üst95] Ali Süleyman Üstünel. Gross–Sobolev derivative, divergence and Ornstein–Uhlenbeck operator. Lecture Notes in Mathematics, 1610:9–18, 1995. CODEN LNMAA2. ISBN 3-540-60170-8 (print), 3-540-44662-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0096330/.

Uhlenbeck:1932:VSF

[UU32] G. E. Uhlenbeck and E. A. Uehling. The velocity of sound in a Fermi–Dirac or Einstein–Bose ideal gas. Physical Review (2), 39 (6):1014–1015, March 1932. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Uehling:1933:TPE

[UU33] E. A. Uehling and G. E. Uhlenbeck. Transport phenomena in Einstein–Bose and Fermi–Dirac gases. I. Physical Review (2), 43 (7):552–561, April 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Uhlenbeck:1930:VDE

[UY30] G. E. Uhlenbeck and L. A. Young. The value of e/m by deflection experiments. Physical Review (2), 36(12):1721–1727, December 1930. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Vakeroudis:2012:HTW

[Vak12] S. Vakeroudis. On hitting times of the winding processes of planar Brownian motion and of Ornstein–Uhlenbeck processes via Bougerol’s identity. Theory of Probability and its Applications, 56(3):485–507, ???? 2012. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

vanderWaals:1935:PZM

[van35] J. D. van der Waals, Jr., editor. Pieter Zeeman: 1865–25 Mei 1935; verhandelingen op 25 Mei 1935 aangeboden aan P. Zeeman. Nijhoff, ’s-Gravenhage, The Netherlands, 1935. xi + 423 pp.

vanBargen:2009:AGS

[vB09] Holger van Bargen. Asymptotic growth of spatial derivatives of isotropic flows. Electronic Journal of Probability, 14:80:2328– 80:2351, 2009. CODEN ???? ISSN 1083-6489. URL http://ejp.

ejpecp.org/article/view/704.

vanBerkel:2013:UGE

[vB13] K. van Berkel. Uhlenbeck, George Eugène (1900–1988). In Biografisch Woordenboek van Nederland, volume 6, page ?? Huygens ING, Amsterdam, The Netherlands, November 12, 2013. URL

http://resources.huygens.knaw.nl/bwn1880-2000/lemmata/
bwn6/uhlenbeck.

vanCalmthout:2016:SGJ

[vC16] Martijn van Calmthout. Sam Goudsmit: de jacht op de atoombom van Hitler. Meulenhoff, Amsterdam, The Netherlands, 2016. ISBN 90-290-8958-X (paperback), 94-023-0744-3 (e-book). 284 + 16 pp. LCCN DS135.N6 G656 2016.

vanCalmthout:2018:SGH

[vCH18] Martijn van Calmthout and Michiel Horn, editors. Sam Goudsmit and the hunt for Hitler’s atom bomb. Prometheus Books, Amherst, NY, USA, 2018. ISBN 1-63388-450-3 (hardcover), 1-63388-451-1 (e-book). 243 + 14 pp. LCCN DS135.N6 G65613 2018.

Veestraeten:2015:RFM

[Vee15] Dirk Veestraeten. A recursion formula for the moments of the first passage time of the Ornstein–Uhlenbeck process. Journal of Applied Probability, 52(2):595–601, June 2015. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

projecteuclid.org/euclid.jap/1437658618.

vanLear:1931:BMS

[vLU31] G. A. van Lear and G. E. Uhlenbeck. The Brownian motion of strings and elastic rods. Physical Review (2), 38(9):1583–1598, November 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Vrins:2016:CFT

[Vri16] Frédéric Vrins. Characteristic function of time-inhomogeneous Lévy-driven Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 116(??):55–61, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715215303849.

VanLier:1937:SCD

[VU37] C. Van Lier and G. E. Uhlenbeck. On the statistical calculation of the density of the energy levels of the nuclei. Physica, 4(7): 531–542, July 1937. CODEN PHYSAG. ISSN 0031-8914 (print),

1873-1767 (electronic). URL https://www.sciencedirect.com/
science/article/pii/S0031891437800981.

Wang:2011:COU

[Wan11a] Feng-Yu Wang. Coupling for Ornstein–Uhlenbeck processes with jumps. Bernoulli: official journal of the Bernoulli Society for Mathematical Statistics and Probability, 17(4):1136–1158, November 2011. CODEN ???? ISSN 1350-7265 (print), 1573-

9759 (electronic). URL http://projecteuclid.org/euclid.bj/
1320417499.

Wang:2011:HIO

[Wan11b] Jian Wang. Harnack inequalities for Ornstein–Uhlenbeck processes driven by Lévy processes. Statistics & Probability Letters, 81(9):1436–1444, September 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715211001593.

Wang:2012:EEL

[Wan12] Jian Wang. On the exponential ergodicity of Lévy–driven Ornstein–Uhlenbeck processes. Journal of Applied Probability, 49 (4):990–1004, December 2012. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/41713824.

Wang:2016:LJO

[Wan16] Yizao Wang. Large jumps of q-Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 118(??):110–116, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301067.

Wang:1997:IRP

[WC97] H. Wang and X. Chen. On the interval recurrence property of (N, d)–Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 33(1):79–84, April 15, 1997. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715296001125.

Wenocur:1990:OUP

[Wen90] Michael L. Wenocur. Ornstein–Uhlenbeck process with quadratic killing. Journal of Applied Probability, 27(3):707–712, September 1990. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3214554.

Weiner:1972:EHN

[WH72] Charles Weiner and Elspeth Hart, editors. Exploring the history of nuclear physics: proceedings of the American Institute of Physics– American Academy of Arts and Sciences conferences on the history of nuclear physics, 1967 and 1969, volume 7 of AIP conference proceedings. American Institute of Physics, Woodbury, NY, USA, 1972. ISSN 0094-243X, 1551-7616, 1935-0465. LCCN QC173 .E88 1972.

Wilkinson:2010:PTS

[Wil10] Michael Wilkinson. Perturbation theory for a stochastic process with Ornstein–Uhlenbeck noise. Journal of Statistical Physics, 139(2):345–353, April 2010. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-010-9944-5.

Wittig:1984:GOU

[Wit84] Timothy A. Wittig. A generalized Ornstein–Uhlenbeck process. Communications in Statistics: Theory and Methods, 13(1):29–43, 1984. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Woolf:1980:SSP

[Woo80] Harry Woolf, editor. Some strangeness in the proportion: a centennial symposium to celebrate the achievements of Albert Einstein. Addison-Wesley, Reading, MA, USA, 1980. ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Wilkinson:2011:SOU

[WP11] Michael Wilkinson and Alain Pumir. Spherical Ornstein– Uhlenbeck processes. Journal of Statistical Physics, 145(1): 113–142, October 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-011-0332-6.

Wang:2015:SOU

[WSW15] Suxin Wang, Shiyu Song, and Yongjin Wang. Skew Ornstein– Uhlenbeck processes and their financial applications. Journal of Computational and Applied Mathematics, 273(??):363–382, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042714003045.

Wang:2015:EPO

[WTZ15] Yinfeng Wang, Yanlin Tang, and Xinsheng Zhang. Estimation of parameters of the Ornstein–Uhlenbeck type processes with continuum of moment conditions. Communications in Statistics: Theory and Methods, 44(24):5189–5203, 2015. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wolfe:1934:SDP

[WU34a] Hugh C. Wolfe and G. E. Uhlenbeck. Spontaneous disintegration of proton or neutron according to the Fermi theory. Physical Review (2), 46(3):237, August 1934. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Wu:1934:DOP

[WU34b] Ta-You Wu and George E. Uhlenbeck. Disintegration of Li^{6} by protons and deutons. Physical Review (2), 45(8):553–554, April 1934. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Wang:1945:TBMc

[WU45a] Ming Chen Wang and G. E. Uhlenbeck. On the theory of the Brownian motion. II. Reviews of Modern Physics, 17:323–342, 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Wang:1945:TBMa

[WU45b] Ming Chen Wang and George E. Uhlenbeck. On the theory of the Brownian motion II. Reviews of Modern Physics, 17 (2–3):323–342, April 1945. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.17.323;   http://
rmp.aps.org/abstract/RMP/v17/i2-3/p323_1.

Wang:1945:TBMb

[WU45c] Ming Chen Wang and George E. Uhlenbeck. On the Theory of the Brownian Motion II. Lancaster Press, Lancaster, PA, USA, 1945. ???? pp.

Wang:1964:U

[WU64] C. S. Wang and G. E. Uhlenbeck. [unknown]. In Studies in Statistical Mechanics [dBU64], pages 743–?? LCCN ????

Wang:2015:AEP

[WX15] Ran Wang and Lihu Xu. Asymptotics of the entropy production rate for d-dimensional Ornstein–Uhlenbeck processes. Journal of Statistical Physics, 160(5):1336–1353, September 2015. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-015-1295-9.

Wong:2011:ODB

[WZ11] Hoi Ying Wong and Jing Zhao. Optimal dividends and bankruptcy procedures: Analysis of the Ornstein–Uhlenbeck process. Journal of Computational and Applied Mathematics, 236(2):150–166, August 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042711003372.

Xing:2012:ASA

[Xin12] Fei Xing. Almost sure asymptotic for Ornstein–Uhlenbeck processes of Poisson potential. Statistics & Probability Letters, 82(12):2091–2102, December 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212002842.

Xing:2012:NTD

[XXY12] Xiaoyu Xing, Yongsheng Xing, and Xuewei Yang. A note on transition density for the reflected Ornstein–Uhlenbeck process. Statistics & Probability Letters, 82(3):586–591, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003737.

Xing:2009:SDT

[XZW09] Xiaoyu Xing, Wei Zhang, and Yongjin Wang. The stationary distributions of two classes of reflected Ornstein–Uhlenbeck processes. Journal of Applied Probability, 46(3):709–720, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/25662456.

Yano:2017:FACb

[Yan17] Ryosuke Yano. Fast and accurate calculation of dilute quantum gas using Uehling–Uhlenbeck model equation. Journal of Computational Physics, 330(??):1010–1021, February 1, 2017. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999116305794.

Young:1930:WBK

[YU30] L. A. Young and G. E. Uhlenbeck. On the Wentzel–Brillouin– Kramers approximate solution of the wave equation. Physical Review (2), 36(7):1154–1167, October 1930. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Yurachkivsky:2017:LDG

[Yur17] Andriy Yurachkivsky. Limit distribution of a generalized Ornstein–Uhlenbeck process. International Journal of Statistics and Probability, 6(1):24–??, ???? 2017. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (electronic). URL http://www.ccsenet.

org/journal/index.php/ijsp/article/view/65098.

Zhang:2012:EPC

[ZGO12] B. Zhang, L. A. Grzelak, and C. W. Oosterlee. Efficient pricing of commodity options with early-exercise under the Ornstein–Uhlenbeck process. Applied Numerical Mathematics: Transactions of IMACS, 62(2):91–111, February 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0168927411001863.

Zhang:2011:EST

[Zha11] Shibin Zhang. Exact simulation of tempered stable Ornstein– Uhlenbeck processes. Journal of Statistical Computation and Simulation, 81(11):1533–1544, 2011. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Ziff:1977:IBE

[ZUK77] R. M. Ziff, G. E. Uhlenbeck, and M. Kac. The ideal Bose–Einstein gas — revisited. Physics Reports, 32C(4):169–248, 1977. CODEN PRPLCM. ISSN 0370-1573 (print), 1873-6270 (electronic).

Zhang:2017:SDO

[ZW17] Zhenzhong Zhang and Wenlong Wang. The stationary distribution of Ornstein–Uhlenbeck process with a two-state Markov switching. Communications in Statistics: Simulation and Computation, 46 (6):4783–4794, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2017:OTL

[ZWB17] Jiang Zhou, Lan Wu, and Yang Bai. Occupation times of Lévydriven Ornstein–Uhlenbeck processes with two-sided exponential jumps and applications. Statistics & Probability Letters, 125(??): 80–90, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217300433.

Zhang:2009:TLT

[ZZ09] Shibin Zhang and Xinsheng Zhang. On the transition law of tempered stable Ornstein–Uhlenbeck processes. Journal of Applied Probability, 46(3):721–731, September 2009. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/25662457.

Zhang:2013:LSE

[ZZ13] Shibin Zhang and Xinsheng Zhang. A least squares estimator for discretely observed Ornstein–Uhlenbeck processes driven by symmetric α-stable motions. Annals of the Institute of Statistical Mathematics (Tokyo), 65(1):89–103, February 2013. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10463-012-0362-0.

Zang:2016:GLB

[ZZ16] Qing-Pei Zang and Li-Xin Zhang. A general lower bound of parameter estimation for reflected Ornstein–Uhlenbeck processes. Journal of Applied Probability, 53(1):22–32, March 2016. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://projecteuclid.org/euclid.jap/1457470555.

Zang:2019:ABT

[ZZ19] Qingpei Zang and Lixin Zhang. Asymptotic behaviour of the trajectory fitting estimator for reflected Ornstein–Uhlenbeck processes. Journal of Theoretical Probability, 32(1):183–201, March 2019. CODEN JTPREO. ISSN 0894-9840 (print), 1572-9230 (electronic).